### v1

# SWAN: Preprocessing SGD Enables Adam-Level Performance On LLM Training With Significant Memory Reduction

Abstract
Adaptive optimizers such as Adam (Kingma & Ba, 2015) have been central to the success of large language models. However, they maintain additional moving average states throughout training, which results in memory requirements several times greater than the model. This overhead imposes constraints on scalability and computational efficiency. On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency, their capability in LLM training is limited (Zhao et al., 2024b).
To address this dilemma, we show that pre-processing SGD is sufficient to reach Adam-level performance on LLMs. Specifically, we propose to preprocess the instantaneous stochastic gradients with two simple operators: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. $\mathtt{GradNorm}$ stabilizes gradient distributions, and $\mathtt{GradWhitening}$ counteracts the local curvature of the loss landscape, respectively. This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any accumulative state variables. Empirically, SWAN has the same memory footprint as SGD, achieving $\approx 50\%$ reduction on total end-to-end memory compared to Adam. In language modeling tasks, SWAN demonstrates the same or even a substantial improvement over Adam. Specifically, when pre-training the LLaMa model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity in less than half tokens seen.
1 Introduction
Adaptive optimizers, such as Adam and its variants (Kingma & Ba, 2015; Loshchilov & Hutter, 2019; Shazeer & Stern, 2018; Pagliardini et al., 2024; Liu et al., 2023; Zhao et al., 2024a), have been central to the success of training large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Dubey et al., 2024; Bi et al., 2024; Bai et al., 2023; Zhang et al., 2022). However, most adaptive optimizers for LLMs are stateful, meaning they require tracking and maintaining internal states. While achieving remarkable empirical success, these states introduce significant memory overhead. For instance, Adam (Kingma & Ba, 2015)–the de facto optimizer for LLM training–involves the tracking of exponential moving averages (EMAs), effectively doubling memory requirements. AdEMAMix (Pagliardini et al., 2024) – an extension of Adam that achieves significant convergence speed boost – requires storing even more states, tripling the memory requirements. This overhead can be significant especially in distributed settings, where the optimizer states could consume a significant amount of the GPU memory (Dubey et al., 2024; Korthikanti et al., 2023). On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency (i.e., they are stateless), their capability to train LLMs is limited (Zhao et al., 2024b), and the performance gap between SGD and Adam is significant (Zhang et al., 2020; Kunstner et al., 2023; 2024).
To address the memory-performance dilemma for LLMs, this work questions the requirement of accumulative state variables during training to achieve good performance in LLMs. We propose to simply preprocess the instantaneous SGD gradient in a history-independent manner, such that it still preserves certain desirable properties of adaptive optimizers. This result in SWAN (SGD with Whitening And Normalization), a novel SGD-like optimizer that eliminates all internal optimizer states and empirically surpasses the performance Adam on several LLM pre-training scenarios.
Our proposed optimizer consists of two stateless operators applied to the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. $\mathtt{GradNorm}$ aims to stabilize the stochasticity of gradient distributions during training, while $\mathtt{GradWhitening}$ seeks to neutralize the local geometry of the loss landscape. Both operators enables SWAN to rely solely on the statistics of the current gradient matrices. This SGD preprocessing approach eliminates all internal optimizer states, and therefore reproduces the memory consumption of SGD, reaching a new state-of-the-art in terms of optimizer memory efficiency for LLMs ($\approx 50\%$ reduction on end-to-end memory footprint vs Adam, and $\approx 100\%$ reduction in optimizer states. See Figure 1). In addition to memory savings, SWAN also demonstrates significant computational efficiency: our empirical evaluations on pre-training LLaMA models on the C4 dataset with multiple model sizes show consistently better performance than Adam and other low-rank memory-efficient optimizers. Remarkably, at the 350M and 1.3B scale, our method achieves 2X faster convergence in terms of tokens seen compared to Adam, marking SWAN the first known stateless optimizer converges as fast/faster than Adam in training large language models.
(a) 350 M LLaMA model Pretraining
(b) 1.3 B LLaMA model Pretraining
(c) End-to-end memory footprint
(d) Effective training throughput (same batch size)
Figure 1: SWAN performance preview on LLM pretraining. (a) and (b): On both 350M and 1.3B LLama architectures, SWAN achieves $>$ 2X speed-up vs Adam in terms of tokens seen. (c): memory footprint. Unlike Zhao et al. (2024a), we directly measure end-to-end memory under full-model training, with batch size = 1 sequence. SWAN achieves near-SGD optimizer memory reduction ($\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states). (d) Training throughput analysis on training 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens. We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the efficiency of optimizer relative to Adam. SWAN closely reproduces the absolute throughput of Adam, while improving the effective throughput.
Our contributions are as follows:
[…]
    1. 1.
      Robustness to ill-conditioned problems. On LLM pretraining, SWAN converges quickly even without learning rate warm-ups (Section 5.3, Section 5.6), a property not observed in Adam.
    2. 2.
      Memory efficiency: SWAN achieves the memory footprint of SGD, that is $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states when compared to Adam.
    3. 3.
      Computational efficiency: consistently outperforms Adam and reaches up to 2$\times$ acceleration and effective throughput.
  • Theoretical analysis of SWAN. We show that (1) $\mathtt{GradNorm}$ can be derived as a strategy to stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix B); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 4.3). Additionally, we highlight that $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 3, Section B.3), offering theoretical advantages over SGD and Adam (Proposition 3).
  • Fast convergence for LLM pretraining tasks. Through experiments (Section 5) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also outperforms Adam and other memory-efficient optimizers. Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
2 Preliminaries
One of the most used adaptive optimizers is the Adam optimizer (Kingma & Ba, 2015). Adam is an example of stateful optimizers, which involves tracking and maintaining internal states. It combines the advantages of two earlier methods: AdaGrad (Duchi et al., 2011), which adapts learning rates based on the historical gradient magnitude, and RMSProp (Tieleman, 2012), which mitigates the aggressive decrease in learning rates by using a decaying average of squared gradients.
Consider a loss function ${\mathcal{L}}_{{\mathbf{W}}}:{\mathcal{X}}\rightarrow\mathbb{R}$, parameterized by weight matrices ${\mathbf{W}}\in\mathbb{R}^{m\times n}$, and denote ${\bm{x}}^{(t)}$ a mini-batch of inputs provided at the $t$-th training step that is sampled from data distribution $p_{\text{data}}({\bm{x}})$. Let ${\mathbf{G}}^{(t)}$ be the stochastic gradient of ${\mathcal{L}}_{{\mathbf{W}}}$ (i.e., a random variable induced by sampling ${\bm{x}}^{(t)}$). Then, Adam can be broken down into the following steps:
${\mathbf{G}}^{(t)}$ $=\nabla_{{\mathbf{W}}}{\mathcal{L}}_{{\mathbf{W}}}({\bm{x}}^{(t)}),\quad{\bm{x}}^{(t)}\sim p_{\text{data}}({\bm{x}})\hskip 28.45274pt$ (stochastic gradient)
[…]
${\bm{\nu}}^{(t)}$ $=\beta_{2}{\bm{\nu}}^{(t-1)}+(1-\beta_{2}){{\mathbf{G}}^{(t)}}^{2},\quad\hskip 28.45274pt$ (EMA second moment)
${\mathbf{W}}^{(t+1)}$ $={\mathbf{W}}^{(t)}-\eta\big{(}\frac{\hat{{\bm{m}}}^{(t)}}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}\big{)}\hskip 28.45274pt$ (weight update)
where ${\bm{m}}^{(t)}$ and ${\bm{\nu}}^{(t)}$ are EMAs of the first and second moments of the gradients; and $\eta$ is a global step size. Essentially, Adam estimates the signal-to-noise (SNR) ratios and use it to adjust learning rates element-wise. Tracking and storing these two EMA estimates triple the total memory consumption required to train a LLM model. For example, for LLaMA 405B model, storing model weights requires 810 GB of memory, while the Adam optimizer states requires an additional 1.6TB of memory.
Desired properties of Adam.
There is a rich literature on understanding adaptive methods’ inner workings and unreasonable effectiveness. Notably, the key desired properties of Adam can be categorized as gradient smoothing, gradient invariance and gradient whitening. (1) Gradient smoothing: the EMA operations in Adam naturally reduce the influence of mini-batch noise Cutkosky & Mehta (2020); Crawshaw et al. (2022); (2) Gradient invariance: recent work suggest the performance gap between SGD and Adam might lie in Adam’s sign-descent-like nature (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023). For example, Adam is robust to the rescaling of gradient diagonals (Kingma & Ba, 2015); and is invariant to any sign-preserving scalings (under $\beta_{1}=\beta_{2}=0$)(Bernstein et al., 2018). And (3) Gradient whitening : it is well-known that the second moment EMA of Adam approximates the diagonal of Fisher information matrix (Kingma & Ba, 2015; Hwang, 2024), and biases the optimization trajectories towards well-conditioned regions (Jiang et al., 2024). For a more comprehensive discussion of these properties, please refer to appendix A.
Adam as SGD preprocessing, and the potential unnecessity of EMAs
Adam can be viewed as history-dependent preprocessing of the gradients ($\{{\mathbf{G}}^{(0)},{\mathbf{G}}^{(1)},...,{\mathbf{G}}^{(t-1)},{\mathbf{G}}^{(t)}\}\rightarrow\frac{\hat{{\bm{m}}}^{(t)}}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}$) to achieve the desired properties above. A key observation is, all of the above properties are achieved through the element-wise operations. That is, each element of the gradient matrix is pre-processed and re-scaled independently. This ignores the interactions and structures among different variables, and we hypothesize this is the reason why extra history information is need to compensate this gap, leading to the requirement of EMA states. Thus, we believe that designing stateless adaptive optimizers is possible, if we can achieve similar properties by designing new matrix-level operations that are preprocesses the instantaneous gradients of SGD (${\mathbf{G}}^{(t)}\rightarrow\tilde{{\mathbf{G}}}^{(t)}$).
3 The SWAN Optimizer: Preprocessing SGD with Normalization and Whitening
Figure 2: Illustration of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ operators. In $\mathtt{GradNorm}$ operator, we perform standardization across the output dimensions (columns), using statistics computed row-wise. In $\mathtt{GradWhitening}$ operator, we treat each column (or row) of the gradient matrix ${\mathbf{G}}$ as a separate data sample. Then, the $\mathtt{GradWhitening}$ operator can be seen as stretching/squeezing the gradient data along all eigen directions, such that the covariance matrix is an identity matrix. Note that the same approach can be applied row-wise as well.
Algorithm 1 SWAN Optimizer
  Input: weight matrix ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ with $m\leq n$. Step size $\eta$. Number of $\mathtt{GradWhitening}$ iteration $k$.
  Initialize step $t\leftarrow 0$
  repeat
     ${\mathbf{G}}^{(t)}\in\mathbb{R}^{m\times n}\leftarrow\nabla_{{\mathbf{W}}^{(t)}}\mathcal{L}^{(t)}({\mathbf{W}}^{(t)})$
     $\tilde{{\mathbf{G}}}^{(t)}\leftarrow\mathtt{GradNorm}({\mathbf{G}}^{(t)})$
[…]
     ${\mathbf{W}}^{(t)}\leftarrow{\mathbf{W}}^{(t-1)}-\eta\Delta{\mathbf{W}}^{(t-1)}$
     $t\leftarrow t+1$
  until convergence criteria met
  return  ${\mathbf{W}}^{(t)}$
Algorithm 2 $\mathtt{GradWhitening}$ Operator
  Input: Matrix ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of Newton Schultz iteration $k$, step size $\beta$
  Initialize ${\mathbf{Y}}\leftarrow{\mathbf{G}}{\mathbf{G}}^{\top}$, ${\mathbf{Z}}\leftarrow{\mathbf{I}}$
  for i = 1,…,k do
     ${\mathbf{Y}}\leftarrow\beta{\mathbf{Y}}(3{\mathbf{I}}-{\mathbf{Z}}{\mathbf{Y}})$,
     ${\mathbf{Z}}\leftarrow\beta(3{\mathbf{I}}-{\mathbf{Z}}{\mathbf{Y}}){\mathbf{Z}}$
  end for
  return  ${\mathbf{Z}}{\mathbf{G}}$
Algorithm 3 $\mathtt{GradNorm}$ Operator
  Input: Matrix ${\mathbf{G}}^{m\times n}$
  return  $\frac{{\mathbf{G}}-\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}\mathbf{1}_{n}^{\top}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\sum_{j=1}^{n}{\mathbf{G}}_{:,j})^{2}}\mathbf{1}_{n}^{\top}}$
Figure 3: SWAN Optimizer. Note that for $\mathtt{GradWhitening}$, in principle, any approximate algorithms can be applied, such as matrix Pade approximant. Here we only show one particular working example based on Newton-Schulz, which introduces minimal overheads.
As discussed in Section 2, we believe the key to designing stateless, adaptive, and effective optimizers lies in designing new matrix-level operations that exploit rich information in the gradient matrix. To this end, we propose a new stateless optimizer (Algorithm 1), SWAN (SGD with Whitening And Normalization). SWAN preprocesses SGD based on two simple operators, $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$, both are stateless, matrix-level operators. When applied in tandem, they achieve similar desirable properties of adaptive optimizers, without the need to store historical gradient moments. In this section we describe the update rules of SWAN.
3.1 SWAN Update Rules
In SWAN (Algorithm 1), the raw SGD gradient ${\mathbf{G}}^{t}$ is processed by the following operations  [^1] Here ${\mathbf{G}}$ is assumed to be the gradient matrix for some parameter block in the model (e.g. a linear layer); enabling us to take advantage of the matrix structure to achieve our smoothing and whitening goals.:
$\begin{cases}\tilde{{\mathbf{G}}}^{(t)}\leftarrow\mathtt{GradNorm}({\mathbf{G}}^{(t)})\\ \Delta{\mathbf{W}}^{(t)}\leftarrow\mathtt{GradWhitening}(\tilde{{\mathbf{G}}}^{(t)})\end{cases}$ (SWAN)
[…]
$\mathtt{GradNorm}$: gradient invariance & stabilization.
Consider the gradient matrix ${\mathbf{G}}\in\mathbb{R}^{m\times n}$ with rows and columns corresponding to input and output dimensions, respectively, of some block of model parameters. Let $1\leq i\leq m$ represent the input indices and $1\leq j\leq n$ represent the output indices. Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, we propose to standardization across the output dimensions at each time step $t$, defined as:
$\mathtt{GradNorm}({\mathbf{G}})=\frac{{\mathbf{G}}-\bar{{\bm{g}}}\mathbf{1}_{n}^{\top}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
where $\bar{{\bm{g}}}=\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}$ is the mean across output dimensions at time $t$; ${\bm{s}}=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\bar{{\bm{g}}})^{2}}$ is the standard deviation across dimensions at time $t$; $\bar{{\bm{g}}}$ and ${\bm{s}}$ are ${m}$-dimensional column vectors; and $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
$\mathtt{GradNorm}$ is an example of operators typically used in the forward pass applied to the backward pass: $\mathtt{GradNorm}$ is LayerNorm (Ba et al., 2016) (without learnable parameters) being applied to the backward gradient rather than forward activations. Similar to Layer Norm, $\mathtt{GradNorm}$ allows the optimizer to be invariant under the following transformations on gradient matrix: i) matrix-wise re-scaling; ii) matrix-wise re-centering; iii) row-wise re-scaling; and iv) row-wise re-centering. Compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) used in signed descent, $\mathtt{GradNorm}$ preserves richer information of the gradient scaling while offering invariance properties. Moreover, in Section 4 we will show that when applied to a transformer architecture, $\mathtt{GradNorm}$ also helps stabilize the gradient distribution across time.
$\mathtt{GradWhitening}$: gradient whitening.
Adam performs approximate gradient whitening element-wise, leading to well-conditioned regions. In contrast, we propose the matrix $\mathtt{GradWhitening}$ operator, which performs whitening at the matrix level. Given the gradient ${\mathbf{G}}$, the $\mathtt{GradWhitening}$ operator simply orthogonalizes ${\mathbf{G}}$  [^2] An equivalent definition is given by $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{U}}{\mathbf{V}}^{\top}$, where ${\mathbf{U}}$ and ${\mathbf{V}}$ are both orthogonal matrices defined via SVD decomposition of ${\mathbf{G}}={\mathbf{U}}{\mathbf{S}}{\mathbf{V}}^{\top}$. :
$\mathtt{GradWhitening}({\mathbf{G}}):=({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ (2)
where the exponent $-\frac{1}{2}$ stands for matrix inverse square root. One can show that $({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ is simply the orthogonalization of ${\mathbf{G}}$, i.e., the closest orthogonal matrix to ${\mathbf{G}}$ (w.r.t the Frobenius norm).
[…]
3.2 Practical Considerations
Various GPU-friendly and fast approximations are developed in the literature of decorrelated batch normalization or feature whitening. We consider the naive Netwon-Schulz (N-S) iteration algorithm (Li et al., 2018; Huang et al., 2019) to compute $\mathtt{GradWhitening}$:
${\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}),\quad{\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}$ (3)
[…]
Computational Overhead.
[…]
Rescaling $\mathtt{GradWhitening}$.
The whitening operation $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, hence might drastically change the effective learning rate. In practice, we propose the following re-scaling before updating the weight:
$\Delta{\mathbf{W}}^{(t)}\leftarrow\frac{\|\tilde{{\mathbf{G}}}^{(t)}\|\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}$ (4)
This will helps us to rescale the norm of whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, and allowing us to get reasonably good performance without extensively tuning learning rates–we found that the learning rates used Adam usually also work well for SWAN. It is also possible to remove rescaling, but that would require additional hyper-parameter search to find the optimal learning-rate for SWAN.
4 Analysis: A LLM Learning Dynamics Perspective
As a new stateless adaptive optimizer, the complete theoretical properties of SWAN is an open question which we leave for future work. However, as a first analysis, we consider SWAN from a learning dynamics perspective, specifically the dynamics of an LLM based upon a simplified transformer block. It is this analysis that led to the design of SWAN.
4.1 Setup
We assume the following simplified transformer block (STB) architecture recently proposed in Tian et al. (2023):
Definition 1 (Simplified Transformer Block (STB)).
Given the input activation ${\bm{x}}\in\mathbb{R}^{M_{C}\times 1}$, query token index $q$, context embedding matrix ${\mathbf{U}}_{C}\in\mathbb{R}^{d\times M_{C}}$, and the query embedding ${\bm{u}}_{q}\in\mathbb{R}^{d\times 1}$, the STB computes the output ${\bm{h}}\in\mathbb{R}^{n\times 1}$ as ${\bm{h}}=\phi\left({\mathbf{W}}^{\top}\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)\right)$, where $M_{C}$ is the context length, the attention logits ${\bm{z}}_{q}\in\mathbb{R}^{M_{C}\times 1}$ are given by $z_{ql}={\bm{u}}_{q}^{\top}{\mathbf{W}}_{Q}^{\top}{\mathbf{W}}_{K}u_{l}$, with ${\mathbf{W}}_{Q},{\mathbf{W}}_{K}\in\mathbb{R}^{d\times d}$ being weight matrices for the queries and keys, respectively, ${\mathbf{W}}\in\mathbb{R}^{d\times n}$ is the weight matrix for the feedforward network, and $\phi$ is a nonlinearity function such as the ReLU.
Given a STB, we consider a loss function ${\mathcal{L}}_{{\mathbf{W}},{\bm{z}}}({\bm{x}}^{(t)})$, where ${\bm{x}}^{(t)}$ is a mini-batch of inputs provided at the $t$-th training step sampled from data distribution $p_{\text{data}}({\bm{x}})$. Standard mini-batch learning dynamics is then given by
$\dot{{\mathbf{W}}}^{(t)}=\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}({\bm{x}}^{(t)})}{\partial{\mathbf{W}}},\quad\dot{{\bm{z}}_{q}}^{(t)}=\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}({\bm{x}}^{(t)})}{\partial{\bm{z}}_{q}}\;.$
In this case, both $\dot{{\mathbf{W}}}^{(t)}$ and $\dot{{\bm{z}}_{q}}$ are viewed as random variables induced by random mini-batch ${\bm{x}}^{(t)}$. For example, for each row $i$, $\dot{{\mathbf{W}}}^{(t)}[i,:]$ can be re-written as $\dot{{\mathbf{W}}}^{(t)}[i,:]=\mathbb{E}[\dot{{\mathbf{W}}}^{(t)}[i,:]]+{\bm{\varepsilon}}_{{\mathbf{W}}}^{(t)}[i,:]$, where ${\bm{\varepsilon}}^{(t)}[i,:]$ is zero mean random variable with covariance $\mathrm{Cov}[\dot{{\mathbf{W}}}^{(t)}[i,:]]$.
4.2 $\mathtt{GradNorm}$ stabilizes gradient distributions of simplified transformers
[…]
Theorem 1 ($\mathtt{GradNorm}$ stabilizes gradient distributions across time for the STB).
Consider the STB (Definition 1). Assuming we inherit the assumptions in Theorem 1 of Tian et al. (2023), as described in Appendix B. Then consider ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, the composition of the MLP project-up matrix and the embedding matrix as a whole. Then, its standardized stochastic gradients $\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}}({\bm{x}}^{(t)})}{{\mathbf{U}}_{C}^{\mathbf{W}}})$ satisfy:
$\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{1})}]=\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{2})}]\quad\text{for all }t_{1},t_{2},\text{and}\,i.$
In other words, the covariance structure of $\tilde{{{\mathbf{G}}}}$ is identical across all time steps $t$, achieving distributional stability across time. The same relationship also holds for the gradient of attention score $\tilde{{\mathbf{G}}}_{{\bm{z}}_{q}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}({\bm{x}}^{(t)})}{\partial{\bm{z}}_{q}})$.
Theorem 1 suggests that $\mathtt{GradNorm}$ implicitly aligns with the dynamics of transformer architectures and removes the time-heterogeneity in gradient covariance structures.
4.3 $\mathtt{GradWhitening}$ Is An Efficient Non-diagonal Second-Order Update
In this section, we show that $\mathtt{GradWhitening}$ is equivalent to non-diagonal approximate second-order method, under a specific Kronecker factorization assumption of the Hessian. The assumption is as below:
Assumption 1 (Assumption of $\mathtt{GradWhitening}$).
At time $t$, the local Hessian ${\mathbf{H}}$ of the loss has shared block-diagonal structure, such that ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
By leveraging this structural assumption, we can now effectively estimate ${\mathbf{H}}$ by only using one single gradient matrix sample $\bm{G}:=[{\bm{g}}_{1},\dots,{\bm{g}}_{n}]\in\mathbb{R}^{m\times n}$. Recall that the Fisher information formulation of Hessian is defined as ${\mathbf{H}}=\mathbb{E}(\text{Vect}({\mathbf{G}})\text{Vect}({\mathbf{G}})^{\top})$ where $\text{Vect}(\cdot)$ denotes the vectorized operation. Under assumption 1, we can estimate ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, using a single sample of the gradient, that is we can compute the following simple estimate $\frac{1}{n}\sum_{i=1}^{n}{\bm{g}}_{i}{\bm{g}}_{i}^{T}=\bm{G}\bm{G}^{\top}$, which approximate the current $\tilde{{\mathbf{H}}}$. Hence, $\mathtt{GradWhitening}$ can be seen as applying a second order update under our structural assumption:
$\mathtt{GradWhitening}({\mathbf{G}})=({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}=\tilde{{\mathbf{H}}}^{-\frac{1}{2}}{\mathbf{G}}$
In Theorem 2 of Appendix, we will further show that the pre-conditioned update rule $\tilde{{\mathbf{H}}}^{-\frac{1}{2}}{\mathbf{G}}$ indeed induces a Newton-like behavior. More importantly, in the following Proposition, we show that the implicit assumption of $\mathtt{GradWhitening}$ is a good abstraction of LLM dynamics since it in fact aligns with the equilibrium Hessian structure in the STB regime:
Proposition 1 (Shared structures in the block-diagonal of Hessians at transformer equilibrium).
Consider a STB (1), trained with full-batch gradient descent. Next, assume we inherit all the assumptions from Theorem 1 of Tian et al. (2023). Then, as $t\xrightarrow[]{}\infty$, we have the following shared Hessian structure along the diagonal blocks:
$\frac{{\mathbf{H}}_{sk,s^{\prime}k}}{\sum_{s,s^{\prime}}{\mathbf{H}}_{sk,s^{\prime}k}}\xrightarrow[]{}\frac{{\mathbf{H}}_{sk^{\prime},s^{\prime}k^{\prime}}}{\sum_{s,s^{\prime}}{\mathbf{H}}_{sk^{\prime},s^{\prime}k^{\prime}}},\quad\forall 1\leq s,s^{\prime}\leq d,1\leq k,k^{\prime}\leq n$ (5)
Where ${\mathbf{H}}({\mathbf{W}})_{sk,s^{\prime}k^{\prime}}=\frac{\partial\mathcal{L}}{\partial{w_{sk}}\partial{w_{s^{\prime}k^{\prime}}}}$.
Proposition 1 shows that, under a simplified setting of the transformer, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares the identical structure, which supports the implicit assumption of $\mathtt{GradWhitening}$. This result is verified in our numerical experiment (Appendix, Figure 9).
Finally, Theorem 3 presented in Appendix will also how $\mathtt{GradWhitening}$ helps to make the convergence rate of SGD more robust to the condition number of local curvatures, and outperforms both SGD and Adam in the ill-conditioned regime.
5 Experiments
In this section, we perform empirical studies on SWAN. The main goals of those studies are as follows:
  • Verifying whether $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ indeed effectively fulfilled their design goals (stabilizing gradient noise, and modeling/counteracting local curvatures). This is performed in Section 5.1 and Section 5.2.
  • Verifying whether SWAN can consistently offer Adam-level or even better performance, while maintaining low memory cost and high-throughput. This is performed in Section 5.3, Section 5.4 and Section 5.5, respectively.
  • Ablation study (Section 5.6).
All experiments run on NVIDIA A100 GPUs.
5.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
Figure 4: (a) - (c): KL Divergence of gradient distributions across training Steps. We use the projection weights in attention and MLP modules of the second layer as an example. The plots compare standard training with $\mathtt{GradNorm}$-augmented training. Lower KL divergence values indicate greater stability in gradient distributions. (d): Training loss curve comparing standard training vs $\mathtt{GradNorm}$-augmented training
To investigate whether $\mathtt{GradNorm}$ indeed stabilizes stochastic gradient distributions as claimed in Theorem 1, we conduct a series of controlled experiments using a scaled-down version of the LLaMA architecture. We aim to assess the impact of $\mathtt{GradNorm}$ on the gradient distribution during training by analyzing the statistical properties of gradients across multiple training steps. Specifically, we employ a small-scale LLaMA-based model with approximately 10 million parameters  (Lialin et al., 2023). Training is conducted on the C4 dataset (Raffel et al., 2020).
Baselines
We compare the following methods:
  • Standard training: this baseline utilize SGD optimizer with a learning rate of $5\times e^{-4}$, and apply a linear learning rate scheduler with a warm-up period comprising 10% of the total training steps (10,000). ]
  • $\mathtt{GradNorm}$-augmented training: apply $\mathtt{GradNorm}$ on noisy gradients before SGD update. We use the same hyperparameters as the SGD baseline.
Methodology
For each method, we measure the change of gradient statistics due to mini-batch sampling noise. For each training step, after the parameter update, we additionally sample 16 different mini-batches (each with batch size 64), and calculate the mean and standard deviation of the corresponding gradients on each of those batches. Then, we estimate the Kullback-Leibler divergence between the current gradient distribution (raw gradient and with $\mathtt{GradNorm}$) relative to the initial distribution at $t=0$. This allows us to quantify the temporal evolution of gradient distributions.
Results
Figure 4 illustrates the KL divergence of gradient distributions over training steps for both standard training and $\mathtt{GradNorm}$-augmented training, relative to the initial distribution at $t=0$. Apart from minor spikes in the beginning of training, $\mathtt{GradNorm}$ consistently and effectively stabilizes the gradient noise distribution. Furthermore, as a by-product, the loss curves demonstrate that models trained with $\mathtt{GradNorm}$ performs indeed improves upon standard training.
Ablation
One question that we might ask is this: is the performance boost of $\mathtt{GradNorm}$ simply due to that fact that $\mathtt{GradNorm}$ might be increasing the effective learning rate. We conduct detailed ablation experiments in Section 5.6, in a larger scale setting, which show this is not the case. These findings validate the hypothesis that $\mathtt{GradNorm}$ serves as an effective mechanism for mitigating the adverse effects of mini-batch noise on gradient distributions, thereby enhancing the stability and performance of the training process.
5.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
We aim to evaluate the pure optimization capability of $\mathtt{GradWhitening}$ processed gradient descent on classic optimization problems. We consider the following three settings:
(a) Quadratic optimization, well-conditioned
(b) Quadratic optimization, ill-conditioned
(c) Rastrigin function optimization
(d) Ablation on initialization
Figure 5: Comparison of convergence rate of different methods on quadratic and non-convex optimization problems. (a): 2500-dimensional quadratic optimization with well-conditioned ${\mathbf{H}}$. (b): 2500-dimensional quadratic optimization with ill-conditioned ${\mathbf{H}}$ (c): 2500-dimensional Rastrigin function optimization. (d): 2500-dimensional Rastrigin function optimization, but forcing all methods to use the same orthogonal initial location.
  • High-dimensional quadratic optimization. We construct quadratic optimization problem of the form in Equation 23, with ${\mathbf{W}}\in\mathbb{R}^{50\times 50}$.
  • Ill-conditioned quadratic optimization. We construct a quadratic optimization problem with the same dimensionality and formulation as above, but this time we deliberately choose an ill-conditioned ${\mathbf{H}}$.
  • Non-convex optimization with multiple local optimas. We consider the following multi-variate Rastrigin function:
    $f({\mathbf{W}})=m^{2}{\mathbf{A}}+\frac{1}{2}\text{Tr}[{\mathbf{W}}^{\top}{\mathbf{W}}]-{\mathbf{A}}\sum_{ij}\text{cos}(2\pi W_{ij})$
    where ${\mathbf{W}}$ is $m\times m$ matrix. This function has $10^{m^{2}}$ possible local optima. This is a complicated stress test for optimization algorithms, where many methods are expected to fail at different regimes. We set $m=50$, and compare the same methods as in the previous experiment.
Baselines
On all three settings (quadratic, quadratic ill-conditioned, and Rastrigin) we compare 5 different methods, including 3 baselines and 2 $\mathtt{GradWhitening}$ based variants. For baselines, we have: gradient descent (with theoretical optimal learning rate in Theorem 4); Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$ with hand-tuned learning rate); and Newton’s method with tuned learning rate. For 2 $\mathtt{GradWhitening}$-based methods, we consider: $\mathtt{GradWhitening}$-augmented GD with orthogonal initialization (i.e., ${{\mathbf{W}}^{(0)}}^{\top}{\mathbf{W}}^{(0)}={\mathbf{I}}$); and without orthogonal initialization. This is to verify whether the claims of Theorem 2 and Theorem 3 is true (i.e., $\mathtt{GradWhitening}$-processed GD behaves like Newton’s method under orthogonal initialization). All methods are initialized with the same initial condition (apart from $\mathtt{GradWhitening}$ with orthogonal initialization, which will project the original initialization onto the orthogonal version).
Results
Results are shown in Figure 5, (a) - (c). Main findings are summarized as follows:
  • On quadratic optimization problems (Figure 5 (a) and (b)), $\mathtt{GradWhitening}$ with orthogonal initialization and Newton’s method converges to optimal solution with 1 step of optimization, confirming our theoretical prediction in Theorem 2 and Theorem 3.
  • On quadratic optimization problems (Figure 5 (a)), when the problem is well-conditioned, standard GD outperforms both Adam and $\mathtt{GradWhitening}$ (non-orthogonal initialization). However, when the problem becomes ill conditioned (Figure 5 (b)), $\mathtt{GradWhitening}$ (non-orthogonal initialization) outperforms GD by a large margin; while GD suffers from convergence.
  • In all three settings (Figure 5, (a) - (c)), $\mathtt{GradWhitening}$ with non-orthogonal initialization consistently outperforms Adam, confirming our theoretical results in Proposition 3.
  • On Rastrigin function optimization (Figure 5 (c)), $\mathtt{GradWhitening}$ performs competitively to Newton’s method, with or without orthogonal initialization.
Ablation
Furthermore, we additionally conduct ablation study on the effect of initialization locations, as shown in fig. 5 (d). In this case, we force all methods to use the same orthogonalized initialization location for optimization. As shown by the result, $\mathtt{GradWhitening}$ still consistently outperforms all baselines, eliminating the potential hypothesis that orthogonalized initialization gave $\mathtt{GradWhitening}$ an unfair advantage.
5.3 SWAN Performance on LLM Pre-training Tasks
Setup
Finally, we present results on the task of pre-training large language models. Below we consider LLaMa architecture used in Zhao et al. (2024a) with RMSNorm and SwiGLU activations. We consider sizes of 60M, 130M, 350M and 1.3B parameters on C4 dataset Raffel et al. (2020), trained with an effective batch size of 130K tokens (512 batchsize and 256 context length). SWAN is applied to all linear modules (in both attention and MLP layers). We run all experiments with BF16 format by default.
Table 1: Comparison with Adam and its memory-efficient low-rank variants on pre-training various sizes of LLaMA models on C4 dataset. Validation perplexity is reported, along with a memory estimate of the total of parameters and optimizer states based on BF16 format. Since we used train scripts and configs based on the code opensourced by Zhao et al. (2024a), all baseline results are taken from Zhao et al. (2024a). The exceptions are the results for Adam and Galore, as we cannot exactly reproduce the results reported in Zhao et al. (2024a); hence we reported results from our own runs instead.
60M 130M 350M 1.3 B
SWAN 32.28 (0.12G) 24.13 (0.25G) 18.22 (0.68G) 15.13 (2.60G)
Adam 33.02 (0.36G) 24.44 (0.76G) 20.11 (2.06G) 16.44 (7.80G)
Galore 33.09 (0.24G) 24.67 (0.52G) 19.74 (1.22G) 15.89 (4.38G)
Low-Rank 78.18 (0.26G) 45.51 (0.54G) 37.41 (1.08G) 142.53 (3.57G)
LoRA 34.99 (0.36G) 33.92 (0.80G) 25.58 (1.76G) 19.21 (6.17G)
ReLoRA 37.04 (0.36G) 29.37 (0.80G) 29.08 (1.76G) 18.33 (6.17G)
SWAN speed up vs Adam 1.2 X 1.23 X > 2X > 2X
$r$ of low-rank methods 128 256 256 512
Training Tokens 1.1B 2.2B 6.4B 13.1B
(a) 60M
(b) 130M
(c) 350M
(d) 1.3B
Figure 6: Comparison of convergence rate of different methods on LLM pretraining tasks.
Baselines
Since the main goal of SWAN is to explore the feasibility of EMA-free adaptive optimizers for LLMs, we mainly compare to Adam and its memory efficient variants. We use the exact evaluation settings of Zhao et al. (2024a); along with Adam, these include:
  • Galore Zhao et al. (2024a): the state-of-the-art memory efficient variant of Adam that based on low-rank projections of gradients.
  • Low-Rank (Kamalakara et al., 2022): Low-rank reparameterization of the weights with the trainable factorization ${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$
  • LoRA: LoRA Hu et al. (2021) applied for pre-training in the exact setting as used in Zhao et al. (2024a).
  • ReLoRA Lialin et al. (2023): a full-rank learning extension of Lora with parameter merging.
Hyperparameters
For all baselines, we directly apply the same hyperparameter configs (including model configs, optimizers, schedulers) open sourced by Zhao et al. (2024a)  [^4] This does not include Adam, since the authors of Zhao et al. (2024a) did not disclose their Adam configs. However, we used the same learning rate sweep procedure as described in their paper. This result in an optimal learning rate of 0.001 for 60M, 130M and 350M models, and 0.0007 for 1.3 B models. These are also consistent with the optimal learning rates found in some other works, such as Pagliardini et al. (2024). . This includes method-specific fine-tuned optimal learning rates obtained via searching on fixed grid Zhao et al. (2024a). For our SWAN method, we simply use a hand-picked default hyperparameter setting out-of-the-box, without fine tuning. These include a un-tuned, default learning rate of 0.001 across all models, and the 10 steps of Newton-Schultz iteration. Notably, thanks to the strong robustness offered by $\mathtt{GradWhitening}$ to ill-conditioned problems, SWAN can safely start training without any learning rate warm-up. This is not the case for other baselines such as Adam, in which they failed to converge without long warm-ups in our ablation experiments. Apart from the warm-up hyperparameter, all the rest of the parameters in learning rate scheduling used by SWAN is the same as all other baselines.
Remark on fair comparison
As empirically shown in Kaddour et al. (2024), it is essential to use the same training budget when comparing efficient optimizers; unfortunately, this fact was not widely respected in some recent work of efficient algorithms. Kaddour et al. (2024) showed that certain state-of-the-art efficient optimizers has shown no gain over Adam under the same training budget. Therefore, in our experiments above, we exactly follow the rule of equal training budget (in our case, defined in terms number of tokens), and forces all methods to have learning rate fully decayed to the same targe learning rate at the end of optimization. For a even fairer comparison, we left the learning rates of SWAN untuned.
Results
Results are shown in Table 1 and Figure 6. SWAN outperforms all baselines in both validation PPLs, as well as estimated memory consumption (SWAN theoretically should consume the same memory as SGD). Furthermore, to our surprise, on both 350M and 1.3B models, SWAN achieved $>$ 2X speed up against Adam in terms of number of steps (hence number of tokens), see Figure 1 and Figure 7 (a). More in-depth discussion of speed-up will be provided in Section 5.4.
5.4 Is the Improvement Multiplicative or Additive?
One natural question when evaluating speedup factors is whether our method provides a multiplicative or additive speedup over Adam. Specifically, multiplicative speedup indicates an optimizer’s performance advantage over a baseline increases proportionally with time, maintaining a consistent speed-up ratio. This is the ideal scenario we would like to achieve. In contrast, additive speed-up implies only a fixed, constant advantage in steps over baseline, which is less desired. To investigate this, we utilize two plots: the speed-up ratio comparison and the perplexity comparison (Figure 7), across different model sizes. Before understanding those plots, we first introduce the following concepts:
(a) Speedup factor of SWAN vs Adam training steps
(b) Ours PPL vs Adam PPL scatter plots
Figure 7: Comparative analysis of SWAN and Adam optimizers: speedup ratios and perplexity metrics across various model sizes. (a) shows how SWAN reduces the number of training steps needed to achieve the same evaluation perplexity as Adam for models ranging from 60M to 1.3B parameters. A speedup ratio greater than one indicates that SWAN reaches target PPL values faster than Adam. (b) presents a direct comparison of perplexity scores between SWAN and Adam. In both plots, we also provide counterfactual additive curves (dashed lines) modeling baselines corresponding to constant step advantages. Together, these plots highlight the nature of SWAN’s speedup over Adam across different model scales.
Speedup ratio definition
The speedup ratio $R(P)$ is defined as the ratio of the number of training steps Adam requires to reach a specific evaluation perplexity (PPL) to the number of steps SWAN requires to achieve the same PPL. Mathematically, for a given PPL threshold $P$, the speedup ratio is expressed as:
$R(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{SWAN{}}}(P)}$ (6)
where $S_{\text{Adam}}(P)$ and $S_{\text{SWAN{}}}(P)$ denote the training steps taken by Adam and SWAN, respectively, to reach perplexity $P$.
Counterfactual additive curve estimation
To determine if the speed-up is additive, we estimate a counterfactual additive curve by assuming a constant step advantage $\Delta$ of SWAN over Adam. This is calculated using the first 10% - 20% (depending on the number of maximum steps) of the training data:
$\Delta=\frac{1}{N}\sum_{i=1}^{N}\left(S_{\text{Adam}}(P_{i})-S_{\text{SWAN{}}}(P_{i})\right)$ (7)
where $N$ is the number of PPL thresholds within the initial 10% - 20% of training steps.
For the speedup ratio plot:
$R_{\text{additive}}(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{Adam}}(P)-\Delta}$ (8)
For the perplexity comparison plot:
$\text{PPL}_{\text{additive}}(S)=\text{PPL}_{\text{Adam}}(S+\Delta)$ (9)
This represents the expected perplexity of SWAN if it only consistently outperforms Adam by $\Delta$ steps.
Results
Results are shown in Figure 7. In both subplots, the actual performance of SWAN is compared against the counterfactual additive curves. If the actual plot lies significantly above the counterfactual additive curve, it indicates a stronger tendency towards multiplicative speedup, meaning SWAN’s performance gains are more than just a fixed step advantage. Analyzing models of sizes 60M, 130M, 350M, and 1.3B parameters, we observe that:
  • For smaller models (60M and 130M parameters), the actual speedup ratios closely follow the counterfactual additive curves, suggesting that the speedup of SWAN over Adam is primarily additive.
  • For larger models (350M and 1.3B parameters), the actual plots significantly exceed the counterfactual additive curves, and has a clear sign of accelerated growth. This pronounced deviation indicates a multiplicative speedup, demonstrating that SWAN offers increasingly substantial efficiency gains as model complexity scales.
5.5 Memory Efficiency and Throughput Analysis
This section, we benchmark the memory footprint of SWAN as well as its throughput.
Memory footprint
We benchmark the memory footprint of SWAN, Adam, and Galore in a practical scenario.. Unlike Zhao et al. (2024a) where memory is benchmarked under a layer-wise training strategy, we directly measure end-to-end memory under full-model training, with batch size = 1 sequence, across different model sizes (1.3B, 7B, and 13B). Results are shown in Figure 1 (c), SWAN achieves near-theoretical limit of optimizer memory efficiency of full-parameter training (assuming no quantization). This means that SWAN can reproduce the memory footprint of vanilla SGD, reaching $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states. This result again highlights the advantage of the stateless design.
Effective throughput
To measure the throughput of SWAN, we consider training a 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens. Here We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the efficiency of optimizer relative to Adam. The purpose of the first metric (absolute throughput) is to measure whether SWAN (especially the use of N-S iterations in $\mathtt{GradWhitening}$) create a significant drop in training throughput; and the second metric (effective throughput) takes into account the fact that different optimizers utilizes training tokens at different efficiency level. Results are shown in Figure 1 (d), SWAN closely reproduces the absolute throughput of Adam, implying that unlike full SVD used in (Zhao et al., 2024a), the computational overhead of $\mathtt{GradWhitening}$ is minimal. As a result, SWAN has doubled the effective throughput vs Adam, due to its high token utilization efficiency as shown in Section 5.3. All results are obtained using $\mathtt{GradWhitening}$ in BF16 format.
5.6 Ablation Studies
How does $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ contribute to the performance boost?
We consider the following 6 ablation settings: SWAN(full), SWAN ($\mathtt{GradNorm}$ only), SWAN ($\mathtt{GradWhitening}$ only), Adam (full), Adam (momentum only), and Adam (second moment only). Results are shown in Figure 8 (a). It suggests that both $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ makes significant contribution towards the final performance of SWAN; and removing any of them results in suboptimal performance. This is similar to how the first and second moment EMAs of Adam jointly contributes to the final performance.
Figure 8: Ablation studies on 130M model. (a) Ablation on the contribution of each components in SWAN and Adam. (b) Ablation on removing $\mathtt{GradNorm}$ and compensate with larger learning rates. (c) Ablation on the effect of learning rate warm-ups.
Is it true that SWAN works only because its $\mathtt{GradNorm}$ component might have increased the effective learning rates?
To answer this question, we remove the $\mathtt{GradNorm}$ component of SWAN, and perform learning rate sweep for it. We start from the default base learning rate used by full SWAN, and then scale it with multipliers ranging from $1$ to $10^{3}$. Results are shown in Figure 8 (b), where the final validation ppl of SWAN without $\mathtt{GradNorm}$ under different learning rate multipliers are shown in blue line; and the performance of the full SWAN under default learning rate is shown as a dashed red reference line. The result suggests that, although increasing the effective learning rates might improve the performance of SWAN without $\mathtt{GradNorm}$, there is still a significant cap between between SWAN without $\mathtt{GradNorm}$ and the full SWAN. Together with results in Section 5.1, we can conclude that the gradient noise stabilization of $\mathtt{GradNorm}$ is indeed essential, and it cannot be trivialized as larger effective learning rates.
How does warm-up affect the performance?
In Section 5.3, we showed that SWAN does not need any learning rate warm-ups and can train with large learning rates (the default 0.001 learning rate used by all methods is considerably large in LLMs settings). Here we perform ablation and compare Adam and SWAN both under warm-ups and not warm-ups. As suggested by the results in Figure 8 (c), we found that for SWAN, training without learning rate warm-ups indeed gives better performance. Meanwhile, SWAN still outperforms Adam (with warm-up) under the same warm-up schedule of Adam. On the contrary, Adam only works with proper learning rate warmup. If we remove warmups, the performance of Adam becomes drastically worse. This confirms our claim that SWAN is more robust to the local geometries of the optimization problems and can work without (and with) learning rate warm ups.
6 Related works
Low-rank memory efficient methods
Low-rank memory-efficient optimizers have become pivotal in training large language models (LLMs) by reducing memory consumption while maintaining model performance. A foundational approach in this domain is LORA introduced by Hu et al. (2021), which fine-tunes pre-trained models using low-rank adaptors for each layer, thereby minimizing the memory footprint. This leads a full line of low-rank adaptation research, but they are mostly focused on post-training adaptations. More recently, ReLoRA Lialin et al. (2023),FLORA Hao et al. (2024) and Galore Zhao et al. (2024a) further push the boundaries of memory-efficient pre-training for LLMs, by introducing the idea of using low-rank projection of gradients to achieve full-rank learning. They have demonstrated significant memory savings on optimizer states, with relatively low impact on model performance. However, these methods often require 30–50 % of the full Adam’s states and may underperform in terms of convergence speed and final model quality (Zhao et al., 2024a). Moreover, it is still an open question on the design choices of these methods, for example: how and why to choose low-rank subspace and projectors? Are these projections properly used, given the non-exchangeability nature between low-rank projection and Adam EMA operator? Why spikes in loss curves are observed during subspace are shifted/mixed (Chen et al., 2024)? These unaddressed questions might make the behavior of low-rank gradient optimizers unpredictable in practice.
Non-adaptive methods with partial EMA removal.
Lion (Chen et al., 2023) and Signum (Bernstein et al., 2018) have introduced effective ways to simplify optimization by partially removing the exponential moving average (EMA) of the second moment. These methods reduce memory and computational overhead while retaining the benefits of momentum through the first moment EMA, offering a practical trade-off between simplicity and performance. However, their reliance on momentum inherently limits their extension to further remove both first and second moment EMAs are removed, as both would reduce to SignSGD (Bernstein et al., 2018), which lacks the adaptive learning rate capability of Adam. This might explain why sign-based methods have limited success in LLM pre-training tasks (Chen et al., 2023; Liu et al., 2023). On the contrary, our approach completely removes all EMA terms while maintaining the adaptive learning rate nature, which is the key to the success of modern adaptive optimizers.
7 Conclusion
The most important insight that we conclude for the paper is, preprocessing SGD might already be enough for LLM training tasks. This presents an elegant solution to address the memory-performance dilemma of training LLMs using adaptive optimizers like Adam. The other insight comes from the fact that the key components of the SWAN optimizer were designed using key insights from LLM dynamics under a simplified setting. In that sense, this paper serves as a proof-of-concept on the value of studying the dynamics of LLM architectures when designing optimizers. We hope both research insights can inspire more research alone this line to further advance this field.
Appendix A Desired properties of adaptive optimizers
There is a rich literature on understanding adaptive methods’ inner workings and unreasonable effectiveness. Using Adam as an example, we first summarize from the literature below the key desired properties of stateful adaptive optimizers that contribute to their empirical success: gradient smoothing, gradient invariance, and gradient whitening. Then we discuss how these understandings will leads to the design of stateless adaptive optimizers.
Gradient Smoothing.
Under the stochastic optimization setting, mini-batch sampling introduces heterogeneous distribution shift on the gradient distribution: ${\mathbf{G}}^{(t)}=\mathbb{E}[{\mathbf{G}}^{(t)}]+{\bm{\varepsilon}}^{(t)}$, where ${\bm{\varepsilon}}^{(t)}$ is time-heterogeneous noise induced by mini-batch sampling. While ${\bm{\varepsilon}}^{(t)}$ helps SGD escapes local optima (Jastrzębski et al., 2017; Zhu et al., 2018), the covariate shift of ${\bm{\varepsilon}}^{(t)}$ over time also present challenges to learning as the model needs to adjust and compensate for this shift, especially under the emergence of heavy tailed gradient distributions (Zhang et al., 2020)  [^5] Such shift cannot be removed by forward covariate-shift reduction architectures such Layer Norm, as it is only invariant to global scaling and re-centering, such as ${\mathbf{W}}^{(t)}=\delta{\mathbf{W}}^{(t)}+\bm{\gamma}\mathbf{1}^{\top}$ for some scalar $\delta$ and incoming vector shift $\bm{\gamma}$ (Ba et al., 2016).. Following this viewpoint, it has been proven that momentum reduces the influence of noises for SGD (Cutkosky & Mehta, 2020; Crawshaw et al., 2022). Therefore we hypothesis that the first moment estimate ${\bm{m}}^{(t)}$ of Adam also effectively stabilizes gradient distribution and reduces effect of ${\bm{\varepsilon}}^{(t)}$. This smoothing stabilizes the variance caused by noisy stochastic gradients across time.
Gradient Invariance.
More recently it has also been identified (Kunstner et al., 2023; 2024) that the major factor contributing to the performance gap between SGD and Adam might lie in Adam’s Sign-descent-like nature (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023). Intuitively, Adam without bias correction under $\beta_{1}=0$ and $\beta_{2}=0$ is equivalent to signed gradient descent ($\Delta{\mathbf{W}}=\text{sign}({\mathbf{G}})$). Indeed, the performance of Adam can be closely reproduced (Kunstner et al., 2023; Crawshaw et al., 2022) or even surpassed (Chen et al., 2023) by variants of signed descent with momentum. Apart from sign-based methods, evidence on performance boost using gradient clipping/normalization was also discussed in the context of understanding Adam (Zhang et al., 2020). Therefore, we hypothesize that one of the key properties of Adam is that it offers invariance over certain transformations on gradients. Particularly, the original Adam is invariant to diagonal rescaling of the gradients (Kingma & Ba, 2015); the signed gradient method is invariant to any scaling that preserves the sign of gradients; and the clipped SGD variant is invariant to extreme gradient magnitude spikes.
Gradient Whitening.
Finally, we argue that the empirical success of adaptive methods also lies in that they model the curvature by first-order information. This is realized by the second moment estimate ${\bm{\nu}}^{(t)}$, which approximates the diagonal of the Fisher information matrix (Kingma & Ba, 2015; Hwang, 2024); helping to counteract local curvatures of the problem. Specifically, Adam computes a trailing estimation of the diagonal coefficients of the Fisher matrix ${\mathbf{F}}=\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ by tracking $\hat{{\mathbf{F}}}=\text{diag}({\mathbf{F}})=\text{diag}[\mathbb{E}[{\bm{g}}^{2}]]$, where ${\bm{g}}=\mathtt{vec}({\mathbf{G}})$ is the vectorized gradient. Interestingly, instead of preconditioning the first moment as $\hat{{\mathbf{F}}}^{-1}\mathtt{vec}({\bm{m}})$, Adam uses a whitening-like preconditioned update $\hat{{\mathbf{F}}}^{-\frac{1}{2}}\mathtt{vec}({\bm{m}})$, suggesting an element-wise approximate whitening of the gradient. It has been shown that such element-wise whitening leads to diagonal approximation to inverse Hessian $\hat{{\mathbf{F}}}^{-\frac{1}{2}}\approx\text{diag}({\mathbf{H}}^{-1})$ (Molybog et al., 2023). Recent empirical studies show that Adam biases optimization trajectories towards regions where the condition number of Hessian is low (Jiang et al., 2024). Therefore, we hypothesize that Adam approximately whitens the gradients element-wise, leading to well-conditioned regions.
Appendix B Theoretical Analysis
B.1 Analyzing the GradNorm: A Transformer Learning Dynamics Perspective
We take the transformer LLM architecture as an example, and demonstrate that $\mathtt{GradNorm}$ effectively removes the time-variant components of gradient noises. For our analysis, we assume the following simplified transformer block (STB) architecture recently proposed in Tian et al. (2023), which contains the major building blocks of transformers and relies on relatively less restrictive assumptions and simplifications, making them suitable for theoretical analysis for transformer dynamics. It is defined as follows:
Definition 1 (Simplified Transformer Block (STB)).
Given the input activation ${\bm{x}}\in\mathbb{R}^{M_{C}\times 1}$, query token index $q$, context embedding matrix ${\mathbf{U}}_{C}\in\mathbb{R}^{d\times M_{C}}$, and the query embedding ${\bm{u}}_{q}\in\mathbb{R}^{d\times 1}$, the simplified transformer block computes the output ${\bm{h}}\in\mathbb{R}^{n\times 1}$ as follows:
${\bm{h}}=\phi\left({\mathbf{W}}^{\top}\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)\right),$
where the attention logits ${\bm{z}}_{q}\in\mathbb{R}^{M_{C}\times 1}$ are given by
$z_{ql}={\bm{u}}_{q}^{\top}{\mathbf{W}}_{Q}^{\top}{\mathbf{W}}_{K}u_{l},$
with ${\mathbf{W}}_{Q},{\mathbf{W}}_{K}\in\mathbb{R}^{d\times d}$ being weight matrices for the queries and keys, respectively, ${\mathbf{W}}\in\mathbb{R}^{d\times n}$ is the weight matrix for the feedforward network, and $\phi$ is a nonlinearity function such as ReLU.
Given a STB, we consider the following standard mini-batch learning dynamics. Define the conditional expectation $\mathbb{E}_{q=m}[\cdot]:=\mathbb{E}[\cdot|q=m]$. Consider the dynamics of the weight matrix $W$ and the attention logits $z_{q}$, if we train the model with a batch of inputs that always end up with query $q[i]=m$. The weight update for $W$ and $z_{q}$ are given by the following noisy updates:
$\dot{{\mathbf{W}}}^{(t)}=\mathbb{E}_{q=m}\left[{\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}\right]+{\bm{\varepsilon}}^{(t)},\quad\dot{{\bm{z}}}_{m}^{(t)}=\mathbb{E}_{q=m}\left[\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}\right]+{\bm{\varsigma}}^{(t)},$ (10)
Where ${\bm{f}}^{(t)}=\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q}^{(t)})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)$, $({\bm{h}}^{(t)})^{\prime}=\phi^{\prime}(({\mathbf{W}}^{(t)})^{\top}{\bm{f}}^{(t)})$ is the derivative of the current activation, ${\mathbf{G}}_{{\bm{h}}}^{(t)}=\nabla_{{\bm{h}}^{(t)}}\mathcal{L}$ is the gradient of the loss function $\mathcal{L}$ with respect to the hidden activation ${\bm{h}}^{(t)}$, and ${\bm{g}}_{{\bm{f}}^{(t)}}^{(t)}=\sum_{k}{\bm{g}}_{{\bm{h}}_{k}^{(t)}}^{(t)}({\bm{h}}^{(t)}_{k})^{\prime}{\bm{w}}_{k}^{(t)}$ is the sum of the gradients with respect to the attention logits. Here, ${\bm{w}}_{k}^{(t)}$ is the $k$-th column of ${\mathbf{W}}^{(t)}$, ${\bm{g}}_{h_{k}^{(t)}}^{(t)}[i]$ be the backpropagated gradient sent to node $k$ at sample $i$. Finally, ${\bm{\varepsilon}}$ and ${\bm{\varsigma}}$ are random noises that are induced by the minibatch sampling.
Remark
Given STB under the dynamics of Equation 10, the distribution of noises ${\bm{\varepsilon}}^{(t)}$ and ${\bm{\varsigma}}^{(t)}$ generally vary across time, since both the integrand within expectation, ${\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}$ and $\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}$ vary across time. However, the following result on $\mathtt{GradNorm}$ suggests that this time variability of ${\bm{\varepsilon}}^{(t)}$ and ${\bm{\varsigma}}^{(t)}$ are completely removed by $\mathtt{GradNorm}$:
Theorem 1 ($\mathtt{GradNorm}$ stabilizes gradient distribution across time for STB).
Consider the STB (Definition 1). Assuming we inherit the assumptions in Theorem 1 of Tian et al. (2023), as described in Appendix B. Then consider ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, the composition of the MLP project-up matrix and the embedding matrix as a whole. Then, its standardized stochastic gradients $\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}}(\top{\bm{x}}^{(t)})}{{\mathbf{U}}_{C}^{\mathbf{W}}})$ satisfy:
$\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{1})}]=\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{2})}]\quad\text{for all }t_{1},t_{2},\text{and}\,i.$
In other words, the covariance structure of $\tilde{{{\mathbf{G}}}}$ is identical across all time steps $t$, achieving distributional stability across time. The same relationship also holds for the gradient of attention score $\tilde{{\mathbf{G}}}_{{\bm{z}}_{q}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}(\top{\bm{x}}^{(t)})}{\partial{\bm{z}}_{q}})$.
Theorem 1 suggests that $\mathtt{GradNorm}$ makes use of the inductive bias induced by transformer architectures and directly removes the time-variant components in gradient covariance structures caused by mini-batch sampling. In practice, $\mathtt{GradNorm}$ makes the following approximation to Theorem 1:For MLP layer $\mathtt{GradNorm}$ is applied on ${\mathbf{W}}$ instead of ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$ as a whole, since such grouping does not exist in practical transformer architectures. However, it still works reasonably well in practice (Section 5.1, Section 5.6).
B.2 Analyzing GradWhitening Pt. I: A Transformer Learning Dynamics Perspective
Figure 9: Normalized Hessian Blocks of size $M_{C}\times M_{C}$ along the diagonal direction of the Hessian, obtained from numerically solving the STB ODE (with $n=12,M_{C}=10$) (1) given by the full-batch dynamics (i.e., removing noise in Equation 10). During all training steps, we analytically track the evolution of Hessian. As predicted by Proposition 1, we see very similar structures across the diagonal blocks of the Hessian.
In this section, we take the perspective of transformer learning dynamics, and show that how $\mathtt{GradWhitening}$ can be naturally derived from the simplified transformer model defined in Definition 1.
Let us first consider the dynamics of the Hessian matrix of STB. We will show that: i), it turns out that at equilibirum, the Hessian is highly structured; ii), under some conditions, $\mathtt{GradWhitening}$ recovers Newton’s method-like behaviors under such Hessian structures. Below we first present the following result on the Hessian structure of the simplified transformer block at equilibrium:
Proposition 1 (Shared structures in the block-diagonal of Hessians at transformer equilibrium).
Consider a STB (1), trained by full-batch dynamics (i.e., removing noise in Equation 10). Assume we inherit all the assumptions from Theorem 1 of Tian et al. (2023). Consider the following Hessian matrix:
${\mathbf{H}}({\mathbf{W}})_{sk,s^{\prime}k^{\prime}}=\frac{\partial\mathcal{L}}{\partial{w_{sk}}\partial{w_{s^{\prime}k^{\prime}}}}$
where $1\leq s,s^{\prime}\leq m$, $1\leq k,k^{\prime}\leq n$. Then, as $t\xrightarrow[]{}\infty$, we have the following shared Hessian structure along the diagonal blocks:
$\frac{{\mathbf{H}}_{sk,s^{\prime}k}}{\sum_{s,s^{\prime}}{\mathbf{H}}_{sk,s^{\prime}k}}\xrightarrow[]{}\frac{{\mathbf{H}}_{sk^{\prime},s^{\prime}k^{\prime}}}{\sum_{s,s^{\prime}}{\mathbf{H}}_{sk^{\prime},s^{\prime}k^{\prime}}},\quad\forall 1\leq s,s^{\prime}\leq d,1\leq k,k^{\prime}\leq n$ (11)
Proposition 1 shows that, under a simplified setting of transformer, because of the winner-takes-all behavior in the attention model, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares the identical structure, after normalization. This result is verified in our numerical experiment as shown in Figure 9.
Given by the inductive bias that equality across diagonal blocks of Hessians at transformer learning equilibrium, in the next theorem we show that $\mathtt{GradWhitening}$ can be derived as a second-order method under this structural assumption
Theorem 2 ($\mathtt{GradWhitening}$ as structural approximation of Hessian).
Consider the optimization problem given by $\min_{W\in\mathbb{R}^{m\times n}}f({\mathbf{W}})$ where ${\mathbf{W}}\in\mathbb{R}^{m\times n}$, with $m<n$. Assume:
  • The local Hessian $H$ of $f$ has shared block-diagonal, such that ${\mathbf{H}}=\text{diag}(\tilde{{\mathbf{H}}},\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}})={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
  • ${\mathbf{W}}^{(t)}$ is initialized as a orthogonal matrix.
Then (under 2nd order approximation) $\mathtt{GradWhitening}$’s modified gradient descent (Equation 13) on $f$ recovers the update of Newton’s method, without explicitly estimate second-order quantities.
To summarize, $\mathtt{GradWhitening}$ recovers a Newton’s method-like behavior under the highly structured Hessian regime (i.e., there exists $\tilde{{\mathbf{H}}}$ such that ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$), which is partially satisfied in transformer equilibrium dynamics (see remark below). Compared to directly applying accelerated second-order methods such as the Gauss-Newton method ($\mathcal{O}(Bm^{2}n^{2}+m^{3}n^{3})$, $B$ being batch size) or conjugate gradient method ($\mathcal{O}(BTmn)$ complexity, $T=$ 10 $\sim$ 100, with worst case complexity $\mathcal{O}(Bm^{2}n^{2})$), $\mathtt{GradWhitening}$ should be much faster ($\mathcal{O}(mn\min\{m,n\})$), especially under the Newton-Schulz iteration.
Remark: limitations
Proposition 1 deviates from the strict assumptions of Theorem 2 in that: Proposition 1 only asserts that diagonal direction of Hessian shares the identical structure, but does not assert that the off-diagonal blocks are zero, without further assumptions on $g_{h_{k}}$. However, given that it is standard practice to ignore off-diagonal elements in many practical optimizers including Adam, we hereby argue that ignoring off-diagonal blocks is already a better approximation.
B.3 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
In this section, we present main results regarding the convergence rate of the $\mathtt{GradWhitening}$ method, understand its implications, and compare it with the lower bounds of GD and Adam.
First, for simplicity, we focus on the following quadratic problem:
$\mathcal{L}({\mathbf{W}})=\frac{1}{2}\text{Tr}({\mathbf{W}}^{\top}{\mathbf{H}}{\mathbf{W}})-\text{Tr}({\mathbf{C}}^{\top}{\mathbf{W}}),$ (12)
where ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ is the parameter matrix, ${\mathbf{H}}\in\mathbb{R}^{m\times m}$ is a positive definite matrix, and ${\mathbf{C}}\in\mathbb{R}^{m\times n}$ is a constant matrix.
For simplicity and without loss of generality, we assume ${\mathbf{C}}=0$. This is because minimizing $\mathcal{L}({\mathbf{W}})=\frac{1}{2}\text{Tr}({\mathbf{W}}^{\top}{\mathbf{H}}{\mathbf{W}})-\text{Tr}({\mathbf{C}}^{\top}{\mathbf{W}})$ is equivalent to minimizing $\mathcal{L}({\mathbf{W}})=\frac{1}{2}\text{Tr}[({\mathbf{W}}-{\mathbf{W}}^{*})^{\top}{\mathbf{H}}({\mathbf{W}}-{\mathbf{W}}^{*})]$, where ${\mathbf{W}}^{*}={\mathbf{H}}^{-1}{\mathbf{C}}$. By defining ${\mathbf{Z}}={\mathbf{W}}-{\mathbf{W}}^{*}$, the problem reduces to minimizing $\mathcal{L}({\mathbf{Z}})=\frac{1}{2}\text{Tr}({\mathbf{Z}}^{\top}{\mathbf{H}}{\mathbf{Z}})$.
Remark
Most results in this note can be easily extended to any loss function that are either i) strongly convex; or ii) has twice differentiable functions and Lipschitz continuous Hessian, by considering their the second order approximation around ${\mathbf{W}}^{*}$.
Next, to understand the effect of $\mathtt{GradWhitening}$, we will examine the gradient flow dynamics induced by $\mathtt{GradWhitening}$. Consider the $\mathtt{GradWhitening}$-modified gradient descent:
$\Delta{\mathbf{W}}^{(t)}=-\eta\mathtt{GradWhitening}({\mathbf{G}}^{(t)})$ (13)
its exact convergence rate is given by the result as below:
Theorem 3 (Contraction factor of $\mathtt{GradWhitening}$).
Consider the quadratic loss function Equation 12. Assume the initialization distribution of ${\mathbf{W}}^{0}$ assigns zero probability to any set of zero Lebesgue measure in $\mathbb{R}^{m\times n}$. Let our update rule be:
${\mathbf{W}}_{\text{whitened}}^{(t+1)}={\mathbf{W}}_{\text{whitened}}^{(t)}-\eta\mathtt{GradWhitening}({\mathbf{G}}^{(t)})$
where the learning rate is $\eta$. Then, with probability 1, we have:
  • The optimal dynamic learning rate to achieve the fastest convergence is given by
    ${\eta^{(t)}}^{*}=\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}}{\text{Tr}[{\mathbf{H}}]}.$ (14)
    where $\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}$ denotes the Schatten $p$-norm with $p=1$ (i.e., sum of singular values).
  • Under ${\eta^{(t)}}^{*}$, the contraction factor of loss function at $t$ is given by:
    $\frac{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t+1)})-\mathcal{L}^{*}}{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})-\mathcal{L}^{*}}=1-\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$ (15)
  • Furthermore, if we additionally enforce ${\mathbf{W}}^{0}\sim V^{m\times n}(\mathbb{R})$, i.e., initialized as an element in Steifel manifold. Then we have
    $\frac{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{t=1})-\mathcal{L}^{*}}{\mathcal{L}({\mathbf{W}}^{0})-\mathcal{L}^{*}}=0$ (16)
That is, $\mathtt{GradWhitening}$ solves the optimization problem (Equation 12) with 1 step iteration.
Theorem 3 has the following key implications.
Convergence rate is condition number agnositc
Unlike the convergence rates of GD and Adam presented in Zhang et al. (2024), as well as Theorem 4 and Corollary 1 in Appendix, the optimal convergence rate (15) of $\mathtt{GradWhitening}$ no longer explicitly depends on the condition number $\kappa$ of $H$. In fact, consider a lower bound $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\geq\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$, since trace of $H$ appear both in the nominator and denominator, we expect that to be more robust to ill-conditioned problems. For example, consider the specific initialization ${\mathbf{W}}_{\text{whitened}}^{(t)}=cI$, it is straightforward to show that $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\geq\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\perp\kappa$, which is completely disentangled from the condition number. Hence $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$ would not shrink as $\kappa\xrightarrow{}\infty$. See Proposition 2 for less extreme situations.
Superlinear convergence with Stiefel manifold initialization
Theorem 3 suggests that if ${\mathbf{W}}_{\text{whitened}}^{(t)}$ is initialized in the Stiefel manifold, then $\mathtt{GradWhitening}$ reaches superlinear convergence rate (= Newton’s method), while being cheaper. In fact, it is straightforward to verify that $\mathtt{GradWhitening}$ reaches optimal solution with 1 step update. This implies $\mathtt{GradWhitening}$ is theoretically the optimal optimization algorithm if ${\mathbf{W}}$ is initialized in the Stiefel manifold.
Estimation and interpretation of optimal learning rate
Compared to the optimal dynamic learning rate of gradient descent $G=\frac{G^{\top}G}{G^{\top}HG}$, the optimal learning rate ${\eta^{(t)}}^{*}$ of $\mathtt{GradWhitening}$ is much easier to compute. $\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]}{\text{Tr}[{\mathbf{H}}]}$ can be seen as balancing the average gradient magnitude against the average curvature. A higher trace of gradient (${\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}$) (strong gradients) relative to ${\mathbf{H}}$ (steep curvature) suggests a larger learning rate, promoting faster updates. Conversely, a higher trace of ${\mathbf{H}}$ would imply a smaller learning rate to ensure stable convergence in regions with high curvature.
Next, we show that the convergence speed of $\mathtt{GradWhitening}$ update is indeed robust to the condition number of local curvature.
Proposition 2 (Robustness of $\mathtt{GradWhitening}$ update convergence rate against the condition number of local Hessian).
Consider the quantity:
$Q:=\frac{\text{Tr}[{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{t}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{T}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$
Assume: i) ,${\mathbf{W}}_{\text{whitened}}^{(t)}\neq{\mathbf{W}}^{*}$; and ii) the norm of ${\mathbf{H}}$ is bounded. Then, there exist some finite positive constant $c$, such that
$Q>c$
This holds even if $\kappa\xrightarrow{}+\infty$, where $\kappa$ is the condition number of ${\mathbf{H}}$.
Below, we provide comparison between $\mathtt{GradWhitening}$ modified gradient descent and Adam. We only consider non-Stiefel initialization for $\mathtt{GradWhitening}$, since with non-Stiefel initialization $\mathtt{GradWhitening}$ is optimal according to Theorem 3. Our results below shows that, for poor conditioned problems $\mathtt{GradWhitening}$ with a properly chosen single global learning rate always outperforms Adam even with optimally tuned sub-group learning rates, in terms of convergence speed.
Proposition 3 ($\mathtt{GradWhitening}$ with single lr vs Adam with tuned group lr).
Consider the optimization problem Equation 12. Assume ${\mathbf{H}}$ is block-diagonal, i.e., ${\mathbf{H}}=\text{diag}({\mathbf{H}}_{1},{\mathbf{H}}_{2},\dots,{\mathbf{H}}_{L}),$ where each ${\mathbf{H}}_{l}\in\mathbb{R}^{m_{l}\times m_{l}}$ is a positive definite matrix for $l=1,2,\dots,L$, and $\sum_{l=1}^{L}m_{l}=m$. Assuming for $\mathtt{GradWhitening}$ we use one global learning rate for all parameters; and for Adam, we use the optimally chosen group learning rate $\eta_{l}$ and initial condition $w_{0}$ for each block ${\mathbf{H}}_{l}$.
Assume either if i) certain regularity conditions are met (see proof in Appendix), or ii), if ${\mathbf{H}}$ is poorly-conditioned (its condition number is large enough). Then: regardless of its initialization, $\mathtt{GradWhitening}$ with a properly chosen learning rate will still have a strictly better convergence speed (i.e., smaller contraction factor) across all blocks $l\in[L]$ than Adam ($\beta_{1}=0,\beta_{2}=1$) under optimal group-wise learning rates and initial condition.
Remark
As pointed out by Zhang et al. (2024) and Da Silva & Gazeau (2020), Adam with $\beta_{2}<1$ will have issues with convergence, which will not be completely removed even with lr decay. Therefore, we will not discuss the case of $\beta_{2}<1$ to avoid the complication.
[…]
Appendix C Proof of Theorem 1
Proof.
We first consider the noiseless, full batch dynamics. Define ${\mathbf{V}}\in\mathbb{R}^{M_{C}\times n}$ as ${\mathbf{V}}:={\mathbf{U}}_{C}^{\top}{\mathbf{W}}$. Then following Theorem 2 in Tian et al. (2023), each column of ${\mathbf{V}}$ satisfies the following differential equation:
$\dot{{\mathbf{V}}}_{[:,j]}=\exp({\mathbf{V}}_{[:,j]}^{2}/2+C)\odot\mathbb{E}_{q}[g_{h_{j}}{\bm{x}}]$ (17)
The corresponding dynamics of attention score is given by:
${\bm{z}}_{q}=\frac{1}{2}\sum_{j}{\mathbf{V}}_{[:,j]}^{2}.$ (18)
Without loss of generality, in this proof we only consider $C=0$.
Now, following the argument of Lemma B.6 of Zhao et al. (2024a), we reparameterize the dynamics row-wise. For this, consider instead
${\mathbf{V}}=\begin{bmatrix}{\bm{u}}_{1}^{\top}\\ {\bm{u}}_{2}^{\top}\\ \vdots\\ {\bm{u}}_{M_{C}}^{\top}\end{bmatrix}$
Then, equation 17 becomes:
$\dot{{\bm{u}}}_{i}=[\exp({{\bm{u}}_{i}}^{2})\cdot\mathbf{1}]{\bm{\mu}}_{i}$ (19)
where ${\bm{\mu}}_{i}\in\mathbb{R}^{n\times 1}$ is given by $[{\bm{\mu}}_{i}]_{j}:=\mathbb{E}_{q}[g_{h_{j}}x_{i}]$. Therefore, it is clear that ${\bm{u}}_{i}$ always move along the direction of ${\bm{\mu}}_{i}$ due to the stationary back-propagated gradient assumption. Hence, $\dot{{\bm{u}}}_{i}=\alpha_{i}(t){\bm{\mu}}_{i}$ for some scalar dynamics $\alpha_{i}(t)$.
Next, consider the mini-batch version of the dynamics. In this case, the packpropagated gradient term $[{\bm{\mu}}_{i}]_{j}:=\mathbb{E}_{q}[g_{h_{j}}x_{i}]$ is corrupted by some i.i.d. mini-batch noise ${\bm{\xi}}$. The noisy row-wise dynamics now becomes:
$\dot{{\bm{u}}}_{i}=\alpha_{i}(t)({\bm{\mu}}_{i}+{\bm{\xi}}_{i})$ (20)
Therefore, after row-wise standardization, the new dynamics becomes
$\dot{\tilde{{\bm{u}}}}_{i}$ $=\frac{\alpha_{i}(t)(\mu_{i}+{\bm{\xi}}_{i})-\alpha_{i}(t)(\frac{1}{n}\sum_{j}\mu_{ij}+\frac{1}{n}\sum_{j}\xi_{ij})}{\alpha_{i}(t)(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij}-\frac{1}{n}\sum_{j}\mu_{ij}-\frac{1}{n}\sum_{j}\xi_{ij})^{2})}$
$=\frac{(\mu_{i}+{\bm{\xi}}_{i})-(\frac{1}{n}\sum_{j}\mu_{ij}+\frac{1}{n}\sum_{j}\xi_{ij})}{(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij}-\frac{1}{n}\sum_{j}\mu_{ij}-\frac{1}{n}\sum_{j}\xi_{ij})^{2})}$
Therefore, the normalized noisy gradient $\dot{\tilde{{\bm{u}}}}_{i}$ no longer depend on the time-variant component $\alpha(t)$. Hence, we have proved:
$\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{1})}]=\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{2})}]\quad\text{for all }t_{1},t_{2},\text{and}\,i.$
The corresponding result for $\tilde{{\mathbf{G}}}_{{\bm{z}}_{q}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}(\top{\bm{x}}^{(t)})}{{\bm{z}}_{q}})$ can be trivially derived due to Equation 18.
Appendix D Proof of Theorem 3
Proof.
We first show that $\nabla\mathcal{L}({\mathbf{W}}^{(0)})={\mathbf{H}}{\mathbf{W}}^{(0)}$ (and hence $\nabla\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})$ with $t\neq\infty$) are non-zero with probability 1 under Assumption of the theorem. Given $\nabla\mathcal{L}({\mathbf{W}}^{(0)})={\mathbf{H}}{\mathbf{W}}^{(0)}$, the set of matrices ${\mathbf{W}}^{(0)}$ such that $\text{Tr}({\mathbf{H}}{\mathbf{W}}^{(0)})=0$ forms a hyperplane in the space of $d\times d$ matrices. Specifically, it is defined by the linear equation: $\text{Tr}({\mathbf{H}}{\mathbf{W}}^{(0)})=0$. Since ${\mathbf{H}}$ is positive definite, at least one entry of ${\mathbf{H}}$ is non-zero. Thus, the hyperplane $\text{Tr}({\mathbf{H}}{\mathbf{W}}^{(0)})=0$ has zero Lebesgue measure in the space of $d\times d$ matrices. Given that ${\mathbf{W}}^{(0)}$ is sampled from a continuous distribution, the probability that $\text{Tr}({\mathbf{H}}{\mathbf{W}}^{(0)})=0$ is zero. Therefore, $\nabla\mathcal{L}({\mathbf{W}}^{(0)})\neq 0$ (and hence $\nabla\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})$ with $t\neq\infty$) with probability 1.
Next, we define the cost-to-go as:
$\mathcal{L}({\mathbf{W}}^{(t)})-\mathcal{L}^{*}=\frac{1}{2}\text{Tr}\left[({\mathbf{W}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}^{(t)}\right],$
and the per-step improvement is (since $\mathcal{L}^{*}=0$ under ${\mathbf{W}}=0$, ):
$\mathcal{L}({\mathbf{W}}^{(t)})-\mathcal{L}({\mathbf{W}}^{(t+1)})=\frac{1}{2}\text{Tr}\left[({\mathbf{W}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}^{(t)}\right]-\frac{1}{2}\text{Tr}\left[({\mathbf{W}}^{(t+1)})^{\top}{\mathbf{H}}{\mathbf{W}}^{(t+1)}\right].$
Substituting the update rule ${\mathbf{W}}^{(t+1)}={\mathbf{W}}^{(t)}-\eta\mathtt{GradWhitening}({\mathbf{G}}_{\text{whitened},l})={\mathbf{W}}^{(t)}-\eta{\mathbf{U}}{\mathbf{V}}^{\top}$, we get:
$\mathcal{L}({\mathbf{W}}^{(t)})-\mathcal{L}({\mathbf{W}}^{(t+1)})=\frac{1}{2}\text{Tr}\left[({\mathbf{W}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}^{(t)}\right]-\frac{1}{2}\text{Tr}\left[({\mathbf{W}}^{(t)}-\eta{\mathbf{U}}{\mathbf{V}}^{\top})^{\top}{\mathbf{H}}({\mathbf{W}}^{(t)}-\eta{\mathbf{U}}{\mathbf{V}}^{\top})\right].$
Expanding the right-hand side, we have
$\mathcal{L}({\mathbf{W}}^{(t)})-\mathcal{L}({\mathbf{W}}^{(t+1)})=\eta\text{Tr}\left[({\mathbf{W}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{U}}{\mathbf{V}}^{\top}\right]-\frac{\eta^{2}}{2}\text{Tr}\left[({\mathbf{U}}{\mathbf{V}}^{\top})^{\top}{\mathbf{H}}({\mathbf{U}}{\mathbf{V}}^{\top})\right].$
Now, noticing that ${\mathbf{G}}={\mathbf{H}}{\mathbf{W}}^{(t)}={\mathbf{U}}\Sigma{\mathbf{V}}^{\top}$, we have:
$\text{Tr}\left[({\mathbf{W}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{U}}{\mathbf{V}}^{\top}\right]=\text{Tr}\left[({\mathbf{H}}{\mathbf{W}}^{(t)})^{\top}{\mathbf{U}}{\mathbf{V}}^{\top}\right]=\text{Tr}\left[({\mathbf{U}}\Sigma{\mathbf{V}}^{\top})^{\top}{\mathbf{U}}{\mathbf{V}}^{\top}\right]=\text{Tr}\left[{\mathbf{V}}\Sigma{\mathbf{U}}^{\top}{\mathbf{U}}{\mathbf{V}}^{\top}\right]=\text{Tr}\left[{\mathbf{V}}\Sigma{\mathbf{V}}^{\top}\right].$
Since ${\mathbf{V}}$ is orthogonal, ${\mathbf{V}}^{\top}{\mathbf{V}}={\mathbf{I}}$, and $\Sigma$ is diagonal, we obtain:
$\text{Tr}\left[({\mathbf{W}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{U}}{\mathbf{V}}^{\top}\right]=\text{Tr}(\Sigma)=\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}.$
Similarly:
$\text{Tr}\left[({\mathbf{U}}{\mathbf{V}}^{\top})^{\top}{\mathbf{H}}({\mathbf{U}}{\mathbf{V}}^{\top})\right]=\text{Tr}\left[{\mathbf{V}}{\mathbf{U}}^{\top}{\mathbf{H}}{\mathbf{U}}{\mathbf{V}}^{\top}\right]=\text{Tr}\left[{\mathbf{V}}\Lambda{\mathbf{V}}^{\top}\right]=\text{Tr}({\mathbf{H}}),$
where $\Lambda$ is the eigenvalue matrix of ${\mathbf{H}}$. Given those intermediate results, we have:
$\frac{\mathcal{L}({\mathbf{W}}^{(t+1)})-\mathcal{L}^{*}}{\mathcal{L}({\mathbf{W}}^{(t)})-\mathcal{L}^{*}}$ $=1-\frac{\mathcal{L}({\mathbf{W}}^{(t)})-\mathcal{L}({\mathbf{W}}^{(t+1)})}{\mathcal{L}({\mathbf{W}}^{(t)})-\mathcal{L}^{*}}$
$=1-\frac{\eta\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}-\frac{\eta^{2}}{2}\text{Tr}({\mathbf{H}})}{\frac{1}{2}\text{Tr}\left[({\mathbf{W}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}^{(t)}\right]}.$
Noticing that this is a quadratic function of $\eta$ and the second order coefficient is positive, it is straightforward to verify via the quadratic formula that the optimal learning rate is given by
$\eta_{t}^{*}=\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}}{\text{Tr}({\mathbf{H}})}.$
Under which the optimal contraction factor is given by
$\frac{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t+1)})-\mathcal{L}^{*}}{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})-\mathcal{L}^{*}}=1-\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}\left[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\right]\text{Tr}({\mathbf{H}})}.$
Finally, if we additionally enforce ${\mathbf{W}}^{(0)}\sim V^{m\times n}(\mathbb{R})$, i.e., we can parameterize ${\mathbf{W}}^{(0)}={\mathbf{O}}$ where ${\mathbf{O}}$ is orthogonal, then it is trivial to verify that $\mathtt{GradWhitening}$ reaches the optimal solution with a 1-step update. To see this, consider the $\mathtt{GradWhitening}$ update:
${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathcal{P}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}\mathcal{P}({\mathbf{H}}{\mathbf{O}}),$
noticing that $\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}=1$, and $\mathcal{P}({\mathbf{H}}{\mathbf{O}})=\mathcal{P}({\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{O}})={\mathbf{Q}}{\mathbf{Q}}^{\top}{\mathbf{O}}={\mathbf{O}}$. Hence:
${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathcal{P}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-{\mathbf{O}}=\mathbf{0}={\mathbf{W}}^{*}.$
Hence, the proof is complete. ∎
Appendix E Proof of Proposition 2
Proof.
Since ${\mathbf{W}}_{\text{whitened}}^{(t)}\neq{\mathbf{W}}^{*}$, the square of the trace of the gradient $\text{Tr}[{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}$ must exceed some positive constant $C_{G}$, that is, $\text{Tr}[{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}>C_{G}$.
On the other hand, because:
  1. 1.
    The quadratic loss term $\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]$ is upper-bounded on $\mathbb{R}^{n\times n}$, and
  2. 2.
    $\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\neq 0$ (due to ${\mathbf{W}}_{\text{whitened}}^{(t)}\neq{\mathbf{W}}^{*}$),
we have that there exists a positive number $0<C_{\mathcal{L}}$ such that
$0<\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]<C_{\mathcal{L}}.$
Finally, since the norm of ${\mathbf{H}}$ is upper bounded, its trace must also be upper bounded by some constant $C_{H}$. Therefore, putting everything together, we have:
$Q>\frac{C_{G}^{2}}{C_{\mathcal{L}}C_{H}}.$
This inequality holds even as the condition number $\kappa\xrightarrow{}+\infty$. ∎
Appendix F Proof of Proposition 3
To prove Proposition 3, we first generalize existing work on the convergence rate lower bound (via contraction factor) of gradient descent and Adam (we only consider $\beta_{2}=1$) under the same setting:
Theorem 4 (Contraction factor lower bound for gradient descent, generalized based on Zhang et al. (2024)).
Consider the optimization problem in Equation 12. Let ${\mathbf{W}}_{GD}^{t}$ be the output of GD after $t$ steps. Then, for any step size $\eta$, there exists an initial condition such that the following lower bound on the contraction rate holds:
$\mathcal{L}({\mathbf{W}}_{GD}^{t+1})-\mathcal{L}^{*}\geq\left(1-\frac{2}{\kappa+1}\right)\left(\mathcal{L}({\mathbf{W}}_{GD}^{t})-\mathcal{L}^{*}\right),$
where $\mathcal{L}^{*}=\mathcal{L}({\mathbf{W}}^{*})$. Furthermore, under optimal $\eta=\frac{2}{\lambda_{1}+\lambda_{m}}$, the bound becomes tight regardless of the settings of ${\mathbf{H}}$, where $\lambda_{1}$ and $\lambda_{m}$ are the largest and smallest eigen values of ${\mathbf{H}}$, respectively.
Proof.
The proposition 1 in Zhang et al. (2024) has shown that the lower bound holds for diagonal positive definite Hessian ${\mathbf{H}}$. To show that the lower bound holds for a general positive definite Hessian ${\mathbf{H}}$ we will reformulate the problem to align with the setup in diagonal case (Prposition 1 of Zhang et al. (2024)).
First, for any positive definite Hessian ${\mathbf{H}}$, we can perform an eigen decomposition ${\mathbf{H}}={\mathbf{U}}{\mathbf{S}}{\mathbf{U}}^{\top}$, where ${\mathbf{U}}$ is an orthogonal matrix and ${\mathbf{S}}$ is a diagonal matrix containing the eigenvalues of ${\mathbf{H}}$. Define a change of variables ${\mathbf{Z}}={\mathbf{U}}^{\top}{\mathbf{W}}$. Then, the optimization problem becomes
$\mathcal{L}({\mathbf{Z}})=\frac{1}{2}\text{Tr}({\mathbf{Z}}^{\top}{\mathbf{S}}{\mathbf{Z}}),$
which reduces the problem to the case of a diagonal ${\mathbf{H}}$ with condition number $\kappa=\frac{\lambda_{1}}{\lambda_{m}}$, where $\lambda_{1}$ and $\lambda_{m}$ are the largest and smallest eigenvalues of ${\mathbf{H}}$, respectively.
Thus, by applying Proposition 1 of Zhang et al. (2024) to this transformed problem, we conclude that there exists initial point such that the lower bound on the contraction rate
$\mathcal{L}({\mathbf{W}}_{GD}^{(t+1)})-\mathcal{L}^{*}\geq\left(1-\frac{2}{\kappa+1}\right)\left(\mathcal{L}({\mathbf{W}}_{GD}^{(t)})-\mathcal{L}^{*}\right)$
holds for the transformed variables ${\mathbf{Z}}$ and, equivalently, for the original variables ${\mathbf{W}}$ since the condition number is preserved under orthogonal transformations.
Therefore, the lower bound for gradient descent applies to any general positive definite Hessian ${\mathbf{H}}$ provided the condition number $\kappa$ remains unchanged.
Finally, under the optimal step size $\eta=\frac{2}{\lambda_{1}+\lambda_{m}}$, the bound becomes tight regardless of the settings of ${\mathbf{H}}$. This is achieved by selecting $\eta$ to minimize the contraction factor, aligning with well-known results regarding the optimal convergence rate of gradient descent on quadratic objectives (Nesterov, 2013).
This completes the proof of Theorem 4. ∎
Corollary 1 (Lower bound on Adam ($\beta_{2}=1$)).
Consider the optimization problem in Equation 12. Assume the weight initialization ${\mathbf{W}}^{0}$ assigned zero probability to any set of zero Lebesgue measure in $\mathbb{R}^{m\times n}$. Let ${\mathbf{W}}_{\text{Adam}}^{(t)}$ be the parameter after $t$ iterations of Adam with hyperparameters $\beta_{1}=0$ and $\beta_{2}=1$. Then, for any step size $\eta$, the following lower bound on the contraction rate holds:
$\mathcal{L}({\mathbf{W}}_{Adam}^{(t+1)})-\mathcal{L}^{*}\geq\left(1-\frac{2}{\kappa^{\prime}({\mathbf{W}}^{0})+1}\right)\left(\mathcal{L}({\mathbf{W}}_{Adam}^{(t)})-\mathcal{L}^{*}\right),$
where $\kappa^{\prime}({\mathbf{W}}^{0})$ is the ${\mathbf{W}}^{0}$-dependent condition number of the preconditioned Hessian $\text{diag}(|{\mathbf{H}}{\mathbf{W}}^{0}|^{-1}){\mathbf{H}}$, and $\mathcal{L}^{*}=\mathcal{L}({\mathbf{W}}^{*})$.
Proof.
The update rule of Adam with $\beta_{1}=0$ and $\beta_{2}=1$ is given by Zhang et al. (2024):
${\mathbf{W}}_{\text{Adam}}^{(t+1)}={\mathbf{W}}_{\text{Adam}}^{(t)}+\eta\,\text{diag}(|{\mathbf{H}}{\mathbf{W}}^{(0)}|^{-1})\,{\mathbf{H}}{\mathbf{W}}_{\text{Adam}}^{(t)}.$
This can be interpreted as gradient descent with a preconditioned Hessian matrix $\text{diag}(|{\mathbf{H}}{\mathbf{W}}^{(0)}|^{-1}){\mathbf{H}}$. By applying Theorem 4, we conclude that the contraction rate for Adam under these settings satisfies the lower bound:
$\mathcal{L}({\mathbf{W}}_{\text{Adam}}^{(t+1)})-\mathcal{L}^{*}\geq\left(1-\frac{2}{\kappa+1}\right)\left(\mathcal{L}({\mathbf{W}}_{\text{Adam}}^{(t)})-\mathcal{L}^{*}\right),$
where $\kappa$ is the condition number of the Hessian matrix ${\mathbf{H}}$.
Therefore, the proof is complete. ∎
Next, We extend our Theorem 3 to block-diagonal Hessian case to prepare for discussions on group learning rates when comparing to Adam.
Corollary 2 (Upper Bound Convergence Rate of SWAN).
Consider the same quadratic loss function $\mathcal{L}({\mathbf{W}})=\frac{1}{2}\text{Tr}({\mathbf{W}}^{\top}{\mathbf{H}}{\mathbf{W}})$ with ${\mathbf{H}}$ being block-diagonal. That is, ${\mathbf{H}}=\text{diag}({\mathbf{H}}_{1},{\mathbf{H}}_{2},\dots,{\mathbf{H}}_{L}),$ where each $H_{l}\in\mathbb{R}^{m_{l}\times n_{l}}$ is a positive definite matrix for $l=1,2,\dots,L$, and $\sum_{l=1}^{L}m_{l}=m$ and $\sum_{l=1}^{L}n_{l}=n$. Assume the initialization distribution of ${\mathbf{W}}^{(0)}$ assignes zero probability to any zero measure set in $\mathbb{R}^{m\times n}$. Let ${\mathbf{W}}_{\text{whitened}}^{(t)}$ be the parameter matrix after $t$ iterations of the SWANoptimizer defined in Theorem 3, with learning rate $\eta$. Then, under the conditions that  [^6] Note that according to Proposition 2, this is always achievable when ${\mathbf{H}}_{l}$ is poorly-conditioned ($\frac{\lambda_{l,d_{l}}}{\lambda_{l,1}}$ is small enough). The lower interval above converges to zero, and one can simply pick e.g., $\eta=\min_{l\in[L]}\frac{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})}{\text{Tr}({\mathbf{H}}_{l})}$.:
$\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}^{2}-\text{Tr}({\mathbf{H}}_{l})\cdot\text{Tr}\left(({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\right)\cdot\frac{2\lambda_{l,m_{l}}}{\lambda_{l,1}+\lambda_{l,m_{l}}}>0,$
where $\lambda_{l,1}$ and $\lambda_{l,m_{l}}$ are the largest and smallest singular value of ${\mathbf{H}}_{l}$, respectively; then there exists a proper learning rate $\eta$ such that: with probability 1, the loss satisfies:
$\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t+1)})-\mathcal{L}^{*}<\max_{l\in[L]}\left(1-\frac{2}{\kappa_{l}+1}\right)\left(\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})-\mathcal{L}^{*}\right),$ (21)
where $\kappa_{l}$ is the condition number of ${\mathbf{H}}_{l}$.
Proof.
Applying the arguments in the proof of Theorem 3 to each block $l$, we have (for simplicity, we will drop the subscript “whitened” when there is no confusion):
$\frac{\mathcal{L}({\mathbf{W}}_{l}^{(t+1)})-\mathcal{L}^{*}}{\mathcal{L}({\mathbf{W}}_{l}^{(t)})-\mathcal{L}^{*}}$ $=1-\frac{\mathcal{L}({\mathbf{W}}_{l}^{(t)})-\mathcal{L}({\mathbf{W}}_{l}^{(t+1)})}{\mathcal{L}({\mathbf{W}}_{l}^{(t)})-\mathcal{L}^{*}}$
$=1-\frac{\eta\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}-\frac{\eta^{2}}{2}\text{Tr}({\mathbf{H}}_{l})}{\frac{1}{2}\text{Tr}\left[({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\right]}.$
It is straightforward to verify via the quadratic formula that if one chooses $\eta$ satisfying:
$\frac{\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}-\sqrt{\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}^{2}-\text{Tr}({\mathbf{H}}_{l})\cdot\text{Tr}(({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})\cdot\frac{2\lambda_{l,m_{l}}}{\lambda_{l,1}+\lambda_{l,m_{l}}}}}{\text{Tr}({\mathbf{H}}_{l})}<\eta$
$<\frac{\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}+\sqrt{\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}^{2}-\text{Tr}({\mathbf{H}}_{l})\cdot\text{Tr}(({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})\cdot\frac{2\lambda_{l,m_{l}}}{\lambda_{l,1}+\lambda_{l,m_{l}}}}}{\text{Tr}({\mathbf{H}}_{l})},$
then we have:
$\frac{\frac{1}{2}\text{Tr}\left[({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\right]}{\eta\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}-\frac{\eta^{2}}{2}\text{Tr}({\mathbf{H}}_{l})}<\frac{\kappa_{l}+1}{2}.$
Rearranging, we obtain:
$\frac{\mathcal{L}({\mathbf{W}}_{l}^{(t+1)})-\mathcal{L}^{*}}{\mathcal{L}({\mathbf{W}}_{l}^{(t)})-\mathcal{L}^{*}}<1-\frac{2}{\kappa_{l}+1}.$
Since ${\mathbf{H}}$ is block-diagonal, the updates for each block $l$ are independent. Summing the loss over all blocks, we obtain:
$\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t+1)})-\mathcal{L}^{*}=\sum_{l=1}^{L}\left(\mathcal{L}({\mathbf{W}}_{l}^{(t+1)})-\mathcal{L}^{*}\right)<\sum_{l=1}^{L}\left(1-\frac{2}{\kappa_{l}+1}\right)\left(\mathcal{L}({\mathbf{W}}_{l}^{(t)})-\mathcal{L}^{*}\right).$
Taking the maximum contraction factor across all blocks:
$\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t+1)})-\mathcal{L}^{*}<$ $\max_{l\in[L]}\left(1-\frac{2}{\kappa_{l}+1}\right)\sum_{l=1}^{L}\left(\mathcal{L}({\mathbf{W}}_{l}^{(t)})-\mathcal{L}^{*}\right)$
$=$ $\max_{l\in[L]}\left(1-\frac{2}{\kappa_{l}+1}\right)\left(\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})-\mathcal{L}^{*}\right).$
Thus, the overall contraction factor for the SWANoptimizer is:
$\rho_{\text{SWAN}}=\max_{l\in[L]}\left(1-\frac{2}{\kappa_{l}+1}\right).$
Finally, we are ready to prove Proposition 3:
Proposition 2 ($\mathtt{GradWhitening}$ with single lr vs Adam with tuned group lr).
Consider the optimization problem Equation 12. Assume ${\mathbf{H}}$ is block-diagonal, i.e., ${\mathbf{H}}=\text{diag}({\mathbf{H}}_{1},{\mathbf{H}}_{2},\dots,{\mathbf{H}}_{L}),$ where each ${\mathbf{H}}_{l}\in\mathbb{R}^{m_{l}\times m_{l}}$ is a positive definite matrix for $l=1,2,\dots,L$, and $\sum_{l=1}^{L}m_{l}=m$. Assuming for $\mathtt{GradWhitening}$ we use one global learning rate for all parameters; and for Adam, we use the optimally chosen group learning rate $\eta_{l}$ and initial condition $w_{0}$ for each block ${\mathbf{H}}_{l}$. Assume either if i) certain regularity conditions are met (see proof in Appendix), or ii), if ${\mathbf{H}}$ is poorly-conditioned (its condition number is large enough). Then: regardless of its initialization, $\mathtt{GradWhitening}$ with a properly chosen learning rate will still have a strictly better convergence speed (i.e., smaller contraction factor) across all blocks $l\in[L]$ than Adam ($\beta_{1}=0,\beta_{2}=1$) under optimal group-wise learning rates and initial condition.
Proof.
For simplicity, we will drop the subscript “whitened” when there is no confusion. Let $\kappa^{\prime}_{l}({\mathbf{W}}_{l}^{(0)})$ denote the ${\mathbf{W}}_{l}^{(0)}$-dependent condition number of the $l$-th block preconditioned Hessian $\text{diag}(\left|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(0)}\right|^{-1}){\mathbf{H}}_{l}$. Let $\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})$ and $\lambda_{l,1}({\mathbf{W}}_{l}^{(0)})$ be the smallest and largest eigenvalues of $\text{diag}(\left|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(0)}\right|^{-1}){\mathbf{H}}_{l}$, respectively. Then,
Case 1:
Under the conditions that:
  1. 1.
    Existence of roots: $\forall l\in[L]$,
    $\|{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\|_{1}^{2}-\text{Tr}({\mathbf{H}}_{l})\cdot\text{Tr}\left(({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\right)\cdot\frac{2\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}{\lambda_{l,1}({\mathbf{W}}_{l}^{(0)})+\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}>0,$
    and
  2. 2.
    Overlap condition:
    $\min_{l\in[L]}\frac{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})+\sqrt{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})^{2}-\text{Tr}({\mathbf{H}}_{l})\cdot\text{Tr}\left(({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\right)\cdot\frac{2\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}{\lambda_{l,1}({\mathbf{W}}_{l}^{(0)})+\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}}}{\text{Tr}({\mathbf{H}}_{l})}$
    $>\max_{l\in[L]}\frac{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})-\sqrt{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})^{2}-\text{Tr}({\mathbf{H}}_{l})\cdot\text{Tr}\left(({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\right)\cdot\frac{2\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}{\lambda_{l,1}({\mathbf{H}}_{l}^{(0)})+\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}}}{\text{Tr}({\mathbf{H}}_{l})}.$
Then, there exists a global learning rate $\eta$, such that for all $l\in[L]$,
$\frac{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t+1)})_{l}-\mathcal{L}^{*}_{l}}{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})_{l}-\mathcal{L}^{*}_{l}}<1-\frac{2}{\kappa^{\prime}_{l}({\mathbf{W}}_{l}^{(0)})+1}\leq\frac{\mathcal{L}({\mathbf{W}}_{\text{Adam}}^{(t+1)})_{l}-\mathcal{L}^{*}_{l}}{\mathcal{L}({\mathbf{W}}_{\text{Adam}}^{(t)})_{l}-\mathcal{L}^{*}_{l}}.$
Case 2:
If ${\mathbf{H}}$ is poorly-conditioned, i.e., $\frac{\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}{\lambda_{l,1}({\mathbf{W}}_{l}^{(0)})}\xrightarrow{}0$, then Proposition 2 asserts that the following term
$\max_{l\in[L]}\frac{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})-\sqrt{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})^{2}-\text{Tr}({\mathbf{H}}_{l})\cdot\text{Tr}\left(({\mathbf{W}}_{l}^{(t)})^{\top}{\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)}\right)\cdot\frac{2\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}{\lambda_{l,1}({\mathbf{W}}_{l}^{(0)})+\lambda_{l,m_{l}}({\mathbf{W}}_{l}^{(0)})}}}{\text{Tr}({\mathbf{H}}_{l})}\xrightarrow{}0,$
and one can simply choose, for example, $\eta=\min_{l\in[L]}\frac{\text{Tr}({\mathbf{H}}_{l}{\mathbf{W}}_{l}^{(t)})}{\text{Tr}({\mathbf{H}}_{l})}$. Under this choice of $\eta$, we still have
$\frac{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t+1)})_{l}-\mathcal{L}^{*}_{l}}{\mathcal{L}({\mathbf{W}}_{\text{whitened}}^{(t)})_{l}-\mathcal{L}^{*}_{l}}<1-\frac{2}{\kappa^{\prime}_{l}({\mathbf{W}}_{l}^{(0)})+1}\leq\frac{\mathcal{L}({\mathbf{W}}_{\text{Adam}}^{(t+1)})_{l}-\mathcal{L}^{*}_{l}}{\mathcal{L}({\mathbf{W}}_{\text{Adam}}^{(t)})_{l}-\mathcal{L}^{*}_{l}}$
for all $l\in[L]$. ∎
Appendix G Proof of Proposition 1
Proof.
First, define ${\mathbf{V}}\in\mathbb{R}^{M_{C}\times n}$ as ${\mathbf{V}}:={\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, and consider the Hessian with respect to ${\mathbf{V}}$ instead of ${\mathbf{W}}$. Notice that although the loss function $\mathcal{L}$ is unknown, its first-order derivatives are known. Specifically, they are given by:
$\frac{\partial\mathcal{L}}{\partial v_{lk}}=\dot{v}_{lk}=\mathbb{E}_{q=m}\left[g_{h_{k}}x_{l}\right]e^{\frac{1}{2}\sum_{s}v_{ls}^{2}}.$
Therefore, the second-order derivatives, i.e., the Hessian matrix ${\mathbf{H}}({\mathbf{V}})$, are:
${\mathbf{H}}({\mathbf{V}})_{lk,l^{\prime}k^{\prime}}=\frac{\partial^{2}\mathcal{L}}{\partial v_{lk}\partial v_{l^{\prime}k^{\prime}}}=\frac{\partial\left[\mathbb{E}_{q=m}\left[g_{h_{k}}x_{l}\right]e^{\frac{1}{2}\sum_{s}v_{ls}^{2}}\right]}{\partial v_{l^{\prime}k^{\prime}}}=\dot{v}_{lk}v_{l^{\prime}k^{\prime}}\delta_{ll^{\prime}},$
where $\delta_{ll^{\prime}}$ is the Kronecker delta, which is $1$ if $l=l^{\prime}$ and 00 otherwise.
Based on Lemma B.6 in Zhao et al. (2024a), as $t\to\infty$, there exists an index subset $O_{l}\subset\{1,\dots,M_{C}\}$ such that:
$v_{l^{*}k}\gg v_{lk},\quad\dot{v}_{l^{*}k}\gg\dot{v}_{lk},\quad\forall l^{*}\in O_{l},\;l\notin O_{l},\;\forall k.$
Consequently,
${\mathbf{H}}({\mathbf{V}})_{l^{*}k,l^{*}k^{\prime}}\gg{\mathbf{H}}({\mathbf{V}})_{lk,l^{\prime}k^{\prime}},\quad\forall l^{*}=l^{\prime*}\in O_{l},\;l,l^{\prime}\notin O_{l},\;\forall k,k^{\prime}.$
After normalization, as $t\to\infty$, we have
$\frac{{\mathbf{H}}({\mathbf{V}})_{lk,l^{\prime}k^{\prime}}}{\sum_{l,l^{\prime}}{\mathbf{H}}({\mathbf{V}})_{lk,l^{\prime}k^{\prime}}}\xrightarrow{t\to\infty}\begin{cases}1,&\text{if }l=l^{\prime}\in O_{l},\\ 0,&\text{otherwise}.\end{cases}$
Reverting back to the ${\mathbf{W}}$ space, we have
${\mathbf{H}}({\mathbf{W}})=({\mathbf{I}}_{K}\otimes{\mathbf{U}}_{C}){\mathbf{H}}({\mathbf{V}})({\mathbf{I}}_{K}\otimes{\mathbf{U}}_{C})^{\top},$
where $\otimes$ denotes the Kronecker product and ${\mathbf{I}}_{K}$ is the identity matrix of appropriate dimensions.
Therefore, for all $1\leq s,s^{\prime}\leq d$ and $1\leq k,k^{\prime}\leq n$, we obtain
$\frac{{\mathbf{H}}({\mathbf{W}})_{sk,s^{\prime}k^{\prime}}}{\sum_{s,s^{\prime}}{\mathbf{H}}({\mathbf{W}})_{sk,s^{\prime}k^{\prime}}}=\frac{{\mathbf{H}}({\mathbf{W}})_{sk^{\prime},s^{\prime}k^{\prime}}}{\sum_{s,s^{\prime}}{\mathbf{H}}({\mathbf{W}})_{sk^{\prime},s^{\prime}k^{\prime}}}\quad\text{as }t\to\infty.$
This holds for all $1\leq s,s^{\prime}\leq d$ and $1\leq k,k^{\prime}\leq n$. ∎
Appendix H Proof of Theorem 2
Proof.
Assume that we have a function $f({\mathbf{W}})$ defined on ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ that we want to minimize, and the global minimum is at ${\mathbf{W}}^{*}$ (without loss of generality, let ${\mathbf{W}}^{*}=0$). Then our problem can be written as:
$\min_{{\mathbf{W}}\in\mathbb{R}^{m\times n}}f({\mathbf{W}})$ (22)
Assume $f$ is a twice differentiable function and has a Lipschitz continuous Hessian around ${\mathbf{W}}^{*}$, then both $\mathtt{GradWhitening}$ and Newton’s method approximate the loss equation 22 up to second order (ignoring first-order terms without loss of generality) as:
$\min_{{\mathbf{w}}\in\mathbb{R}^{mn}}\frac{1}{2}{\mathbf{w}}^{T}{\mathbf{H}}{\mathbf{w}},$ (23)
where ${\mathbf{w}}\in\mathbb{R}^{mn\times 1}$ is the flattened version of ${\mathbf{W}}$, and ${\mathbf{H}}\in\mathbb{R}^{mn\times mn}$ is the Hessian matrix of $f$ at ${\mathbf{W}}^{*}$. We may apply second-order Newton’s method to optimize equation 23, which gives the following update:
${\mathbf{w}}^{t+1}={\mathbf{w}}^{t}-\eta{\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t}).$ (24)
However, this update is very expensive and requires $\mathcal{O}(m^{3}n^{3})$ computation. SWANcan then be viewed as a fast approximation to estimate the update of equation 24. Under our assumption on Hessian structure, we have:
${\mathbf{H}}=\text{diag}(\tilde{{\mathbf{H}}},\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}})$ (25)
where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
Under this assumption, the optimization problem equation 23 can be rewritten as:
$\min_{{\mathbf{W}}\in\mathbb{R}^{m\times n}}\frac{1}{2}\text{Tr}\left[{\mathbf{W}}^{\top}\tilde{{\mathbf{H}}}{\mathbf{W}}\right],$ (26)
For this model, the gradient of ${\mathbf{W}}$ is given by:
$\nabla f({\mathbf{W}})=\tilde{{\mathbf{H}}}{\mathbf{W}}.$
Let ${\mathbf{W}}$ have rank $m$, and assume $n>m$. We say ${\mathbf{W}}$ is now over-parameterized in the sense that its dimensionality is now larger than the dimensionality of the second-order information $\tilde{{\mathbf{H}}}$. It is now clear that the second-order information is losslessly compressed into gradients: one can exactly recover the Hessian via
${\mathbf{H}}=\text{diag}(\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\dots,\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\dots,\nabla f({\mathbf{W}}){\mathbf{W}}^{-1}).$
Given this fact, one can easily estimate Newton’s update ${\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t})$ without explicitly estimating ${\mathbf{H}}^{-1}$ (or $\tilde{{\mathbf{H}}}$). Notice that:
${\mathbf{H}}={\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}$
where ${\mathbf{I}}_{m}$ is the $m\times m$ identity matrix and $\otimes$ denotes the Kronecker product. Given that
${\mathbf{H}}^{-1}={\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}^{-1},$
we have
${\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t})=({\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}^{-1})\,\text{vec}(\nabla f({\mathbf{W}}^{t}))=\text{vec}\left(\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t})\right).$
Reshaping back to matrix form, we can rewrite the update rule as:
$\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t}).$
Next, since $\tilde{{\mathbf{H}}}$ is positive definite:
$\nabla f({\mathbf{W}}^{t})=\tilde{{\mathbf{H}}}{\mathbf{W}}^{t}={\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{W}}^{t},$
where ${\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}$ is the eigen decomposition of $\tilde{{\mathbf{H}}}$. Assume further that ${\mathbf{W}}^{t}$ is initialized as orthogonal; one can immediately recognize that ${\mathbf{Q}}$ is also the left singular vector matrix of $\nabla f({\mathbf{W}}^{t})$. Therefore, the matrix inversion of $\tilde{{\mathbf{H}}}$ required in Newton’s update equation 24 can be easily achieved by simply deleting the $\Lambda$ term. That is,
$\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t})={\mathbf{Q}}\Lambda^{-1}{\mathbf{Q}}^{\top}{\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{W}}^{t}$
$=$ ${\mathbf{Q}}\left[{\mathbf{Q}}^{\top}{\mathbf{W}}^{t}\right]\overset{{\mathbf{U}}:={\mathbf{Q}},{\mathbf{V}}:=({\mathbf{W}}^{t})^{\top}{\mathbf{Q}}}{=}{\mathbf{U}}{\mathbf{V}}^{\top}=\mathcal{P}\left(\nabla f({\mathbf{W}}^{t})\right)={\mathbf{G}}_{\text{whitened}}.$
This is exactly the $\mathtt{GradWhitening}$-modified gradient update. ∎
Appendix I Implementation details
Setup
We describe the implementation setups for SWAN used in LLM pre-training tasks. To enable a more straightforward and comparable analysis, we simply replicate the setting of Zhao et al. (2024a), under exactly the same model configs and optimizer configs. This include the same model architecture, tokenizer, batch size, context length, learning rate scheduler, etc. The only exception is that since SWAN does not need warm up, we use zero-warm up ratio for SWAN. But we still ensure that the learning rates of all methods are fully decayed to the same target learning rate (0.1 $\times$ inital learning rate). Across all model sizes, we use the same hyperparameter settings and the same learning rate tuning procedure. The only exception is that for model of size 1.3B: as we already know that larger model requires smaller learning rates, we conduct learning search for Adam and Galore over a smaller but more fine grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$. As a result, the optimal learning rate found for Adam on 1.3B is 0.0007.
Applying SWAN on LLMs
Since SWAN utilizes matrix-level operations on gradients, it can only be applied on 2D parameters. Therefore, in our experiments we only apply $SWAN{}$ on all linear projection weights in transformer blocks. Similar to Galore (Zhao et al., 2024a), the rest of the non-Linear parameters still uses Adam as default choice. Therefore, we follow the learning rate setup of Galore, where we fix a global learning rate 0.001 across all model sizes and all modules. Then, for the linear projection modules where SWAN is applied, we simply apply a scaling factor $\alpha$ on top of the global learning rate. We choose $\alpha=1$ for all model sizes except for 1.3 B case, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$. Note that this choice is quite random; and it is not extensively grid-searched.
Applying $\mathtt{GradWhitening}$
We found that one strategy that lead to better convergence is to first convert the gradient matrice to FP32, perform $\mathtt{GradWhitening}$, and then convert back to BF16. This is the default setup in Section 5.3. In practice, we found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence. We found that setting $\beta=0.8$ with iteration number $k=10$ generally give satisfactory performance. This can be further optimized and we leave it for future work. Alternatively, $\mathtt{GradWhitening}$ can also be done in BF16 precision. The setup we use for BF16 precision is $\beta=0.85$, $k=6$. These hyperparameters are selected on randomly sampled matrices via minimizing inverse square root estimation loss; then directly applied to LLM without any further tuning. We found that this usually translates well.

### v2

@@
-SWAN: Preprocessing SGD Enables Adam-Level Performance On LLM Training With Significant Memory Reduction
+SWAN: SGD with Normalization and Whitening Enables Stateless LLM Training
 Abstract
 Adaptive optimizers such as Adam (Kingma & Ba, 2015) have been central to the success of large language models.
-However, they maintain additional moving average states throughout training, which results in memory requirements several times greater than the model.
+However, they often require maintaining optimizer states throughout training, which can result in memory requirements several times greater than the model footprint.
 This overhead imposes constraints on scalability and computational efficiency.
-On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency, their capability in LLM training is limited (Zhao et al., 2024b).
-To address this dilemma, we show that pre-processing SGD is sufficient to reach Adam-level performance on LLMs.
-Specifically, we propose to preprocess the instantaneous stochastic gradients with two simple operators: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$.
-$\mathtt{GradNorm}$ stabilizes gradient distributions, and $\mathtt{GradWhitening}$ counteracts the local curvature of the loss landscape, respectively.
-This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any accumulative state variables.
+Stochastic Gradient Descent (SGD), in contrast, is a stateless optimizer, as it does not track state variables during training.
+Consequently, it achieves optimal memory efficiency.
+However, its capability in LLM training is limited (Zhao et al., 2024b).
+In this work, we show that pre-processing SGD in a stateless manner can achieve the same performance as the Adam optimizer for LLM training, while drastically reduces the memory cost.
+Specifically, we propose to pre-process the instantaneous stochastic gradients using normalization and whitening.
+We show that normalization stabilizes gradient distributions, and whitening counteracts the local curvature of the loss landscape.
+This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any optimizer states.
 Empirically, SWAN has the same memory footprint as SGD, achieving $\approx 50\%$ reduction on total end-to-end memory compared to Adam.
-In language modeling tasks, SWAN demonstrates the same or even a substantial improvement over Adam.
-Specifically, when pre-training the LLaMa model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity in less than half tokens seen.
+In language modeling tasks, SWAN demonstrates comparable or even better performance than Adam: when pre-training the LLaMA model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity using half as many tokens.
 1 Introduction
-Adaptive optimizers, such as Adam and its variants (Kingma & Ba, 2015; Loshchilov & Hutter, 2019; Shazeer & Stern, 2018; Pagliardini et al., 2024; Liu et al., 2023; Zhao et al., 2024a), have been central to the success of training large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Dubey et al., 2024; Bi et al., 2024; Bai et al., 2023; Zhang et al., 2022).
+Adaptive optimizers, such as Adam and its variants (Kingma & Ba, 2015; Loshchilov & Hutter, 2019; Shazeer & Stern, 2018; Pagliardini et al., 2024; Liu et al., 2023; Zhao et al., 2024a), have been central to the success of training large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023b; Dubey et al., 2024; Bi et al., 2024; Bai et al., 2023; Zhang et al., 2022).
 However, most adaptive optimizers for LLMs are stateful, meaning they require tracking and maintaining internal states.
@@
 For instance, Adam (Kingma & Ba, 2015) – the de facto optimizer for LLM training – involves the tracking of exponential moving averages (EMAs), effectively doubling memory requirements.
-AdEMAMix (Pagliardini et al., 2024) – an extension of Adam that achieves significant convergence speed boost – requires storing even more states, tripling the memory requirements.
+AdEMAMix (Pagliardini et al., 2024), an extension of Adam that achieves significant convergence speed boost, requires storing even more states, tripling the memory requirements.
 This overhead can be significant especially in distributed settings, where the optimizer states could consume a significant amount of the GPU memory (Dubey et al., 2024; Korthikanti et al., 2023).
-On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency (i.e., they are stateless), their capability to train LLMs is limited (Zhao et al., 2024b), and the performance gap between SGD and Adam is significant (Zhang et al., 2020; Kunstner et al., 2023; 2024).
-To address the memory-performance dilemma for LLMs, this work questions the requirement of accumulative state variables during training to achieve good performance in LLMs.
-We propose to simply preprocess the instantaneous SGD gradient in a history-independent manner, such that it still preserves certain desirable properties of adaptive optimizers.
-This result in SWAN (SGD with Whitening And Normalization), a novel SGD-like optimizer that eliminates all internal optimizer states and empirically surpasses the performance Adam on several LLM pre-training scenarios.
-Our proposed optimizer consists of two stateless operators applied to the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$.
-$\mathtt{GradNorm}$ aims to stabilize the stochasticity of gradient distributions during training, while $\mathtt{GradWhitening}$ seeks to neutralize the local geometry of the loss landscape.
-Both operators enables SWAN to rely solely on the statistics of the current gradient matrices.
-This SGD preprocessing approach eliminates all internal optimizer states, and therefore reproduces the memory consumption of SGD, reaching a new state-of-the-art in terms of optimizer memory efficiency for LLMs ($\approx 50\%$ reduction on end-to-end memory footprint vs Adam, and $\approx 100\%$ reduction in optimizer states.
-See Figure 1).
-In addition to memory savings, SWAN also demonstrates significant computational efficiency: our empirical evaluations on pre-training LLaMA models on the C4 dataset with multiple model sizes show consistently better performance than Adam and other low-rank memory-efficient optimizers.
-Remarkably, at the 350M and 1.3B scale, our method achieves 2X faster convergence in terms of tokens seen compared to Adam, marking SWAN the first known stateless optimizer converges as fast/faster than Adam in training large language models.
+On the other hand, while stochastic gradient descent (SGD) is optimal in terms of memory efficiency (i.e., it is stateless), their capability to train LLMs is limited (Zhao et al., 2024b; Zhang et al., 2020; Kunstner et al., 2023; 2024).
+Therefore, a natural question arises:
+Can LLMs be trained efficiently using stateless optimizers?
+There is a growing body of research that has contributed to answering this question positively by developing novel optimizers that reduce the memory requirements associated with tracking internal state variables during the training of LLMs while achieving similar or even speedup boost performance compared to Adam.
+For instance, some methods rely solely on tracking the first moment of gradients (Xu et al., 2024a; Jordan et al., 2024), while others introduce an additional one-dimensional tracking variable on top of first moments (Zhang et al., 2024b; Zhao et al., 2024c).
+Alternatively, approaches focusing exclusively on pre-conditioner tracking have also been proposed (Pooladzandi & Li, 2024; Li, 2017).
+Another line of work focuses on using low-rank approximations to store the first and second moments, thereby reducing the memory cost associated with tracking optimizer states (Lialin et al., 2023; Hao et al., 2024; Zhao et al., 2024a).
+More recently, Zhu et al.
+(2024) introduced an almost stateless optimizer for efficient LLM training, which proposes to only track rank-1 state variables during training.
+In this work, we address this question by proposing to simply pre-process the instantaneous stochastic gradient in a stateless manner.
+The result is SWAN (SGD with Whitening And Normalization), a novel stochastic optimizer that eliminates all internal optimizer states and empirically achieves comparable or even better performance compared to Adam on several LLM pre-training tasks.
+Our optimizer consists of combining two well-known operators to pre-process the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$.
+$\mathtt{GradNorm}$ applies a row-wise standardization on the gradient matrix, while $\mathtt{GradWhitening}$ orthogonalizes the normalized gradient matrix.
+We show that these operators aims at stabilizing the stochasticity of gradient distributions during training, and neutralizing the local geometry of the loss landscape, respectively.
+When applied together, both operators enables SWAN to rely solely on the statistics of the current gradient matrices.
+This approach eliminates the need to track state variables, thereby matching the memory footprint of SGD.
+In addition to memory savings, SWAN also demonstrates significant computational efficiency: our empirical evaluations on pre-training LLaMA (Touvron et al., 2023a) models on the C4 dataset with multiple model sizes show consistently the same or better performance than Adam and other low-rank optimizers.
+Remarkably, at the 350M and 1.3B scale, our method achieves up to 2X faster convergence in terms of tokens seen compared to Adam.
+Our contributions are summarized below:
+[…]
+    1. 1.
+      Robustness to ill-conditioned problems: on LLM pre-training, SWAN consistently and efficiently converges without learning rate warm-up (Section 6.1, Section 6.4).
+    2. 2.
+      Memory efficiency: SWAN achieves the memory footprint of SGD, that is $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states when compared to Adam.
+    3. 3.
+      Computational efficiency: consistently achieves the same or better performance when compared to Adam across several scales and can significantly improve token-efficiency during training up to 2$\times$ at the 1B scale for the LLaMA model.
+  • Theoretical analysis of SWAN.
+We show that (1) $\mathtt{GradNorm}$ can stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix B); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 5.3).
+Additionally, we highlight that in the quadratic case, $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 2, Section B.1).
+  • Fast convergence for LLM pretraining.
+Through experiments (Section 6) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also achieves similar or even better performance compared to Adam.
+Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
 (a) 350 M LLaMA model Pretraining
@@
 (c): memory footprint.
-Unlike Zhao et al.
-(2024a), we directly measure end-to-end memory under full-model training, with batch size = 1 sequence.
+We directly measure end-to-end memory under full-model training, with batch size = 1 sequence.
+This is different to the layer-wise training setting Zhao et al.
+(2024a).
 SWAN achieves near-SGD optimizer memory reduction ($\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states).
 (d) Training throughput analysis on training 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens.
-We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the efficiency of optimizer relative to Adam.
+We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the token efficiency of optimizer relative to Adam.
 SWAN closely reproduces the absolute throughput of Adam, while improving the effective throughput.
-Our contributions are as follows:
-[…]
-    1. 1.
-      Robustness to ill-conditioned problems. On LLM pretraining, SWAN converges quickly even without learning rate warm-ups (Section 5.3, Section 5.6), a property not observed in Adam.
-    2. 2.
-      Memory efficiency: SWAN achieves the memory footprint of SGD, that is $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states when compared to Adam.
-    3. 3.
-      Computational efficiency: consistently outperforms Adam and reaches up to 2$\times$ acceleration and effective throughput.
-  • Theoretical analysis of SWAN.
-We show that (1) $\mathtt{GradNorm}$ can be derived as a strategy to stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix B); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 4.3).
-Additionally, we highlight that $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 3, Section B.3), offering theoretical advantages over SGD and Adam (Proposition 3).
-  • Fast convergence for LLM pretraining tasks.
-Through experiments (Section 5) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also outperforms Adam and other memory-efficient optimizers.
-Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
+2 Related works
+Towards Stateless Optimizers for LLM Training.
+Adaptive optimizers generally rely on tracking internal state variables to perform weight updates, which can substantially increase memory consumption when training large models.
+Several recent works have successfully managed to reduce the memory requirements associated with storing additional state variables for training LLMs.
+Muon (Jordan et al., 2024; Bernstein & Newhouse, 2024b), a newly proposed optimizer, has demonstrated strong acceleration and memory saving for LLM training by simplifying shampoo-like optimizers (Gupta et al., 2018; Anil et al., 2021; Shi et al., 2023; Wang et al., 2024; Vyas et al., 2024; Peirson et al., 2022; Lin et al., 2024) and requiring only the tracking of a first-moment estimate.
+SGD-Sal (Xu et al., 2024b) only stores the first moment with a learning rate estimated at the beginning of training.
+Sign-based methods such as (Chen et al., 2023) have also demonstrated success on training transformer-based models by only tracking first moments.
+There are also several works that aim to enhance the memory efficiency of Adam by reducing the memory cost associated with second moments.
+Adam-mini (Zhang et al., 2024b) significantly reduces memory usage by storing only scalar values for each parameter block, while Adalayer (Zhao et al., 2024c) retains only the scalar average of the second moment for each layer.
+Alternatively, PSGD (Pooladzandi & Li, 2024; Li, 2017) focuses on exclusively tracking a pre-conditioner, eliminating the need to track a first moment estimate.
+Finally,  (Zhu et al., 2024) proposes an nearly stateless optimizer to train LLMs, which only requires tracking rank-1 first and second moments.
+However, all the aforementioned optimizers still require the storage of state variables.
+In contrast, SWAN completely eliminates the need to store internal states for both the first and second moments by employing a combination of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ steps, which is discussed next.
+Pre-processing Gradients.
+Gradient pre-processing is a common technique used to enhance performance of optimizers.
+Various pre-processing procedures have been proposed in the literature, such as signed gradient (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023; Kunstner et al., 2023), gradient clipping (Zhang et al., 2020), normalization (Zhang et al., 2020; You et al., 2019), and whitening (Yang & Laaksonen, 2008; Kingma & Ba, 2015; Hwang, 2024; Jordan et al., 2024; Bernstein & Newhouse, 2024c; a; Carlson et al., 2015).
+In this work, we particularly focus on normalization and whitening.
+We apply normalization row-wise on gradient matrices, similar to (You et al., 2019); together with gradient whitening under a specific structural assumption of the Fisher Information (Hwang, 2024; Martens et al., 2018), recovering the orthogonalization step used in (Jordan et al., 2024; Tuddenham et al., 2022).
+Our key result is that composing normalization and whitening on stochastic gradients is sufficient to enable the efficient training of LLMs in a completely stateless manner.
+Compared to Lamb (You et al., 2019), our normalization operation is applied on raw gradients instead of Adam states.
+Compared to Muon (Jordan et al., 2024), SWAN removes first-moment tracking and instead uses normalization.
+We show empirically that removing any one of these two pre-processing steps from SWAN results in significant performance degradation (Section 6.4).
+Low-rank methods.
+Low-rank optimization techniques have been explored in the context of large language model (LLM) training as a means to reduce memory consumption.
+These methods focus on applying low-rank approximations to model weights, gradients, and/or optimizer state variables.
+A seminal work in this domain is LORA (Hu et al., 2021) to fine-tune pre-trained models using additional low-rank weight matrices at each layer, thereby significantly reducing memory usage to update the weights.
+More recently, methods such as ReLoRA (Lialin et al., 2023), FLORA (Hao et al., 2024), and Galore (Zhao et al., 2024a) have advanced low-rank optimization techniques for memory-efficient LLM pre-training.
+These approaches leverage low-rank gradient projections to enable full-rank learning, thereby achieving memory savings without compromising model capacity.
+Notably, they have achieved substantial reductions in the memory consumption of optimizer states, with only minimal impact on model performance.
+While these approaches effectively reduce the memory footprint of LLM training, they still necessitate storing internal states, resulting in higher memory consumption compared to SWAN.
 3 Preliminaries
-One of the most used adaptive optimizers is the Adam optimizer (Kingma & Ba, 2015).
-Adam is an example of stateful optimizers, which involves tracking and maintaining internal states.
-It combines the advantages of two earlier methods: AdaGrad (Duchi et al., 2011), which adapts learning rates based on the historical gradient magnitude, and RMSProp (Tieleman, 2012), which mitigates the aggressive decrease in learning rates by using a decaying average of squared gradients.
+Adam (Kingma & Ba, 2015) is the current standard choice for adaptive optimizers across a multitude of machine learning training tasks, including LLM pre-training.
+Adam is an example of a stateful optimizer, which means it accumulates and stores internal states throughout training.
+It combines the advantages of two earlier methods: AdaGrad (Duchi et al., 2011), which adapts learning rates based on the historical gradients’ magnitudes, and RMSProp (Tieleman, 2012), which mitigates the aggressive decrease in learning rates by using a decaying average of squared gradients.
 Consider a loss function ${\mathcal{L}}_{{\mathbf{W}}}:{\mathcal{X}}\rightarrow\mathbb{R}$, parameterized by weight matrices ${\mathbf{W}}\in\mathbb{R}^{m\times n}$, and denote ${\bm{x}}^{(t)}$ a mini-batch of inputs provided at the $t$-th training step that is sampled from data distribution $p_{\text{data}}({\bm{x}})$.
@@
 where ${\bm{m}}^{(t)}$ and ${\bm{\nu}}^{(t)}$ are EMAs of the first and second moments of the gradients; and $\eta$ is a global step size.
-Essentially, Adam estimates the signal-to-noise (SNR) ratios and use it to adjust learning rates element-wise.
+Intuitively, Adam estimates the signal-to-noise (SNR) ratios and use it to adjust learning rates element-wise.
 Tracking and storing these two EMA estimates triple the total memory consumption required to train a LLM model.
@@
 There is a rich literature on understanding adaptive methods’ inner workings and unreasonable effectiveness.
-Notably, the key desired properties of Adam can be categorized as gradient smoothing, gradient invariance and gradient whitening.
-(1) Gradient smoothing: the EMA operations in Adam naturally reduce the influence of mini-batch noise Cutkosky & Mehta (2020); Crawshaw et al.
-(2022); (2) Gradient invariance: recent work suggest the performance gap between SGD and Adam might lie in Adam’s sign-descent-like nature (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023).
-For example, Adam is robust to the rescaling of gradient diagonals (Kingma & Ba, 2015); and is invariant to any sign-preserving scalings (under $\beta_{1}=\beta_{2}=0$)(Bernstein et al., 2018).
-And (3) Gradient whitening : it is well-known that the second moment EMA of Adam approximates the diagonal of Fisher information matrix (Kingma & Ba, 2015; Hwang, 2024), and biases the optimization trajectories towards well-conditioned regions (Jiang et al., 2024).
+Notably, the key desired properties of Adam can be categorized as gradient whitening, gradient smoothing, and gradient invariance.
+  • Gradient whitening: it is known that the inverse second moment $\frac{1}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}$ of Adam performs gradient whitening by approximating the square root inverse of the diagonal of Fisher information matrix (Kingma & Ba, 2015; Hwang, 2024).
+This step biases the optimization trajectories towards well-conditioned regions (Jiang et al., 2024) and provides a better approximation to the geodesic flow when compared with the natural gradient update (Yang & Laaksonen, 2008).
+  • Gradient smoothing: the EMA operations in Adam naturally reduce the influence of mini-batch noise (Cutkosky & Mehta, 2020; Crawshaw et al., 2022);
+  • Gradient invariance: recent work suggest the performance gap between SGD and Adam might lie in Adam’s sign-descent-like nature (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023).
+For example, Adam is robust to the rescaling of gradient diagonals (Kingma & Ba, 2015); and is invariant to any sign-preserving scalings (under $\beta_{1}=\beta_{2}=0$(Bernstein et al., 2018).
 For a more comprehensive discussion of these properties, please refer to appendix A.
-Adam as SGD preprocessing, and the potential unnecessity of EMAs
-Adam can be viewed as history-dependent preprocessing of the gradients ($\{{\mathbf{G}}^{(0)},{\mathbf{G}}^{(1)},...,{\mathbf{G}}^{(t-1)},{\mathbf{G}}^{(t)}\}\rightarrow\frac{\hat{{\bm{m}}}^{(t)}}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}$) to achieve the desired properties above.
-A key observation is, all of the above properties are achieved through the element-wise operations.
-That is, each element of the gradient matrix is pre-processed and re-scaled independently.
-This ignores the interactions and structures among different variables, and we hypothesize this is the reason why extra history information is need to compensate this gap, leading to the requirement of EMA states.
-Thus, we believe that designing stateless adaptive optimizers is possible, if we can achieve similar properties by designing new matrix-level operations that are preprocesses the instantaneous gradients of SGD (${\mathbf{G}}^{(t)}\rightarrow\tilde{{\mathbf{G}}}^{(t)}$).
+Adam as SGD pre-processing.
+Adam can be viewed as history-dependent pre-processing of the gradients ($\{{\mathbf{G}}^{(0)},{\mathbf{G}}^{(1)},...,{\mathbf{G}}^{(t)}\}\rightarrow\frac{\hat{{\bm{m}}}^{(t)}}{\sqrt{\hat{{\bm{\nu}}}^{(t)}}+{\epsilon}}$) to achieve the desired properties described above.
+A key observation is that all of these properties are achieved through element-wise operations, where each element of the gradient matrix is independently pre-processed and re-scaled.
+This approach does not take into account the interactions and structures between different variables, and we hypothesize that this is the reason why additional history information is necessary to bridge this gap, ultimately leading to the requirement for EMA states.
+Thus, we believe that designing stateless adaptive optimizers is possible if we can achieve similar properties by applying matrix-level operations that pre-process the instantaneous stochastic gradients of SGD (${\mathbf{G}}^{(t)}\rightarrow\tilde{{\mathbf{G}}}^{(t)}$).
 4 The SWAN Optimizer: Preprocessing SGD with Normalization and Whitening
@@
 In $\mathtt{GradNorm}$ operator, we perform standardization across the output dimensions (columns), using statistics computed row-wise.
-In $\mathtt{GradWhitening}$ operator, we treat each column (or row) of the gradient matrix ${\mathbf{G}}$ as a separate data sample.
-Then, the $\mathtt{GradWhitening}$ operator can be seen as stretching/squeezing the gradient data along all eigen directions, such that the covariance matrix is an identity matrix.
-Note that the same approach can be applied row-wise as well.
+In $\mathtt{GradWhitening}$ operator (illustration adapted from Huang et al.
+(2019)), we treat each column of the gradient matrix ${\mathbf{G}}$ as a separate data sample.
+Then, $\mathtt{GradWhitening}$ can be seen as stretching/squeezing the gradient data along all eigen directions, such that the covariance matrix is the identity matrix.
 Algorithm 1 SWAN Optimizer
-  Input: weight matrix ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ with $m\leq n$. Step size $\eta$. Number of $\mathtt{GradWhitening}$ iteration $k$.
+  Input: weight matrix ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ with $m\leq n$. Step size $\eta$. Number of $\mathtt{GradWhitening}$ iteration $k$ (default = 10).
   Initialize step $t\leftarrow 0$
@@
 Algorithm 2 $\mathtt{GradWhitening}$ Operator
-  Input: Matrix ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of Newton Schultz iteration $k$, step size $\beta$
+  Input: ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of iterations $k$. Step size $\beta$ (default = 0.8).
   Initialize ${\mathbf{Y}}\leftarrow{\mathbf{G}}{\mathbf{G}}^{\top}$, ${\mathbf{Z}}\leftarrow{\mathbf{I}}$
@@
 Algorithm 3 $\mathtt{GradNorm}$ Operator
-  Input: Matrix ${\mathbf{G}}^{m\times n}$
-  return  $\frac{{\mathbf{G}}-\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}\mathbf{1}_{n}^{\top}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\sum_{j=1}^{n}{\mathbf{G}}_{:,j})^{2}}\mathbf{1}_{n}^{\top}}$
+  Input: ${\mathbf{G}}^{m\times n}$
+  return  $\frac{{\mathbf{G}}-\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}\mathbf{1}_{n}^{\top}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\frac{1}{n}\sum_{k=1}^{n}{\mathbf{G}}_{:,k})^{2}}\mathbf{1}_{n}^{\top}}$
 Figure 3: SWAN Optimizer.
-Note that for $\mathtt{GradWhitening}$, in principle, any approximate algorithms can be applied, such as matrix Pade approximant.
+Note that for $\mathtt{GradWhitening}$, in principle, any approximate algorithms can be applied, such as matrix Pade approximate Song et al.
+(2022).
 Here we only show one particular working example based on Newton-Schulz, which introduces minimal overheads.
-As discussed in Section 2, we believe the key to designing stateless, adaptive, and effective optimizers lies in designing new matrix-level operations that exploit rich information in the gradient matrix.
-To this end, we propose a new stateless optimizer (Algorithm 1), SWAN (SGD with Whitening And Normalization).
-SWAN preprocesses SGD based on two simple operators, $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$, both are stateless, matrix-level operators.
+As discussed in Section 3, we believe the key to designing stateless, adaptive, and effective optimizers lies in the incorporation of matrix-level operations that exploit rich information contained in the gradient matrix.
+To this end, we compose two well-known matrix operators, namely normalization and whitening.
 When applied in tandem, they achieve similar desirable properties of adaptive optimizers, without the need to store historical gradient moments.
-In this section we describe the update rules of SWAN.
+The result is SWAN (SGD with Whitening And Normalization), a new stateless optimizer which we describe next.
 4.1 SWAN Update Rules
-In SWAN (Algorithm 1), the raw SGD gradient ${\mathbf{G}}^{t}$ is processed by the following operations  [^1] Here ${\mathbf{G}}$ is assumed to be the gradient matrix for some parameter block in the model (e.g.
+In SWAN (Algorithm 1), the raw SGD gradient ${\mathbf{G}}^{t}$ is processed by the operations below [^1] Here ${\mathbf{G}}$ is assumed to be the gradient matrix for some parameter block in the model (e.g.
 a linear layer); enabling us to take advantage of the matrix structure to achieve our smoothing and whitening goals.:
@@
 […]
-$\mathtt{GradNorm}$: gradient invariance & stabilization.
+On the $\mathtt{GradNorm}$ Step.
 Consider the gradient matrix ${\mathbf{G}}\in\mathbb{R}^{m\times n}$ with rows and columns corresponding to input and output dimensions, respectively, of some block of model parameters.
 Let $1\leq i\leq m$ represent the input indices and $1\leq j\leq n$ represent the output indices.
-Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, we propose to standardization across the output dimensions at each time step $t$, defined as:
-$\mathtt{GradNorm}({\mathbf{G}})=\frac{{\mathbf{G}}-\bar{{\bm{g}}}\mathbf{1}_{n}^{\top}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
-where $\bar{{\bm{g}}}=\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}$ is the mean across output dimensions at time $t$; ${\bm{s}}=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\bar{{\bm{g}}})^{2}}$ is the standard deviation across dimensions at time $t$; $\bar{{\bm{g}}}$ and ${\bm{s}}$ are ${m}$-dimensional column vectors; and $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
-$\mathtt{GradNorm}$ is an example of operators typically used in the forward pass applied to the backward pass: $\mathtt{GradNorm}$ is LayerNorm (Ba et al., 2016) (without learnable parameters) being applied to the backward gradient rather than forward activations.
-Similar to Layer Norm, $\mathtt{GradNorm}$ allows the optimizer to be invariant under the following transformations on gradient matrix: i) matrix-wise re-scaling; ii) matrix-wise re-centering; iii) row-wise re-scaling; and iv) row-wise re-centering.
+Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, as done in Adam, we propose to standardize across the output dimensions at each time step $t$, as follows:
+$\mathtt{GradNorm}({\mathbf{G}}):=\frac{{\mathbf{G}}-\bar{{\bm{g}}}\mathbf{1}_{n}^{\top}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
+where $\bar{{\bm{g}}}:=\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}$ is the mean across output dimension; ${\bm{s}}:=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\bar{{\bm{g}}})^{2}}$ is the standard deviation across dimension; $\bar{{\bm{g}}}$ and ${\bm{s}}$ are ${m}$-dimensional column vectors; $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
+$\mathtt{GradNorm}$ is the forward pass operator LayerNorm (Ba et al., 2016) (without learnable parameters) applied on backward gradients.
+Similar operations have been considered in the literature to process gradients (You et al., 2019).
+$\mathtt{GradNorm}$ allows the optimizer to be invariant under the following transformations on the gradient matrix: i) matrix-wise re-scaling; ii) matrix-wise re-centering; iii) row-wise re-scaling; and iv) row-wise re-centering.
 Compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) used in signed descent, $\mathtt{GradNorm}$ preserves richer information of the gradient scaling while offering invariance properties.
-Moreover, in Section 4 we will show that when applied to a transformer architecture, $\mathtt{GradNorm}$ also helps stabilize the gradient distribution across time.
-$\mathtt{GradWhitening}$: gradient whitening.
-Adam performs approximate gradient whitening element-wise, leading to well-conditioned regions.
-In contrast, we propose the matrix $\mathtt{GradWhitening}$ operator, which performs whitening at the matrix level.
-Given the gradient ${\mathbf{G}}$, the $\mathtt{GradWhitening}$ operator simply orthogonalizes ${\mathbf{G}}$  [^2] An equivalent definition is given by $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{U}}{\mathbf{V}}^{\top}$, where ${\mathbf{U}}$ and ${\mathbf{V}}$ are both orthogonal matrices defined via SVD decomposition of ${\mathbf{G}}={\mathbf{U}}{\mathbf{S}}{\mathbf{V}}^{\top}$.
-:
-$\mathtt{GradWhitening}({\mathbf{G}}):=({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ (2)
+In Section 5 we will show that when applied to a simplified transformer architecture, $\mathtt{GradNorm}$ also helps stabilize the gradient distribution.
+On the $\mathtt{GradWhitening}$ Step.
+As discussed in Section 3, Adam relies on a second moment estimate to perform element-wise gradient whitening.
+More formally, the second moment of Adam estimate a diagonal approximation of the Fisher information matrix (FIM) via (Kingma & Ba, 2015; Hwang, 2024):
+$\mathbb{E}(\text{Vec}({\mathbf{G}})\text{Vec}({\mathbf{G}})^{\top})\approx\text{Diag}[\text{Vec}[\mathbb{E}({\mathbf{G}}^{\odot 2})]]$
+where $\text{Vec}(\cdot)$ denotes the vectorized operation, $\text{Diag}(\cdot)$ denotes the operation that produces a diagonal matrix given a vector, and $\odot$ denotes the Hadamard product.
+Hence, whitening using the inverse square root of this diagonal FIM is equivalent to element-wise rescaling with $\frac{1}{\sqrt{\mathbb{E}({\mathbf{G}}^{\odot 2})}}$.
+Here, we consider the following non-diagonal approximation:
+$\mathbb{E}(\text{Vec}({\mathbf{G}})\text{Vec}({\mathbf{G}})^{\top})\approx{\mathbf{I}}_{n\times n}\otimes{\mathbf{G}}{\mathbf{G}}^{\top}\;,$
+where $\otimes$ denotes Kronecker product. This leads to the following whitening step:
+$\mathtt{GradWhitening}({\mathbf{G}}):=({\mathbf{G}}{\mathbf{G}}^{\top})^{-1/2}{\mathbf{G}}$ (2)
 where the exponent $-\frac{1}{2}$ stands for matrix inverse square root.
-One can show that $({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ is simply the orthogonalization of ${\mathbf{G}}$, i.e., the closest orthogonal matrix to ${\mathbf{G}}$ (w.r.t the Frobenius norm).
+$({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ is simply the orthogonalization of ${\mathbf{G}}$, i.e., the closest orthogonal matrix to ${\mathbf{G}}$ (w.r.t the Frobenius norm).
+The derivation of Equation 2 as structured FIM/Hessian approximation is discussed in Section 5.3.
 […]
 4.2 Practical Considerations
-Various GPU-friendly and fast approximations are developed in the literature of decorrelated batch normalization or feature whitening.
-We consider the naive Netwon-Schulz (N-S) iteration algorithm (Li et al., 2018; Huang et al., 2019) to compute $\mathtt{GradWhitening}$:
+Computing $\mathtt{GradWhitening}$ exactly can be expensive, as it involves solving the matrix square-root inverse.
+We hence directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation.
+This is given by (Song et al., 2022; Li et al., 2018):
 ${\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}),\quad{\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}$ (3)
-[…]
+where ${\mathbf{Y}}_{0}={\mathbf{G}}{\mathbf{G}}^{\top}$ , ${\mathbf{Z}}_{0}={\mathbf{I}}$.
+At convergence, $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{Z}}{\mathbf{G}}$ (Algorithm 2).
+See Appendix H for implementation details.
+Note that the same N-S procedure has been discussed in (Mei et al., 2023) and  (Jackson, 2023) for preconditioned optimizers.
+Another N-S procedure has also been introduced in (Bernstein & Newhouse, 2024b), and optimized in (Jordan et al., 2024) under BF16 precision.
 Computational Overhead.
@@
 Rescaling $\mathtt{GradWhitening}$.
-The whitening operation $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, hence might drastically change the effective learning rate.
+The whitening operation $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, and as such might drastically change the effective learning rate.
 In practice, we propose the following re-scaling before updating the weight:
 $\Delta{\mathbf{W}}^{(t)}\leftarrow\frac{\|\tilde{{\mathbf{G}}}^{(t)}\|\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}$ (4)
-This will helps us to rescale the norm of whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, and allowing us to get reasonably good performance without extensively tuning learning rates–we found that the learning rates used Adam usually also work well for SWAN.
-It is also possible to remove rescaling, but that would require additional hyper-parameter search to find the optimal learning-rate for SWAN.
+This helps rescale the norm of the whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, and allowing us to get reasonably good performance without extensively tuning learning rates; we found that the learning rates used by Adam tend to work well for SWAN.
+We can remove rescaling, but doing so would require additional hyper-parameter search to find the optimal learning rate.
 5 Analysis: A LLM Learning Dynamics Perspective
@@
 However, as a first analysis, we consider SWAN from a learning dynamics perspective, specifically the dynamics of an LLM based upon a simplified transformer block.
-It is this analysis that led to the design of SWAN.
+It is this analysis that led to the design choices for SWAN.
 5.1 Setup
-We assume the following simplified transformer block (STB) architecture recently proposed in Tian et al. (2023):
+We consider the simplified transformer block (STB) architecture proposed in (Tian et al., 2023).
 Definition 1 (Simplified Transformer Block (STB)).
@@
 For example, for each row $i$, $\dot{{\mathbf{W}}}^{(t)}[i,:]$ can be re-written as $\dot{{\mathbf{W}}}^{(t)}[i,:]=\mathbb{E}[\dot{{\mathbf{W}}}^{(t)}[i,:]]+{\bm{\varepsilon}}_{{\mathbf{W}}}^{(t)}[i,:]$, where ${\bm{\varepsilon}}^{(t)}[i,:]$ is zero mean random variable with covariance $\mathrm{Cov}[\dot{{\mathbf{W}}}^{(t)}[i,:]]$.
-4.2 $\mathtt{GradNorm}$ stabilizes gradient distributions of simplified transformers
-[…]
+5.2 $\mathtt{GradNorm}$: stabilizing gradient distributions of STB
+Below we show that, based on the dynamics of the STB, $\mathtt{GradNorm}$ stabilizes ${\bm{\varepsilon}}_{{\mathbf{W}}}^{(t)}$.
 Theorem 1 ($\mathtt{GradNorm}$ stabilizes gradient distributions across time for the STB).
@@
 Theorem 1 suggests that $\mathtt{GradNorm}$ implicitly aligns with the dynamics of transformer architectures and removes the time-heterogeneity in gradient covariance structures.
-4.3 $\mathtt{GradWhitening}$ Is An Efficient Non-diagonal Second-Order Update
-In this section, we show that $\mathtt{GradWhitening}$ is equivalent to non-diagonal approximate second-order method, under a specific Kronecker factorization assumption of the Hessian.
-The assumption is as below:
+5.3 $\mathtt{GradWhitening}$: An Efficient Non-diagonal Second-Order Update
+Here, we show that $\mathtt{GradWhitening}$ is equivalent to a non-diagonal second-order method under a specific Kronecker factorization assumption of the Hessian/FIM. The assumption is as below:
 Assumption 1 (Assumption of $\mathtt{GradWhitening}$).
 At time $t$, the local Hessian ${\mathbf{H}}$ of the loss has shared block-diagonal structure, such that ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
-By leveraging this structural assumption, we can now effectively estimate ${\mathbf{H}}$ by only using one single gradient matrix sample $\bm{G}:=[{\bm{g}}_{1},\dots,{\bm{g}}_{n}]\in\mathbb{R}^{m\times n}$.
-Recall that the Fisher information formulation of Hessian is defined as ${\mathbf{H}}=\mathbb{E}(\text{Vect}({\mathbf{G}})\text{Vect}({\mathbf{G}})^{\top})$ where $\text{Vect}(\cdot)$ denotes the vectorized operation.
-Under assumption 1, we can estimate ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, using a single sample of the gradient, that is we can compute the following simple estimate $\frac{1}{n}\sum_{i=1}^{n}{\bm{g}}_{i}{\bm{g}}_{i}^{T}=\bm{G}\bm{G}^{\top}$, which approximate the current $\tilde{{\mathbf{H}}}$.
+Approximating Hessian with a Kronecker factorization is not new and has been extensively studied in the literature (Martens & Grosse, 2015; George et al., 2018; Gao et al., 2023; 2021; Koroko et al., 2022; Eschenhagen et al., 2024; Gupta et al., 2018).
+Here, this specific structure is useful in our context as 1) it enables estimation of Hessian/FIM without temporal EMAs, nor mini-batch statistics, and 2) it aligns with the statistical property of STB, as shown later in Proposition 1.
+By leveraging assumption 1, we can now effectively estimate ${\mathbf{H}}$ by only using one single gradient matrix sample $\bm{G}:=[{\bm{g}}_{1},\dots,{\bm{g}}_{n}]\in\mathbb{R}^{m\times n}$.
+Recall that the Fisher information formulation of Hessian is defined as ${\mathbf{H}}=\mathbb{E}[\text{Vec}({\mathbf{G}})\text{Vec}({\mathbf{G}})^{\top}]$ where $\text{Vec}(\cdot)$ denotes the vectorized operation.
+Under assumption 1, we can estimate ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$ by computing the following simple estimate $\frac{1}{n}\sum_{i=1}^{n}{\bm{g}}_{i}{\bm{g}}_{i}^{\top}=\bm{G}\bm{G}^{\top}$, which approximate the current $\tilde{{\mathbf{H}}}$.
 Hence, $\mathtt{GradWhitening}$ can be seen as applying a second order update under our structural assumption:
 $\mathtt{GradWhitening}({\mathbf{G}})=({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}=\tilde{{\mathbf{H}}}^{-\frac{1}{2}}{\mathbf{G}}$
-In Theorem 2 of Appendix, we will further show that the pre-conditioned update rule $\tilde{{\mathbf{H}}}^{-\frac{1}{2}}{\mathbf{G}}$ indeed induces a Newton-like behavior.
-More importantly, in the following Proposition, we show that the implicit assumption of $\mathtt{GradWhitening}$ is a good abstraction of LLM dynamics since it in fact aligns with the equilibrium Hessian structure in the STB regime:
+In the following Proposition, we show that the assumption 1 of $\mathtt{GradWhitening}$ aligns with the equilibrium Hessian structure in the STB regime.
 Proposition 1 (Shared structures in the block-diagonal of Hessians at transformer equilibrium).
@@
 Where ${\mathbf{H}}({\mathbf{W}})_{sk,s^{\prime}k^{\prime}}=\frac{\partial\mathcal{L}}{\partial{w_{sk}}\partial{w_{s^{\prime}k^{\prime}}}}$.
-Proposition 1 shows that, under a simplified setting of the transformer, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares the identical structure, which supports the implicit assumption of $\mathtt{GradWhitening}$.
+Proposition 1 shows that, under a simplified setting of the transformer, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares an identical structure, which supports the assumption of $\mathtt{GradWhitening}$.
 This result is verified in our numerical experiment (Appendix, Figure 9).
-Finally, Theorem 3 presented in Appendix will also how $\mathtt{GradWhitening}$ helps to make the convergence rate of SGD more robust to the condition number of local curvatures, and outperforms both SGD and Adam in the ill-conditioned regime.
+Finally, Theorem 2 presented in Appendix will also show how $\mathtt{GradWhitening}$ helps to make the convergence rate of SGD more robust to the condition number of local curvatures, and outperforms both SGD and Adam in the ill-conditioned regime.
 6 Experiments
-In this section, we perform empirical studies on SWAN. The main goals of those studies are as follows:
-  • Verifying whether $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ indeed effectively fulfilled their design goals (stabilizing gradient noise, and modeling/counteracting local curvatures).
-This is performed in Section 5.1 and Section 5.2.
-  • Verifying whether SWAN can consistently offer Adam-level or even better performance, while maintaining low memory cost and high-throughput.
-This is performed in Section 5.3, Section 5.4 and Section 5.5, respectively.
-  • Ablation study (Section 5.6).
+In this section, we report empirical results for SWAN. The main purposes are:
+  • Benchmarking the performance, training speed-ups and memory footprint of SWAN on pre-training LLaMA with C4 dataset (Raffel et al., 2020) against Adam.
+Results are presented in Section 6.1, Section 6.2, and Section 6.3.
+  • Investigating the contribution of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ to the performance of SWAN through ablation studies, as outlined in Section 6.4.
+  • Separately evaluating whether $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ achieve their intended purposes of stabilizing gradient noise and modeling/ counteracting local curvature, respectively.
+Detailed results are reported in Section 6.5.1 and Section 6.5.2.
 All experiments run on NVIDIA A100 GPUs.
-5.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
-Figure 4: (a) - (c): KL Divergence of gradient distributions across training Steps.
-We use the projection weights in attention and MLP modules of the second layer as an example.
-The plots compare standard training with $\mathtt{GradNorm}$-augmented training.
-Lower KL divergence values indicate greater stability in gradient distributions.
-(d): Training loss curve comparing standard training vs $\mathtt{GradNorm}$-augmented training
-To investigate whether $\mathtt{GradNorm}$ indeed stabilizes stochastic gradient distributions as claimed in Theorem 1, we conduct a series of controlled experiments using a scaled-down version of the LLaMA architecture.
-We aim to assess the impact of $\mathtt{GradNorm}$ on the gradient distribution during training by analyzing the statistical properties of gradients across multiple training steps.
-Specifically, we employ a small-scale LLaMA-based model with approximately 10 million parameters  (Lialin et al., 2023).
-Training is conducted on the C4 dataset (Raffel et al., 2020).
-Baselines
-We compare the following methods:
-  • Standard training: this baseline utilize SGD optimizer with a learning rate of $5\times e^{-4}$, and apply a linear learning rate scheduler with a warm-up period comprising 10% of the total training steps (10,000).
-]
-  • $\mathtt{GradNorm}$-augmented training: apply $\mathtt{GradNorm}$ on noisy gradients before SGD update. We use the same hyperparameters as the SGD baseline.
-Methodology
-For each method, we measure the change of gradient statistics due to mini-batch sampling noise.
-For each training step, after the parameter update, we additionally sample 16 different mini-batches (each with batch size 64), and calculate the mean and standard deviation of the corresponding gradients on each of those batches.
-Then, we estimate the Kullback-Leibler divergence between the current gradient distribution (raw gradient and with $\mathtt{GradNorm}$) relative to the initial distribution at $t=0$.
-This allows us to quantify the temporal evolution of gradient distributions.
-Results
-Figure 4 illustrates the KL divergence of gradient distributions over training steps for both standard training and $\mathtt{GradNorm}$-augmented training, relative to the initial distribution at $t=0$.
-Apart from minor spikes in the beginning of training, $\mathtt{GradNorm}$ consistently and effectively stabilizes the gradient noise distribution.
-Furthermore, as a by-product, the loss curves demonstrate that models trained with $\mathtt{GradNorm}$ performs indeed improves upon standard training.
-Ablation
-One question that we might ask is this: is the performance boost of $\mathtt{GradNorm}$ simply due to that fact that $\mathtt{GradNorm}$ might be increasing the effective learning rate.
-We conduct detailed ablation experiments in Section 5.6, in a larger scale setting, which show this is not the case.
-These findings validate the hypothesis that $\mathtt{GradNorm}$ serves as an effective mechanism for mitigating the adverse effects of mini-batch noise on gradient distributions, thereby enhancing the stability and performance of the training process.
-5.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
-We aim to evaluate the pure optimization capability of $\mathtt{GradWhitening}$ processed gradient descent on classic optimization problems. We consider the following three settings:
-(a) Quadratic optimization, well-conditioned
-(b) Quadratic optimization, ill-conditioned
-(c) Rastrigin function optimization
-(d) Ablation on initialization
-Figure 5: Comparison of convergence rate of different methods on quadratic and non-convex optimization problems.
-(a): 2500-dimensional quadratic optimization with well-conditioned ${\mathbf{H}}$.
-(b): 2500-dimensional quadratic optimization with ill-conditioned ${\mathbf{H}}$ (c): 2500-dimensional Rastrigin function optimization.
-(d): 2500-dimensional Rastrigin function optimization, but forcing all methods to use the same orthogonal initial location.
-  • High-dimensional quadratic optimization. We construct quadratic optimization problem of the form in Equation 23, with ${\mathbf{W}}\in\mathbb{R}^{50\times 50}$.
-  • Ill-conditioned quadratic optimization.
-We construct a quadratic optimization problem with the same dimensionality and formulation as above, but this time we deliberately choose an ill-conditioned ${\mathbf{H}}$.
-  • Non-convex optimization with multiple local optimas. We consider the following multi-variate Rastrigin function:
-    $f({\mathbf{W}})=m^{2}{\mathbf{A}}+\frac{1}{2}\text{Tr}[{\mathbf{W}}^{\top}{\mathbf{W}}]-{\mathbf{A}}\sum_{ij}\text{cos}(2\pi W_{ij})$
-    where ${\mathbf{W}}$ is $m\times m$ matrix.
-This function has $10^{m^{2}}$ possible local optima.
-This is a complicated stress test for optimization algorithms, where many methods are expected to fail at different regimes.
-We set $m=50$, and compare the same methods as in the previous experiment.
-Baselines
-On all three settings (quadratic, quadratic ill-conditioned, and Rastrigin) we compare 5 different methods, including 3 baselines and 2 $\mathtt{GradWhitening}$ based variants.
-For baselines, we have: gradient descent (with theoretical optimal learning rate in Theorem 4); Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$ with hand-tuned learning rate); and Newton’s method with tuned learning rate.
-For 2 $\mathtt{GradWhitening}$-based methods, we consider: $\mathtt{GradWhitening}$-augmented GD with orthogonal initialization (i.e., ${{\mathbf{W}}^{(0)}}^{\top}{\mathbf{W}}^{(0)}={\mathbf{I}}$); and without orthogonal initialization.
-This is to verify whether the claims of Theorem 2 and Theorem 3 is true (i.e., $\mathtt{GradWhitening}$-processed GD behaves like Newton’s method under orthogonal initialization).
-All methods are initialized with the same initial condition (apart from $\mathtt{GradWhitening}$ with orthogonal initialization, which will project the original initialization onto the orthogonal version).
-Results
-Results are shown in Figure 5, (a) - (c). Main findings are summarized as follows:
-  • On quadratic optimization problems (Figure 5 (a) and (b)), $\mathtt{GradWhitening}$ with orthogonal initialization and Newton’s method converges to optimal solution with 1 step of optimization, confirming our theoretical prediction in Theorem 2 and Theorem 3.
-  • On quadratic optimization problems (Figure 5 (a)), when the problem is well-conditioned, standard GD outperforms both Adam and $\mathtt{GradWhitening}$ (non-orthogonal initialization).
-However, when the problem becomes ill conditioned (Figure 5 (b)), $\mathtt{GradWhitening}$ (non-orthogonal initialization) outperforms GD by a large margin; while GD suffers from convergence.
-  • In all three settings (Figure 5, (a) - (c)), $\mathtt{GradWhitening}$ with non-orthogonal initialization consistently outperforms Adam, confirming our theoretical results in Proposition 3.
-  • On Rastrigin function optimization (Figure 5 (c)), $\mathtt{GradWhitening}$ performs competitively to Newton’s method, with or without orthogonal initialization.
-Ablation
-Furthermore, we additionally conduct ablation study on the effect of initialization locations, as shown in fig.
-5 (d).
-In this case, we force all methods to use the same orthogonalized initialization location for optimization.
-As shown by the result, $\mathtt{GradWhitening}$ still consistently outperforms all baselines, eliminating the potential hypothesis that orthogonalized initialization gave $\mathtt{GradWhitening}$ an unfair advantage.
 6.1 SWAN Performance on LLM Pre-training Tasks
 Setup
-Finally, we present results on the task of pre-training large language models.
-Below we consider LLaMa architecture used in Zhao et al.
-(2024a) with RMSNorm and SwiGLU activations.
-We consider sizes of 60M, 130M, 350M and 1.3B parameters on C4 dataset Raffel et al.
-(2020), trained with an effective batch size of 130K tokens (512 batchsize and 256 context length).
-SWAN is applied to all linear modules (in both attention and MLP layers).
-We run all experiments with BF16 format by default.
+We evaluate SWAN on large language model (LLM) pre-training tasks using a LLaMa-based architecture Touvron et al.
+(2023a) with RMSNorm and SwiGLU activations (Zhang & Sennrich, 2019; Gao et al., 2023).
+We consider models with 60M, 130M, 350M, and 1.3B parameters, all trained on the C4 dataset Raffel et al.
+(2020) using an effective token batch size of 130K tokens (batch size 512, context length 256).
+Following the setup of Zhao et al.
+(2024a), SWAN is applied to all linear modules in both attention and MLP blocks.
+Training uses BF16 precision by default.
+The other evaluation settings are the same as Zhao et al.
+(2024a).
 Table 1: Comparison with Adam and its memory-efficient low-rank variants on pre-training various sizes of LLaMA models on C4 dataset.
@@
 $r$ of low-rank methods 128 256 256 512
-Training Tokens 1.1B 2.2B 6.4B 13.1B
+Training Steps 10K 20K 60K 100K
 (a) 60M
@@
 Baselines
-Since the main goal of SWAN is to explore the feasibility of EMA-free adaptive optimizers for LLMs, we mainly compare to Adam and its memory efficient variants.
-We use the exact evaluation settings of Zhao et al.
-(2024a); along with Adam, these include:
-  • Galore Zhao et al. (2024a): the state-of-the-art memory efficient variant of Adam that based on low-rank projections of gradients.
-  • Low-Rank (Kamalakara et al., 2022): Low-rank reparameterization of the weights with the trainable factorization ${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$
-  • LoRA: LoRA Hu et al. (2021) applied for pre-training in the exact setting as used in Zhao et al. (2024a).
-  • ReLoRA Lialin et al. (2023): a full-rank learning extension of Lora with parameter merging.
+Since SWAN aims to explore EMA-free adaptive optimizers for LLMs, we focus on comparing it with Adam and other low-rank optimizers. The baselines are:
+  • Adam (Kingma & Ba, 2015): one standard choice for training large models.
+  • Galore Zhao et al. (2024a): a memory-efficient Adam variant with low-rank gradient projections.
+  • Low-Rank (Kamalakara et al., 2022): a low-rank factorization approach (${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$).
+  • LoRA Hu et al. (2021): the LoRA method applied for pre-training as in Zhao et al. (2024a).
+  • ReLoRA Lialin et al. (2023): a full-rank extension of LoRA with parameter merging.
 Hyperparameters
-For all baselines, we directly apply the same hyperparameter configs (including model configs, optimizers, schedulers) open sourced by Zhao et al.
-(2024a)  [^4] This does not include Adam, since the authors of Zhao et al.
+We use the publicly available hyperparameter configurations from Zhao et al.
+(2024a) for all baselines [^2] This does not include Adam, since the authors of Zhao et al.
 (2024a) did not disclose their Adam configs.
 However, we used the same learning rate sweep procedure as described in their paper.
-This result in an optimal learning rate of 0.001 for 60M, 130M and 350M models, and 0.0007 for 1.3 B models.
-These are also consistent with the optimal learning rates found in some other works, such as Pagliardini et al.
-(2024).
-.
-This includes method-specific fine-tuned optimal learning rates obtained via searching on fixed grid Zhao et al.
-(2024a).
-For our SWAN method, we simply use a hand-picked default hyperparameter setting out-of-the-box, without fine tuning.
-These include a un-tuned, default learning rate of 0.001 across all models, and the 10 steps of Newton-Schultz iteration.
-Notably, thanks to the strong robustness offered by $\mathtt{GradWhitening}$ to ill-conditioned problems, SWAN can safely start training without any learning rate warm-up.
-This is not the case for other baselines such as Adam, in which they failed to converge without long warm-ups in our ablation experiments.
-Apart from the warm-up hyperparameter, all the rest of the parameters in learning rate scheduling used by SWAN is the same as all other baselines.
-Remark on fair comparison
-As empirically shown in Kaddour et al.
-(2024), it is essential to use the same training budget when comparing efficient optimizers; unfortunately, this fact was not widely respected in some recent work of efficient algorithms.
-Kaddour et al.
-(2024) showed that certain state-of-the-art efficient optimizers has shown no gain over Adam under the same training budget.
-Therefore, in our experiments above, we exactly follow the rule of equal training budget (in our case, defined in terms number of tokens), and forces all methods to have learning rate fully decayed to the same targe learning rate at the end of optimization.
-For a even fairer comparison, we left the learning rates of SWAN untuned.
+This result in an optimal learning rate of 0.001 for 60M, 130M and 350M models, and 0.0007 for 1.3 B models., including method-specific learning rates selected through a grid search.
+For SWAN, we use a default setting without extensive tuning.
+We match the global learning rate (0.001) used by Adam, except for the 1.3B model (Adam diverges at 0.001, see Appendix H).
+We use 10 steps of Newton–Schulz iteration for $\mathtt{GradWhitening}$.
+Unlike Adam, SWAN does not require a warm-up phase to ensure stable training.
+All other scheduling parameters remain the same as in the baselines.
+Full details can be found in Appendix H.
+Fair Comparison
+We follow Kaddour et al.
+(2024) by allocating the same total training budget (in terms of processed tokens) for all methods.
+By the end of training, all methods decay to the same target learning rate.
+We also leave SWAN’s learning rates largely untuned for a more direct comparison.
 Results
-Results are shown in Table 1 and Figure 6.
-SWAN outperforms all baselines in both validation PPLs, as well as estimated memory consumption (SWAN theoretically should consume the same memory as SGD).
-Furthermore, to our surprise, on both 350M and 1.3B models, SWAN achieved $>$ 2X speed up against Adam in terms of number of steps (hence number of tokens), see Figure 1 and Figure 7 (a).
-More in-depth discussion of speed-up will be provided in Section 5.4.
+Table 1 and Figure 4 show the results.
+SWAN achieves lower validation perplexities than the baselines with lower estimated memory consumptions.
+On the 350M and 1.3B models, SWAN reaches over a 2$\times$ speedup (in steps or tokens) relative to Adam (Figure 1 and Figure 5 (a)).
+Further details are in Section 6.2.
 6.2 Is the Improvement Multiplicative or Additive?
-One natural question when evaluating speedup factors is whether our method provides a multiplicative or additive speedup over Adam.
-Specifically, multiplicative speedup indicates an optimizer’s performance advantage over a baseline increases proportionally with time, maintaining a consistent speed-up ratio.
-This is the ideal scenario we would like to achieve.
-In contrast, additive speed-up implies only a fixed, constant advantage in steps over baseline, which is less desired.
-To investigate this, we utilize two plots: the speed-up ratio comparison and the perplexity comparison (Figure 7), across different model sizes.
-Before understanding those plots, we first introduce the following concepts:
+A key question in assessing speedup factors is whether the improvement over Adam is multiplicative or additive.
+A multiplicative speedup implies that the optimizer’s relative advantage remains proportionally consistent over time, while an additive speedup suggests a less desired constant step advantage.
+We explore these scenarios using two plots, the speed-up ratio comparison and the perplexity comparison (Figure 5), across different model sizes.
 (a) Speedup factor of SWAN vs Adam training steps
@@
 Speedup Ratio Definition
-The speedup ratio $R(P)$ is defined as the ratio of the number of training steps Adam requires to reach a specific evaluation perplexity (PPL) to the number of steps SWAN requires to achieve the same PPL.
-Mathematically, for a given PPL threshold $P$, the speedup ratio is expressed as:
-$R(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{SWAN{}}}(P)}$ (6)
-where $S_{\text{Adam}}(P)$ and $S_{\text{SWAN{}}}(P)$ denote the training steps taken by Adam and SWAN, respectively, to reach perplexity $P$.
+We define the speedup ratio $R(P)$ for a given perplexity (PPL) threshold $P$ as the ratio of the number of training steps Adam requires to reach a specific evaluation perplexity (PPL) to the number of required steps for SWAN:
+$R(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{SWAN{}}}(P)},$ (6)
+where $S_{\text{Adam}}(P)$ and $S_{\text{SWAN{}}}(P)$ are the training steps required by Adam and SWAN, respectively, to reach perplexity $P$.
 Counterfactual Additive Curve Estimation
-To determine if the speed-up is additive, we estimate a counterfactual additive curve by assuming a constant step advantage $\Delta$ of SWAN over Adam.
-This is calculated using the first 10% - 20% (depending on the number of maximum steps) of the training data:
-$\Delta=\frac{1}{N}\sum_{i=1}^{N}\left(S_{\text{Adam}}(P_{i})-S_{\text{SWAN{}}}(P_{i})\right)$ (7)
-where $N$ is the number of PPL thresholds within the initial 10% - 20% of training steps.
-For the speedup ratio plot:
-$R_{\text{additive}}(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{Adam}}(P)-\Delta}$ (8)
-For the perplexity comparison plot:
-$\text{PPL}_{\text{additive}}(S)=\text{PPL}_{\text{Adam}}(S+\Delta)$ (9)
+To test whether the speedup is additive, we compute a counterfactual additive curve by assuming SWAN gains a fixed step advantage $\Delta$ over Adam in early training (approximately the first 10%–20% of total steps):
+$\Delta=\frac{1}{N}\sum_{i=1}^{N}\left(S_{\text{Adam}}(P_{i})-S_{\text{SWAN{}}}(P_{i})\right),$ (7)
+where $N$ is the number of PPL thresholds considered. We then use $\Delta$ to define the counterfactual additive speedup ratio:
+$R_{\text{additive}}(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{Adam}}(P)-\Delta},$ (8)
+and the counterfactual additive perplexity estimate:
+$\text{PPL}_{\text{additive}}(S)=\text{PPL}_{\text{Adam}}\left(S+\Delta\right).$ (9)
 This represents the expected perplexity of SWAN if it only consistently outperforms Adam by $\Delta$ steps.
 Results
-Results are shown in Figure 7.
-In both subplots, the actual performance of SWAN is compared against the counterfactual additive curves.
-If the actual plot lies significantly above the counterfactual additive curve, it indicates a stronger tendency towards multiplicative speedup, meaning SWAN’s performance gains are more than just a fixed step advantage.
-Analyzing models of sizes 60M, 130M, 350M, and 1.3B parameters, we observe that:
-  • For smaller models (60M and 130M parameters), the actual speedup ratios closely follow the counterfactual additive curves, suggesting that the speedup of SWAN over Adam is primarily additive.
-  • For larger models (350M and 1.3B parameters), the actual plots significantly exceed the counterfactual additive curves, and has a clear sign of accelerated growth.
-This pronounced deviation indicates a multiplicative speedup, demonstrating that SWAN offers increasingly substantial efficiency gains as model complexity scales.
+Figure 5 compares SWAN’s actual performance against the counterfactual additive curves.
+If SWAN’s actual curves exceed these additive estimates, it indicates a tendency towards a multiplicative speedup, instead of the additive advantage.
+We summarize the observations for model sizes of 60M, 130M, 350M, and 1.3B parameters:
+  • For smaller models (60M and 130M), the actual speedup trajectories align closely with the additive baseline, indicating a primarily additive speedup.
+  • For larger models (350M and 1.3B), the actual curves rise noticeably above the additive estimates, suggesting a multiplicative speedup.
+This indicates that SWAN yields increasing efficiency gains as model size grows.
 6.3 Memory Efficiency and Throughput Analysis
-This section, we benchmark the memory footprint of SWAN as well as its throughput.
+In this section, we measure both the memory footprint and throughput of SWAN.
 Memory Footprint
-We benchmark the memory footprint of SWAN, Adam, and Galore in a practical scenario..
+We compare SWAN, Adam, and Galore on a single A100 GPU.
 Unlike Zhao et al.
-(2024a) where memory is benchmarked under a layer-wise training strategy, we directly measure end-to-end memory under full-model training, with batch size = 1 sequence, across different model sizes (1.3B, 7B, and 13B).
-Results are shown in Figure 1 (c), SWAN achieves near-theoretical limit of optimizer memory efficiency of full-parameter training (assuming no quantization).
-This means that SWAN can reproduce the memory footprint of vanilla SGD, reaching $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states.
-This result again highlights the advantage of the stateless design.
+(2024a), which reports layer-wise training memory usage, we measure total end-to-end memory consumption under full-model training using batch size of $1$ for LLaMA with 1.3B, 7B, and 13B parameters.
+As shown in Figure 1 (c), SWAN’s memory usage is close to that of vanilla SGD, providing nearly a 50% reduction in total memory by removing the needs for first and second moments compared to Adam.
+This observation underlines the benefit of a stateless design.
 Effective Throughput
-To measure the throughput of SWAN, we consider training a 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens.
-Here We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the efficiency of optimizer relative to Adam.
-The purpose of the first metric (absolute throughput) is to measure whether SWAN (especially the use of N-S iterations in $\mathtt{GradWhitening}$) create a significant drop in training throughput; and the second metric (effective throughput) takes into account the fact that different optimizers utilizes training tokens at different efficiency level.
-Results are shown in Figure 1 (d), SWAN closely reproduces the absolute throughput of Adam, implying that unlike full SVD used in (Zhao et al., 2024a), the computational overhead of $\mathtt{GradWhitening}$ is minimal.
-As a result, SWAN has doubled the effective throughput vs Adam, due to its high token utilization efficiency as shown in Section 5.3.
-All results are obtained using $\mathtt{GradWhitening}$ in BF16 format.
+We also assess throughput when training a 1.3B-parameter model on 4 A100 GPUs with a token batch size of 130K tokens. We use two metrics:
+  • Absolute throughput: number of tokens processed per second.
+  • Effective throughput: absolute throughput adjusted by SWAN’s token efficiency relative to Adam.
+These metrics evaluate whether $\mathtt{GradWhitening}$ (which uses Newton-Schulz iterations) significantly reduces training speed and also account for the fact that some optimizers make more effective use of training tokens.
+As shown in Figure 1 (d), SWAN achieves an absolute throughput comparable to Adam, indicating that the computational overhead of $\mathtt{GradWhitening}$ is minimal.
+Consequently, SWAN exhibits a higher effective throughput than Adam because of its more efficient use of tokens during training.
+All experiments use BF16 precision for $\mathtt{GradWhitening}$.
 6.4 Ablation Studies
-How does $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ contribute to the performance boost?
-We consider the following 6 ablation settings: SWAN(full), SWAN ($\mathtt{GradNorm}$ only), SWAN ($\mathtt{GradWhitening}$ only), Adam (full), Adam (momentum only), and Adam (second moment only).
-Results are shown in Figure 8 (a).
-It suggests that both $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ makes significant contribution towards the final performance of SWAN; and removing any of them results in suboptimal performance.
-This is similar to how the first and second moment EMAs of Adam jointly contributes to the final performance.
+Effect of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ on Performance
+We consider six ablation settings: (1) SWAN(full), (2) SWAN ($\mathtt{GradNorm}$ only), (3) SWAN ($\mathtt{GradWhitening}$ only), (4) Adam (full), (5) Adam (momentum only), and (6) Adam (second moment only).
+As shown in Figure 6 (a), both $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ contribute to SWAN’s final performance.
+Removing either results in performance degradation.
+Similarly, Adam also requires first and second moment estimates for optimal performance.
 Figure 6: Ablation studies on 130M model.
@@
 (c) Ablation on the effect of learning rate warm-ups.
-Is it true that SWAN works only because its $\mathtt{GradNorm}$ component might have increased the effective learning rates?
-To answer this question, we remove the $\mathtt{GradNorm}$ component of SWAN, and perform learning rate sweep for it.
-We start from the default base learning rate used by full SWAN, and then scale it with multipliers ranging from $1$ to $10^{3}$.
-Results are shown in Figure 8 (b), where the final validation ppl of SWAN without $\mathtt{GradNorm}$ under different learning rate multipliers are shown in blue line; and the performance of the full SWAN under default learning rate is shown as a dashed red reference line.
-The result suggests that, although increasing the effective learning rates might improve the performance of SWAN without $\mathtt{GradNorm}$, there is still a significant cap between between SWAN without $\mathtt{GradNorm}$ and the full SWAN.
-Together with results in Section 5.1, we can conclude that the gradient noise stabilization of $\mathtt{GradNorm}$ is indeed essential, and it cannot be trivialized as larger effective learning rates.
+Does SWAN Succeed Only by Increasing Effective Learning Rates though $\mathtt{GradNorm}$?
+To answer this question, we remove the $\mathtt{GradNorm}$ operator from SWAN and run a learning rate sweep.
+Starting with the default learning rate of full SWAN, we apply multipliers from $1$ to $10^{3}$.
+In Figure 6 (b), the blue line shows the final validation perplexity for SWAN without $\mathtt{GradNorm}$ at different learning rates multipliers, while the dashed red line represents the performance of full SWAN at the default learning rate.
+Although raising the learning rate can improve the performance of SWAN without $\mathtt{GradNorm}$, the gap to full SWAN remains large.
+This indicates that $\mathtt{GradNorm}$’s gradient noise stabilization is essential and cannot be replaced simply by increasing the learning rate.
 How does warm-up affect the performance?
-In Section 5.3, we showed that SWAN does not need any learning rate warm-ups and can train with large learning rates (the default 0.001 learning rate used by all methods is considerably large in LLMs settings).
-Here we perform ablation and compare Adam and SWAN both under warm-ups and not warm-ups.
-As suggested by the results in Figure 8 (c), we found that for SWAN, training without learning rate warm-ups indeed gives better performance.
-Meanwhile, SWAN still outperforms Adam (with warm-up) under the same warm-up schedule of Adam.
-On the contrary, Adam only works with proper learning rate warmup.
-If we remove warmups, the performance of Adam becomes drastically worse.
-This confirms our claim that SWAN is more robust to the local geometries of the optimization problems and can work without (and with) learning rate warm ups.
-6 Related works
-Low-rank memory efficient methods
-Low-rank memory-efficient optimizers have become pivotal in training large language models (LLMs) by reducing memory consumption while maintaining model performance.
-A foundational approach in this domain is LORA introduced by Hu et al.
-(2021), which fine-tunes pre-trained models using low-rank adaptors for each layer, thereby minimizing the memory footprint.
-This leads a full line of low-rank adaptation research, but they are mostly focused on post-training adaptations.
-More recently, ReLoRA Lialin et al.
-(2023),FLORA Hao et al.
-(2024) and Galore Zhao et al.
-(2024a) further push the boundaries of memory-efficient pre-training for LLMs, by introducing the idea of using low-rank projection of gradients to achieve full-rank learning.
-They have demonstrated significant memory savings on optimizer states, with relatively low impact on model performance.
-However, these methods often require 30–50 % of the full Adam’s states and may underperform in terms of convergence speed and final model quality (Zhao et al., 2024a).
-Moreover, it is still an open question on the design choices of these methods, for example: how and why to choose low-rank subspace and projectors?
-Are these projections properly used, given the non-exchangeability nature between low-rank projection and Adam EMA operator?
-Why spikes in loss curves are observed during subspace are shifted/mixed (Chen et al., 2024)?
-These unaddressed questions might make the behavior of low-rank gradient optimizers unpredictable in practice.
-Non-adaptive methods with partial EMA removal.
-Lion (Chen et al., 2023) and Signum (Bernstein et al., 2018) have introduced effective ways to simplify optimization by partially removing the exponential moving average (EMA) of the second moment.
-These methods reduce memory and computational overhead while retaining the benefits of momentum through the first moment EMA, offering a practical trade-off between simplicity and performance.
-However, their reliance on momentum inherently limits their extension to further remove both first and second moment EMAs are removed, as both would reduce to SignSGD (Bernstein et al., 2018), which lacks the adaptive learning rate capability of Adam.
-This might explain why sign-based methods have limited success in LLM pre-training tasks (Chen et al., 2023; Liu et al., 2023).
-On the contrary, our approach completely removes all EMA terms while maintaining the adaptive learning rate nature, which is the key to the success of modern adaptive optimizers.
+Section 6.1 shows that SWAN can train with no warm-up phase even under a relatively large learning rate (0.001).
+Here, we compare Adam and SWAN with and without warm-ups.
+As seen in Figure 6 (c), SWAN without the warm-up phase gives better performance, and it still outperforms Adam under Adam’s own warm-up schedule.
+On the other hand, Adam’s performance decreases drastically without a proper warm-up.
+These findings suggest that SWAN is more robust to warm-up schedule and can train effectively with or without it.
+6.5 Analysis of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
+6.5.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
+Figure 7: KL divergence comparison of gradient distributions against initial gradient distribution across training Steps.
+We use the projection weights in attention and MLP modules of the second layer as an example.
+The plots compare standard training with $\mathtt{GradNorm}$-augmented training.
+Lower KL divergence values indicate greater stability in gradient distributions.
+To examine whether $\mathtt{GradNorm}$ stabilizes the distribution of the stochastic gradients as suggested by Theorem 1, we conduct controlled experiments using a scaled-down LLaMA-based model (about 10 million parameters) (Lialin et al., 2023), trained on the C4 dataset.
+Our goal is to measure how $\mathtt{GradNorm}$ affects the distribution of stochastic gradients over multiple training steps.
+Specifically, we employ a small-scale LLaMA-based model with approximately 10 million parameters  (Lialin et al., 2023).
+Training is conducted on the C4 dataset (Raffel et al., 2020).
+Baselines
+We compare:
+  • Standard training: This uses an SGD optimizer with a learning rate of $5\times 10^{-4}$ and a linear learning rate scheduler, including a 10% warm-up of total training steps (10,000 steps).
+  • $\mathtt{GradNorm}$-processed training: This applies $\mathtt{GradNorm}$ to pre-process the stochastic gradient before the parameter update.
+All other settings match the standard training baseline.
+Methodology
+We measure gradient statistics in the presence of mini-batch noise.
+At step $t=0$, we sample 16 additional mini-batches (batch size 64 each) and compute the mean and standard deviation of the corresponding raw or $\mathtt{GradNorm}$ gradients in each batch, and obtain the approximated initial gradient distribution.
+After each training step of baseline methods, we perform the same procedure and calculate the Kullback-Leibler divergence between the resulting gradient distributions and the initial gradient distributions.
+This process tracks how the gradient distribution changes over time.
+Results
+Figure 7 shows the KL divergence of gradient distributions for standard and $\mathtt{GradNorm}$-augmented training, relative to the corresponding initial approximated distributions.
+Apart from early spikes, $\mathtt{GradNorm}$ reduces fluctuations in the gradient distribution throughout training.
+6.5.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
+This subsection evaluates the optimization performance of gradient descent when combined with $\mathtt{GradWhitening}$. We use three classic problem settings:
+(a) Quadratic optimization, well-conditioned
+(b) Quadratic optimization, ill-conditioned
+(c) Rastrigin function optimization
+(d) Effect of initialization
+Figure 8: Comparison of convergence rate of different methods on quadratic and non-convex optimization problems.
+(a): 2500-dimensional quadratic optimization with well-conditioned ${\mathbf{H}}$.
+(b): 2500-dimensional quadratic optimization with ill-conditioned ${\mathbf{H}}$ (c): 2500-dimensional Rastrigin function optimization.
+(d): 2500-dimensional Rastrigin function optimization, but forcing all methods to use the same orthogonal initial location.
+  • High-dimensional quadratic optimization. A quadratic problem of the form in Equation 10, where ${\mathbf{W}}\in\mathbb{R}^{50\times 50}$.
+  • Ill-conditioned quadratic optimization. Same setup as above, but with a deliberately chosen ill-conditioned ${\mathbf{H}}$.
+  • Non-convex optimization with multiple local optima. We use the multivariate Rastrigin function:
+    $f({\mathbf{W}})=m^{2}{\mathbf{A}}+\frac{1}{2}\text{Tr}[{\mathbf{W}}^{\top}{\mathbf{W}}]-{\mathbf{A}}\sum_{ij}\text{cos}(2\pi W_{ij}),$
+    where ${\mathbf{W}}$ is an $m\times m$ matrix and $m=50$. This function has $10^{m^{2}}$ possible local optima.
+Baselines
+We compare five methods on all three problems: gradient descent (GD) with the theoretical optimal learning rate in Theorem 3, Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and a hand-tuned learning rate, Newton’s method with a tuned learning rate, and two $\mathtt{GradWhitening}$-based variants (with and without orthogonal initialization).
+This is to verify, under orthogonal initialization, $\mathtt{GradWhitening}$-processed GD behaves similarly to Newton’s method, as discussed in Section 5.3 and Theorem 2.
+All methods share the same initialization, except for the orthogonal $\mathtt{GradWhitening}$ variant, which projects the initial parameters onto an orthogonal matrix.
+Results
+From Figure 8 (a)–(c), we summarizes the following outcomes:
+  • Quadratic problems (Figure 8 (a) and (b)).
+$\mathtt{GradWhitening}$ with orthogonal initialization and Newton’s method converge to optimum in one step, aligning with the theoretical predictions in Section 5.3 and Theorem 2.
+  • Well- vs.
+ill-conditioned cases (Figure 8 (a) and (b)).
+In the well-conditioned setting (a), standard GD outperforms both Adam and $\mathtt{GradWhitening}$ (non-orthogonal initialization).
+In the ill-conditioned setting (b), $\mathtt{GradWhitening}$ (non-orthogonal initialization) outperforms GD by a large margin, while GD experiences slow convergence.
+  • Comparison with Adam (Figure 8 (a)–(c)). In all three settings, $\mathtt{GradWhitening}$ with non-orthogonal initialization consistently outperforms Adam, consistent with Proposition 3.
+  • Rastrigin function (Figure 8 (c)). On this non-convex problem, $\mathtt{GradWhitening}$ performs comparably to Newton’s method, with or without orthogonal initialization.
+  • Effect of initialization (Figure 8 (d)) Furthermore, we force all methods to share the same orthogonalized initialization.
+As shown by the result, $\mathtt{GradWhitening}$ GD with orthogonal initialization still consistently outperforms all baselines, confirming that this initialization is only beneficial to $\mathtt{GradWhitening}$ GD among all baselines.
 7 Conclusion
-The most important insight that we conclude for the paper is, preprocessing SGD might already be enough for LLM training tasks.
-This presents an elegant solution to address the memory-performance dilemma of training LLMs using adaptive optimizers like Adam.
-The other insight comes from the fact that the key components of the SWAN optimizer were designed using key insights from LLM dynamics under a simplified setting.
-In that sense, this paper serves as a proof-of-concept on the value of studying the dynamics of LLM architectures when designing optimizers.
-We hope both research insights can inspire more research alone this line to further advance this field.
+We introduced SWAN, a stateless optimizer for LLM training that combines two well-known operators applied to raw gradients, $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$, to stabilize the stochasticity of gradient distribution and neutralizing the local geometry of loss landscape, respectively.
+Through theoretical analysis and empirical evidence, we showed that SWAN reduces memory usage while achieving on-par or even better performance compared to Adam on LLaMA pre-training tasks.
+Notably, SWAN achieves 2×2$\times$2 × speedups compared to Adam in terms of tokens processed when training 350M- and 1.3B-parameter models.
+These findings serve as a proof-of-concept that the stateless approach has the potential to serve as a practical and efficient alternative to other optimizers that require tracking internal states.
+Future work may explore other design choices for stateless optimizers and further expand SWAN’s applicability to other complex training regimes beyond standard LLM pre-training.
 Appendix A Desired properties of adaptive optimizers
@@
 Therefore, we hypothesize that Adam approximately whitens the gradients element-wise, leading to well-conditioned regions.
-Appendix B Theoretical Analysis
-B.1 Analyzing the GradNorm: A Transformer Learning Dynamics Perspective
-We take the transformer LLM architecture as an example, and demonstrate that $\mathtt{GradNorm}$ effectively removes the time-variant components of gradient noises.
-For our analysis, we assume the following simplified transformer block (STB) architecture recently proposed in Tian et al.
-(2023), which contains the major building blocks of transformers and relies on relatively less restrictive assumptions and simplifications, making them suitable for theoretical analysis for transformer dynamics.
-It is defined as follows:
-Definition 1 (Simplified Transformer Block (STB)).
-Given the input activation ${\bm{x}}\in\mathbb{R}^{M_{C}\times 1}$, query token index $q$, context embedding matrix ${\mathbf{U}}_{C}\in\mathbb{R}^{d\times M_{C}}$, and the query embedding ${\bm{u}}_{q}\in\mathbb{R}^{d\times 1}$, the simplified transformer block computes the output ${\bm{h}}\in\mathbb{R}^{n\times 1}$ as follows:
-${\bm{h}}=\phi\left({\mathbf{W}}^{\top}\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)\right),$
-where the attention logits ${\bm{z}}_{q}\in\mathbb{R}^{M_{C}\times 1}$ are given by
-$z_{ql}={\bm{u}}_{q}^{\top}{\mathbf{W}}_{Q}^{\top}{\mathbf{W}}_{K}u_{l},$
-with ${\mathbf{W}}_{Q},{\mathbf{W}}_{K}\in\mathbb{R}^{d\times d}$ being weight matrices for the queries and keys, respectively, ${\mathbf{W}}\in\mathbb{R}^{d\times n}$ is the weight matrix for the feedforward network, and $\phi$ is a nonlinearity function such as ReLU.
-Given a STB, we consider the following standard mini-batch learning dynamics.
-Define the conditional expectation $\mathbb{E}_{q=m}[\cdot]:=\mathbb{E}[\cdot|q=m]$.
-Consider the dynamics of the weight matrix $W$ and the attention logits $z_{q}$, if we train the model with a batch of inputs that always end up with query $q[i]=m$.
-The weight update for $W$ and $z_{q}$ are given by the following noisy updates:
-$\dot{{\mathbf{W}}}^{(t)}=\mathbb{E}_{q=m}\left[{\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}\right]+{\bm{\varepsilon}}^{(t)},\quad\dot{{\bm{z}}}_{m}^{(t)}=\mathbb{E}_{q=m}\left[\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}\right]+{\bm{\varsigma}}^{(t)},$ (10)
-Where ${\bm{f}}^{(t)}=\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q}^{(t)})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)$, $({\bm{h}}^{(t)})^{\prime}=\phi^{\prime}(({\mathbf{W}}^{(t)})^{\top}{\bm{f}}^{(t)})$ is the derivative of the current activation, ${\mathbf{G}}_{{\bm{h}}}^{(t)}=\nabla_{{\bm{h}}^{(t)}}\mathcal{L}$ is the gradient of the loss function $\mathcal{L}$ with respect to the hidden activation ${\bm{h}}^{(t)}$, and ${\bm{g}}_{{\bm{f}}^{(t)}}^{(t)}=\sum_{k}{\bm{g}}_{{\bm{h}}_{k}^{(t)}}^{(t)}({\bm{h}}^{(t)}_{k})^{\prime}{\bm{w}}_{k}^{(t)}$ is the sum of the gradients with respect to the attention logits.
-Here, ${\bm{w}}_{k}^{(t)}$ is the $k$-th column of ${\mathbf{W}}^{(t)}$, ${\bm{g}}_{h_{k}^{(t)}}^{(t)}[i]$ be the backpropagated gradient sent to node $k$ at sample $i$.
-Finally, ${\bm{\varepsilon}}$ and ${\bm{\varsigma}}$ are random noises that are induced by the minibatch sampling.
-Remark
-Given STB under the dynamics of Equation 10, the distribution of noises ${\bm{\varepsilon}}^{(t)}$ and ${\bm{\varsigma}}^{(t)}$ generally vary across time, since both the integrand within expectation, ${\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}$ and $\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}$ vary across time.
-However, the following result on $\mathtt{GradNorm}$ suggests that this time variability of ${\bm{\varepsilon}}^{(t)}$ and ${\bm{\varsigma}}^{(t)}$ are completely removed by $\mathtt{GradNorm}$:
-Theorem 1 ($\mathtt{GradNorm}$ stabilizes gradient distribution across time for STB).
-Consider the STB (Definition 1).
-Assuming we inherit the assumptions in Theorem 1 of Tian et al.
-(2023), as described in Appendix B.
-Then consider ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, the composition of the MLP project-up matrix and the embedding matrix as a whole.
-Then, its standardized stochastic gradients $\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}}(\top{\bm{x}}^{(t)})}{{\mathbf{U}}_{C}^{\mathbf{W}}})$ satisfy:
-$\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{1})}]=\mathrm{Cov}[\tilde{{\mathbf{G}}}_{{\mathbf{U}}_{C}^{\top}{\mathbf{W}}}[i,:]^{(t_{2})}]\quad\text{for all }t_{1},t_{2},\text{and}\,i.$
-In other words, the covariance structure of $\tilde{{{\mathbf{G}}}}$ is identical across all time steps $t$, achieving distributional stability across time.
-The same relationship also holds for the gradient of attention score $\tilde{{\mathbf{G}}}_{{\bm{z}}_{q}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}(\top{\bm{x}}^{(t)})}{\partial{\bm{z}}_{q}})$.
-Theorem 1 suggests that $\mathtt{GradNorm}$ makes use of the inductive bias induced by transformer architectures and directly removes the time-variant components in gradient covariance structures caused by mini-batch sampling.
-In practice, $\mathtt{GradNorm}$ makes the following approximation to Theorem 1:For MLP layer $\mathtt{GradNorm}$ is applied on ${\mathbf{W}}$ instead of ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$ as a whole, since such grouping does not exist in practical transformer architectures.
-However, it still works reasonably well in practice (Section 5.1, Section 5.6).
-B.2 Analyzing GradWhitening Pt. I: A Transformer Learning Dynamics Perspective
-Figure 9: Normalized Hessian Blocks of size $M_{C}\times M_{C}$ along the diagonal direction of the Hessian, obtained from numerically solving the STB ODE (with $n=12,M_{C}=10$) (1) given by the full-batch dynamics (i.e., removing noise in Equation 10).
-During all training steps, we analytically track the evolution of Hessian.
-As predicted by Proposition 1, we see very similar structures across the diagonal blocks of the Hessian.
-In this section, we take the perspective of transformer learning dynamics, and show that how $\mathtt{GradWhitening}$ can be naturally derived from the simplified transformer model defined in Definition 1.
-Let us first consider the dynamics of the Hessian matrix of STB.
-We will show that: i), it turns out that at equilibirum, the Hessian is highly structured; ii), under some conditions, $\mathtt{GradWhitening}$ recovers Newton’s method-like behaviors under such Hessian structures.
-Below we first present the following result on the Hessian structure of the simplified transformer block at equilibrium:
-Proposition 1 (Shared structures in the block-diagonal of Hessians at transformer equilibrium).
-Consider a STB (1), trained by full-batch dynamics (i.e., removing noise in Equation 10).
-Assume we inherit all the assumptions from Theorem 1 of Tian et al.
-(2023).
-Consider the following Hessian matrix:
-${\mathbf{H}}({\mathbf{W}})_{sk,s^{\prime}k^{\prime}}=\frac{\partial\mathcal{L}}{\partial{w_{sk}}\partial{w_{s^{\prime}k^{\prime}}}}$
-where $1\leq s,s^{\prime}\leq m$, $1\leq k,k^{\prime}\leq n$. Then, as $t\xrightarrow[]{}\infty$, we have the following shared Hessian structure along the diagonal blocks:
-$\frac{{\mathbf{H}}_{sk,s^{\prime}k}}{\sum_{s,s^{\prime}}{\mathbf{H}}_{sk,s^{\prime}k}}\xrightarrow[]{}\frac{{\mathbf{H}}_{sk^{\prime},s^{\prime}k^{\prime}}}{\sum_{s,s^{\prime}}{\mathbf{H}}_{sk^{\prime},s^{\prime}k^{\prime}}},\quad\forall 1\leq s,s^{\prime}\leq d,1\leq k,k^{\prime}\leq n$ (11)
-Proposition 1 shows that, under a simplified setting of transformer, because of the winner-takes-all behavior in the attention model, the Hessian will also converge to an equilibrium solution where the $M_{C}\times M_{C}$ blocks over the diagonal direction of Hessian shares the identical structure, after normalization.
-This result is verified in our numerical experiment as shown in Figure 9.
-Given by the inductive bias that equality across diagonal blocks of Hessians at transformer learning equilibrium, in the next theorem we show that $\mathtt{GradWhitening}$ can be derived as a second-order method under this structural assumption
-Theorem 2 ($\mathtt{GradWhitening}$ as structural approximation of Hessian).
-Consider the optimization problem given by $\min_{W\in\mathbb{R}^{m\times n}}f({\mathbf{W}})$ where ${\mathbf{W}}\in\mathbb{R}^{m\times n}$, with $m<n$. Assume:
-  • The local Hessian $H$ of $f$ has shared block-diagonal, such that ${\mathbf{H}}=\text{diag}(\tilde{{\mathbf{H}}},\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}})={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$, where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
-  • ${\mathbf{W}}^{(t)}$ is initialized as a orthogonal matrix.
-Then (under 2nd order approximation) $\mathtt{GradWhitening}$’s modified gradient descent (Equation 13) on $f$ recovers the update of Newton’s method, without explicitly estimate second-order quantities.
-To summarize, $\mathtt{GradWhitening}$ recovers a Newton’s method-like behavior under the highly structured Hessian regime (i.e., there exists $\tilde{{\mathbf{H}}}$ such that ${\mathbf{H}}={\mathbf{I}}_{n\times n}\otimes\tilde{{\mathbf{H}}}$), which is partially satisfied in transformer equilibrium dynamics (see remark below).
-Compared to directly applying accelerated second-order methods such as the Gauss-Newton method ($\mathcal{O}(Bm^{2}n^{2}+m^{3}n^{3})$, $B$ being batch size) or conjugate gradient method ($\mathcal{O}(BTmn)$ complexity, $T=$ 10 $\sim$ 100, with worst case complexity $\mathcal{O}(Bm^{2}n^{2})$), $\mathtt{GradWhitening}$ should be much faster ($\mathcal{O}(mn\min\{m,n\})$), especially under the Newton-Schulz iteration.
-Remark: limitations
-Proposition 1 deviates from the strict assumptions of Theorem 2 in that: Proposition 1 only asserts that diagonal direction of Hessian shares the identical structure, but does not assert that the off-diagonal blocks are zero, without further assumptions on $g_{h_{k}}$.
-However, given that it is standard practice to ignore off-diagonal elements in many practical optimizers including Adam, we hereby argue that ignoring off-diagonal blocks is already a better approximation.
-B.3 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
+Appendix B Additional Analysis
+B.1 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
 In this section, we present main results regarding the convergence rate of the $\mathtt{GradWhitening}$ method, understand its implications, and compare it with the lower bounds of GD and Adam.
@@
 Unlike the convergence rates of GD and Adam presented in Zhang et al.
-(2024), as well as Theorem 4 and Corollary 1 in Appendix, the optimal convergence rate (15) of $\mathtt{GradWhitening}$ no longer explicitly depends on the condition number $\kappa$ of $H$.
+(2024a), as well as Theorem 3 and Corollary 1 in Appendix, the optimal convergence rate (13) of $\mathtt{GradWhitening}$ no longer explicitly depends on the condition number $\kappa$ of $H$.
 In fact, consider a lower bound $\frac{\|{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}\|_{1}^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}\geq\frac{\text{Tr}[H{\mathbf{W}}_{\text{whitened}}^{(t)}]^{2}}{\text{Tr}[({\mathbf{W}}_{\text{whitened}}^{(t)})^{\top}{\mathbf{H}}{\mathbf{W}}_{\text{whitened}}^{(t)}]\text{Tr}[{\mathbf{H}}]}$, since trace of $H$ appear both in the nominator and denominator, we expect that to be more robust to ill-conditioned problems.
@@
 As pointed out by Zhang et al.
-(2024) and Da Silva & Gazeau (2020), Adam with $\beta_{2}<1$ will have issues with convergence, which will not be completely removed even with lr decay.
+(2024a) and Da Silva & Gazeau (2020), Adam with $\beta_{2}<1$ will have issues with convergence, which will not be completely removed even with lr decay.
 Therefore, we will not discuss the case of $\beta_{2}<1$ to avoid the complication.
-[…]
+B.2 Numerical Verification of Proposition 1
+Given a STB, we consider the following standard full-batch learning dynamics (Tian et al., 2023).
+Define the conditional expectation $\mathbb{E}_{q=m}[\cdot]:=\mathbb{E}[\cdot|q=m]$.
+Consider the dynamics of the weight matrix $W$ and the attention logits $z_{q}$, if we train the model with a batch of inputs that always end up with query $q[i]=m$.
+The weight update for $W$ and $z_{q}$ are given by the following noisy updates:
+$\dot{{\mathbf{W}}}^{(t)}=\mathbb{E}_{q=m}\left[{\bm{f}}^{(t)}({\mathbf{G}}_{{\bm{h}}}\odot{\bm{h}}^{\prime(t)})^{\top}\right],\quad\dot{{\bm{z}}}_{m}^{(t)}=\mathbb{E}_{q=m}\left[\left(\frac{\partial{\bm{b}}}{\partial{\bm{z}}_{m}^{(t)}}\right)^{\top}{\mathbf{U}}_{C}^{\top}{\bm{g}}_{f}^{(t)}\right],$ (15)
+Where ${\bm{f}}^{(t)}=\left({\mathbf{U}}_{C}\left(\exp({\bm{z}}_{q}^{(t)})\odot{\bm{x}}\right)+{\bm{u}}_{q}\right)$, $({\bm{h}}^{(t)})^{\prime}=\phi^{\prime}(({\mathbf{W}}^{(t)})^{\top}{\bm{f}}^{(t)})$ is the derivative of the current activation, ${\mathbf{G}}_{{\bm{h}}}^{(t)}=\nabla_{{\bm{h}}^{(t)}}\mathcal{L}$ is the gradient of the loss function $\mathcal{L}$ with respect to the hidden activation ${\bm{h}}^{(t)}$, and ${\bm{g}}_{{\bm{f}}^{(t)}}^{(t)}=\sum_{k}{\bm{g}}_{{\bm{h}}_{k}^{(t)}}^{(t)}({\bm{h}}^{(t)}_{k})^{\prime}{\bm{w}}_{k}^{(t)}$ is the sum of the gradients with respect to the attention logits.
+Here, ${\bm{w}}_{k}^{(t)}$ is the $k$-th column of ${\mathbf{W}}^{(t)}$, ${\bm{g}}_{h_{k}^{(t)}}^{(t)}[i]$ be the backpropagated gradient sent to node $k$ at sample $i$.
+Then, we numerically solving the STB ODE with $n=12,M_{C}=10$ in Equation 15.
+During all training steps, we analytically track the evolution of Hessian of $rmW$.
+Results are shown in Figure 9.
+As predicted by Proposition 1, we see very similar structures across the diagonal blocks of the Hessian.
+Figure 9: Normalized Hessian Blocks of size $M_{C}\times M_{C}$ along the diagonal direction of the Hessian, obtained from numerically solving the STB ODE (with $n=12,M_{C}=10$) (1) given by the full-batch dynamics (i.e., removing noise in Equation 15).
+During all training steps, we analytically track the evolution of Hessian.
+As predicted by Proposition 1, we see very similar structures across the diagonal blocks of the Hessian.
 Appendix C Proof of Theorem 1
@@
 To see this, consider the $\mathtt{GradWhitening}$ update:
-${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathcal{P}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}\mathcal{P}({\mathbf{H}}{\mathbf{O}}),$
-noticing that $\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}=1$, and $\mathcal{P}({\mathbf{H}}{\mathbf{O}})=\mathcal{P}({\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{O}})={\mathbf{Q}}{\mathbf{Q}}^{\top}{\mathbf{O}}={\mathbf{O}}$.
+${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}}),$
+noticing that $\frac{\|{\mathbf{H}}{\mathbf{O}}\|_{1}}{\text{Tr}({\mathbf{H}})}=1$, and $\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}})=\mathcal{P}({\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{O}})={\mathbf{Q}}{\mathbf{Q}}^{\top}{\mathbf{O}}={\mathbf{O}}$.
 Hence:
-${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathcal{P}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-{\mathbf{O}}=\mathbf{0}={\mathbf{W}}^{*}.$
+${\mathbf{W}}_{\text{whitened}}^{(1)}={\mathbf{O}}-\eta^{*}\mathtt{GradWhitening}({\mathbf{H}}{\mathbf{O}})={\mathbf{O}}-{\mathbf{O}}=\mathbf{0}={\mathbf{W}}^{*}.$
 Hence, the proof is complete. ∎
@@
 To prove Proposition 3, we first generalize existing work on the convergence rate lower bound (via contraction factor) of gradient descent and Adam (we only consider $\beta_{2}=1$) under the same setting:
-Theorem 4 (Contraction factor lower bound for gradient descent, generalized based on Zhang et al. (2024)).
+Theorem 3 (Contraction factor lower bound for gradient descent, generalized based on Zhang et al. (2024a)).
 Consider the optimization problem in Equation 10.
@@
 The proposition 1 in Zhang et al.
-(2024) has shown that the lower bound holds for diagonal positive definite Hessian ${\mathbf{H}}$.
+(2024a) has shown that the lower bound holds for diagonal positive definite Hessian ${\mathbf{H}}$.
 To show that the lower bound holds for a general positive definite Hessian ${\mathbf{H}}$ we will reformulate the problem to align with the setup in diagonal case (Prposition 1 of Zhang et al.
-(2024)).
+(2024a)).
 First, for any positive definite Hessian ${\mathbf{H}}$, we can perform an eigen decomposition ${\mathbf{H}}={\mathbf{U}}{\mathbf{S}}{\mathbf{U}}^{\top}$, where ${\mathbf{U}}$ is an orthogonal matrix and ${\mathbf{S}}$ is a diagonal matrix containing the eigenvalues of ${\mathbf{H}}$.
@@
 which reduces the problem to the case of a diagonal ${\mathbf{H}}$ with condition number $\kappa=\frac{\lambda_{1}}{\lambda_{m}}$, where $\lambda_{1}$ and $\lambda_{m}$ are the largest and smallest eigenvalues of ${\mathbf{H}}$, respectively.
-Thus, by applying Proposition 1 of Zhang et al. (2024) to this transformed problem, we conclude that there exists initial point such that the lower bound on the contraction rate
+Thus, by applying Proposition 1 of Zhang et al. (2024a) to this transformed problem, we conclude that there exists initial point such that the lower bound on the contraction rate
 $\mathcal{L}({\mathbf{W}}_{GD}^{(t+1)})-\mathcal{L}^{*}\geq\left(1-\frac{2}{\kappa+1}\right)\left(\mathcal{L}({\mathbf{W}}_{GD}^{(t)})-\mathcal{L}^{*}\right)$
@@
 Proof.
-The update rule of Adam with $\beta_{1}=0$ and $\beta_{2}=1$ is given by Zhang et al. (2024):
+The update rule of Adam with $\beta_{1}=0$ and $\beta_{2}=1$ is given by Zhang et al. (2024a):
 ${\mathbf{W}}_{\text{Adam}}^{(t+1)}={\mathbf{W}}_{\text{Adam}}^{(t)}+\eta\,\text{diag}(|{\mathbf{H}}{\mathbf{W}}^{(0)}|^{-1})\,{\mathbf{H}}{\mathbf{W}}_{\text{Adam}}^{(t)}.$
@@
 This holds for all $1\leq s,s^{\prime}\leq d$ and $1\leq k,k^{\prime}\leq n$. ∎
-Appendix H Proof of Theorem 2
-Proof.
-Assume that we have a function $f({\mathbf{W}})$ defined on ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ that we want to minimize, and the global minimum is at ${\mathbf{W}}^{*}$ (without loss of generality, let ${\mathbf{W}}^{*}=0$).
-Then our problem can be written as:
-$\min_{{\mathbf{W}}\in\mathbb{R}^{m\times n}}f({\mathbf{W}})$ (22)
-Assume $f$ is a twice differentiable function and has a Lipschitz continuous Hessian around ${\mathbf{W}}^{*}$, then both $\mathtt{GradWhitening}$ and Newton’s method approximate the loss equation 22 up to second order (ignoring first-order terms without loss of generality) as:
-$\min_{{\mathbf{w}}\in\mathbb{R}^{mn}}\frac{1}{2}{\mathbf{w}}^{T}{\mathbf{H}}{\mathbf{w}},$ (23)
-where ${\mathbf{w}}\in\mathbb{R}^{mn\times 1}$ is the flattened version of ${\mathbf{W}}$, and ${\mathbf{H}}\in\mathbb{R}^{mn\times mn}$ is the Hessian matrix of $f$ at ${\mathbf{W}}^{*}$.
-We may apply second-order Newton’s method to optimize equation 23, which gives the following update:
-${\mathbf{w}}^{t+1}={\mathbf{w}}^{t}-\eta{\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t}).$ (24)
-However, this update is very expensive and requires $\mathcal{O}(m^{3}n^{3})$ computation.
-SWANcan then be viewed as a fast approximation to estimate the update of equation 24.
-Under our assumption on Hessian structure, we have:
-${\mathbf{H}}=\text{diag}(\tilde{{\mathbf{H}}},\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}})$ (25)
-where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
-Under this assumption, the optimization problem equation 23 can be rewritten as:
-$\min_{{\mathbf{W}}\in\mathbb{R}^{m\times n}}\frac{1}{2}\text{Tr}\left[{\mathbf{W}}^{\top}\tilde{{\mathbf{H}}}{\mathbf{W}}\right],$ (26)
-For this model, the gradient of ${\mathbf{W}}$ is given by:
-$\nabla f({\mathbf{W}})=\tilde{{\mathbf{H}}}{\mathbf{W}}.$
-Let ${\mathbf{W}}$ have rank $m$, and assume $n>m$.
-We say ${\mathbf{W}}$ is now over-parameterized in the sense that its dimensionality is now larger than the dimensionality of the second-order information $\tilde{{\mathbf{H}}}$.
-It is now clear that the second-order information is losslessly compressed into gradients: one can exactly recover the Hessian via
-${\mathbf{H}}=\text{diag}(\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\dots,\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\dots,\nabla f({\mathbf{W}}){\mathbf{W}}^{-1}).$
-Given this fact, one can easily estimate Newton’s update ${\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t})$ without explicitly estimating ${\mathbf{H}}^{-1}$ (or $\tilde{{\mathbf{H}}}$). Notice that:
-${\mathbf{H}}={\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}$
-where ${\mathbf{I}}_{m}$ is the $m\times m$ identity matrix and $\otimes$ denotes the Kronecker product. Given that
-${\mathbf{H}}^{-1}={\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}^{-1},$
-we have
-${\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t})=({\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}^{-1})\,\text{vec}(\nabla f({\mathbf{W}}^{t}))=\text{vec}\left(\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t})\right).$
-Reshaping back to matrix form, we can rewrite the update rule as:
-$\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t}).$
-Next, since $\tilde{{\mathbf{H}}}$ is positive definite:
-$\nabla f({\mathbf{W}}^{t})=\tilde{{\mathbf{H}}}{\mathbf{W}}^{t}={\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{W}}^{t},$
-where ${\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}$ is the eigen decomposition of $\tilde{{\mathbf{H}}}$.
-Assume further that ${\mathbf{W}}^{t}$ is initialized as orthogonal; one can immediately recognize that ${\mathbf{Q}}$ is also the left singular vector matrix of $\nabla f({\mathbf{W}}^{t})$.
-Therefore, the matrix inversion of $\tilde{{\mathbf{H}}}$ required in Newton’s update equation 24 can be easily achieved by simply deleting the $\Lambda$ term.
-That is,
-$\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t})={\mathbf{Q}}\Lambda^{-1}{\mathbf{Q}}^{\top}{\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{W}}^{t}$
-$=$ ${\mathbf{Q}}\left[{\mathbf{Q}}^{\top}{\mathbf{W}}^{t}\right]\overset{{\mathbf{U}}:={\mathbf{Q}},{\mathbf{V}}:=({\mathbf{W}}^{t})^{\top}{\mathbf{Q}}}{=}{\mathbf{U}}{\mathbf{V}}^{\top}=\mathcal{P}\left(\nabla f({\mathbf{W}}^{t})\right)={\mathbf{G}}_{\text{whitened}}.$
-This is exactly the $\mathtt{GradWhitening}$-modified gradient update. ∎
-Appendix I Implementation details
+Appendix H Implementation details
 Setup
@@
 (2024a), under exactly the same model configs and optimizer configs.
-This include the same model architecture, tokenizer, batch size, context length, learning rate scheduler, etc.
-The only exception is that since SWAN does not need warm up, we use zero-warm up ratio for SWAN.
-But we still ensure that the learning rates of all methods are fully decayed to the same target learning rate (0.1 $\times$ inital learning rate).
+This includes the same model architecture, tokenizer, batch size, context length, learning rate scheduler, etc.
+The only exception is that since SWAN does not need warm-up, we use zero-warm-up ratio for SWAN.
+But we still ensure that the learning rates of all methods are fully decayed to the same target learning rate (0.1 $\times$ initial learning rate).
 Across all model sizes, we use the same hyperparameter settings and the same learning rate tuning procedure.
-The only exception is that for model of size 1.3B: as we already know that larger model requires smaller learning rates, we conduct learning search for Adam and Galore over a smaller but more fine grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$.
+The only exception is that for a model of size 1.3B: as we already know that a larger model requires smaller learning rates, we conduct a learning search for Adam over a smaller but more fine-grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$.
 As a result, the optimal learning rate found for Adam on 1.3B is 0.0007.
 Applying SWAN on LLMs
-Since SWAN utilizes matrix-level operations on gradients, it can only be applied on 2D parameters.
-Therefore, in our experiments we only apply $SWAN{}$ on all linear projection weights in transformer blocks.
-Similar to Galore (Zhao et al., 2024a), the rest of the non-Linear parameters still uses Adam as default choice.
+Since SWAN utilizes matrix-level operations on gradients, it can only be applied to 2D parameters.
+Therefore, in our experiments, we only apply $SWAN{}$ on all linear projection weights in transformer blocks.
+Similar to Galore (Zhao et al., 2024a), the rest of the non-linear parameters still uses Adam as the default choice.
 Therefore, we follow the learning rate setup of Galore, where we fix a global learning rate 0.001 across all model sizes and all modules.
 Then, for the linear projection modules where SWAN is applied, we simply apply a scaling factor $\alpha$ on top of the global learning rate.
-We choose $\alpha=1$ for all model sizes except for 1.3 B case, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$.
-Note that this choice is quite random; and it is not extensively grid-searched.
-Applying $\mathtt{GradWhitening}$
-We found that one strategy that lead to better convergence is to first convert the gradient matrice to FP32, perform $\mathtt{GradWhitening}$, and then convert back to BF16.
-This is the default setup in Section 5.3.
-In practice, we found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence.
-We found that setting $\beta=0.8$ with iteration number $k=10$ generally give satisfactory performance.
-This can be further optimized and we leave it for future work.
+We choose $\alpha=1$ for all model sizes except for 1.3 B case.
+This is because we observe that the optimal learning rate of Adam under 1.3B becomes smaller than 0.001, hence we also reduce the learning rate on SWAN, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$.
+Note that this choice is quite arbitrary, and it is not extensively grid-searched.
+This is to show how the method work out-of-the box, without extensive problem-specific tuning.
+Implementation of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
+For $\mathtt{GradNorm}$, we simply set $\bar{{\bm{g}}}=0$ to simplify computation, as we found in practice this often has small scales.
+For $\mathtt{GradWhitening}$, before N-S iteration, we further normalize its input matrix by its Frobenius norm.
+This is applied in all our ablation studies as well, regardless of whether the gradient is processed by $\mathtt{GradNorm}$.
+For the N-S iteration, we found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence.
+We set $\beta=0.8$ with iteration number $k=10$ and it generally gives satisfactory performance.
+$\mathtt{GradWhitening}$ can be performed in both FP32 and BF16 format.
+In FP32 format, we first convert the gradient matrice to FP32, perform $\mathtt{GradWhitening}$, and then convert back to BF16.
+This is the default setup in Section 6.1 unless stated otherwise.
 Alternatively, $\mathtt{GradWhitening}$ can also be done in BF16 precision.
@@
 We found that this usually translates well.
+Note that this might be further optimized similar to (Jordan et al., 2024), we leave it to future work as N-S performance optimization is orthogonal to our key contribution.

### v3

@@
 However, its capability in LLM training is limited (Zhao et al., 2024b).
-In this work, we show that pre-processing SGD in a stateless manner can achieve the same performance as the Adam optimizer for LLM training, while drastically reduces the memory cost.
-Specifically, we propose to pre-process the instantaneous stochastic gradients using normalization and whitening.
-We show that normalization stabilizes gradient distributions, and whitening counteracts the local curvature of the loss landscape.
+In this work, we show that pre-processing SGD using normalization and whitening in a stateless manner can achieve the same performance as the Adam optimizer for LLM training, while maintaining the same memory footprint of SGD.
+Specifically, we show that normalization stabilizes gradient distributions, and whitening counteracts the local curvature of the loss landscape.
 This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any optimizer states.
-Empirically, SWAN has the same memory footprint as SGD, achieving $\approx 50\%$ reduction on total end-to-end memory compared to Adam.
+Empirically, SWAN achieves $\approx 50\%$ reduction on total end-to-end memory compared to Adam.
 In language modeling tasks, SWAN demonstrates comparable or even better performance than Adam: when pre-training the LLaMA model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity using half as many tokens.
@@
 Therefore, a natural question arises:
-Can LLMs be trained efficiently using stateless optimizers?
+Can LLMs be trained efficiently, without storing any optimizer states?
 There is a growing body of research that has contributed to answering this question positively by developing novel optimizers that reduce the memory requirements associated with tracking internal state variables during the training of LLMs while achieving similar or even speedup boost performance compared to Adam.
@@
 Another line of work focuses on using low-rank approximations to store the first and second moments, thereby reducing the memory cost associated with tracking optimizer states (Lialin et al., 2023; Hao et al., 2024; Zhao et al., 2024a).
-More recently, Zhu et al.
-(2024) introduced an almost stateless optimizer for efficient LLM training, which proposes to only track rank-1 state variables during training.
+Finally, the concurrent work of  (Zhu et al., 2024) proposed an approximate gradient scaling method, that tracks the channel-wise or tensor-wise surrogate learning rates, further improving the memory efficiency.
 In this work, we address this question by proposing to simply pre-process the instantaneous stochastic gradient in a stateless manner.
@@
 Our optimizer consists of combining two well-known operators to pre-process the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$.
-$\mathtt{GradNorm}$ applies a row-wise standardization on the gradient matrix, while $\mathtt{GradWhitening}$ orthogonalizes the normalized gradient matrix.
+$\mathtt{GradNorm}$ applies a row-wise normalization on the gradient matrix, while $\mathtt{GradWhitening}$ orthogonalizes the normalized gradient matrix.
 We show that these operators aims at stabilizing the stochasticity of gradient distributions during training, and neutralizing the local geometry of the loss landscape, respectively.
@@
     1. 1.
-      Robustness to ill-conditioned problems: on LLM pre-training, SWAN consistently and efficiently converges without learning rate warm-up (Section 6.1, Section 6.4).
+      Memory efficiency: SWAN only requires the memory footprint of SGD (w/o momentum), that is: $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states compared to Adam.
+This can be further reduced with LOMO technique (Lv et al., 2023) (Figure 1 (c)).
     2. 2.
-      Memory efficiency: SWAN achieves the memory footprint of SGD, that is $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states when compared to Adam.
+      Sample efficiency: through experiments (Section 6) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also achieves similar or even better performance compared to Adam.
+Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
     3. 3.
-      Computational efficiency: consistently achieves the same or better performance when compared to Adam across several scales and can significantly improve token-efficiency during training up to 2$\times$ at the 1B scale for the LLaMA model.
-  • Theoretical analysis of SWAN.
-We show that (1) $\mathtt{GradNorm}$ can stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix B); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 5.3).
-Additionally, we highlight that in the quadratic case, $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 2, Section B.1).
-  • Fast convergence for LLM pretraining.
-Through experiments (Section 6) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also achieves similar or even better performance compared to Adam.
-Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
-(a) 350 M LLaMA model Pretraining
-(b) 1.3 B LLaMA model Pretraining
+      Computational efficiency: we propose a efficient scheme to accelerate the computation of SWAN (Section 4.2), achieving a similar throughput to Adam without the need for distributed computation (Section 6.2) required by recent optimizers such as shampoo Gupta et al.
+(2018).
+    4. 4.
+      Robustness to hyperparameters: the hyperparameters of SWAN are lazily tuned via heuristics; and they are shared across all model sizes.
+Even when compared to tuned-baselines obtained with learning rate sweeps, SWAN still delivers strong performances on LLM pretraining.
+This is a promising indication of applicability to real-world problems.
+  • Theoretical consistency with LLM dynamnics.
+We show that (1) $\mathtt{GradNorm}$ can stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix C); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 5.3).
+Additionally, we highlight that in the quadratic case, $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 2, Section C.1).
+Table 1: The summary of memory consumption breakdown of optimizers, and the corresponding learning rate expressiveness.
+We assume all optimizers are applied on a 2D tensor of $m$ by $n$ ($m<n$).
+The optimizer state is defined as all intermediate variables maintained over time; while expressiveness is defined as the number of adaptive learning rates that a optimizer can model per tensor.
+Adam SGD SGD-Sal Apollo (rank $r$) Apollo (mini) GaLore (rank $r$) SWAN
+Optimizer States $\downarrow$ $2mn$ 00 $mn$ $2nr+2+(mr)$  [^1] The factor $mr$ comes from the storage of random projection tensors.
+It can be reduced by only storing the random seeds; however, this might cause challenges in a distributed setting.
+$2n+2+(m)$ $2nr+mr$ 00
+Model weights $\downarrow$ $mn$ $mn$ $mn$ $mn$ $mn$ $mn$ $mn$
+Expressiveness $\uparrow$ $mn$ $1$ blockwise $n$ $1$ $mn$ $mn$
+(a) 350M model pretraining
+(b) 1.3B model Pretraining
 (c) End-to-end memory footprint
-(d) Effective training throughput (same batch size)
-Figure 1: SWAN performance preview on LLM pretraining.
-(a) and (b): On both 350M and 1.3B LLama architectures, SWAN achieves $>$ 2X speed-up vs Adam in terms of tokens seen.
+(d) Throughput w/o distributed processing
+Figure 1: SWAN Performance on LLM Pretraining.
+(a) and (b): On both 350M and 1.3B LLama architectures, SWAN achieves over 2X speed-up compared to Adam in terms of tokens seen.
 (c): Memory footprint.
-We directly measure end-to-end memory under full-model training, with batch size = 1 sequence.
-This is different to the layer-wise training setting Zhao et al.
-(2024a).
-SWAN achieves near-SGD optimizer memory reduction ($\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states).
-(d) Training throughput analysis on training 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens.
-We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the token efficiency of optimizer relative to Adam.
-SWAN closely reproduces the absolute throughput of Adam, while improving the effective throughput.
+We measure end-to-end memory usage during full-model training with a batch size of 1 sequence.
+SWAN achieves nearly 100% reduction in optimizer states memory, and 50% total memory reduction (up to 70% when combined with per-layer training technique Zhao et al.
+(2024a); Lv et al.
+(2023)), and (d): Training throughput of a 1.3B model without distributed gradient pre-processing.
+With our acceleration scheme (denoteed by SWAN), we closely match the throughput of Adam without the need for distributed computation (Shi et al., 2023).
+The rightmost bar shows the effective throughput, adjusted for the token efficiency of the optimizer relative to Adam.
 2 Related works
-Towards Stateless Optimizers for LLM Training.
+Towards Stateless LLM Training.
 Adaptive optimizers generally rely on tracking internal state variables to perform weight updates, which can substantially increase memory consumption when training large models.
@@
 Alternatively, PSGD (Pooladzandi & Li, 2024; Li, 2017) focuses on exclusively tracking a pre-conditioner, eliminating the need to track a first moment estimate.
-Finally,  (Zhu et al., 2024) proposes an nearly stateless optimizer to train LLMs, which only requires tracking rank-1 first and second moments.
+Finally,  (Zhu et al., 2024) proposes a optimizer based on approximate gradient scaling to train LLMs, which requires tracking channel-wise or tensor-wise approximation to element-wise adaptive learning rates.
 However, all the aforementioned optimizers still require the storage of state variables.
@@
 Pre-processing Gradients.
-Gradient pre-processing is a common technique used to enhance performance of optimizers.
+Gradient pre-processing is a common technique to improve optimizer performance.
 Various pre-processing procedures have been proposed in the literature, such as signed gradient (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023; Kunstner et al., 2023), gradient clipping (Zhang et al., 2020), normalization (Zhang et al., 2020; You et al., 2019), and whitening (Yang & Laaksonen, 2008; Kingma & Ba, 2015; Hwang, 2024; Jordan et al., 2024; Bernstein & Newhouse, 2024c; a; Carlson et al., 2015).
-In this work, we particularly focus on normalization and whitening.
-We apply normalization row-wise on gradient matrices, similar to (You et al., 2019); together with gradient whitening under a specific structural assumption of the Fisher Information (Hwang, 2024; Martens et al., 2018), recovering the orthogonalization step used in (Jordan et al., 2024; Tuddenham et al., 2022).
-Our key result is that composing normalization and whitening on stochastic gradients is sufficient to enable the efficient training of LLMs in a completely stateless manner.
-Compared to Lamb (You et al., 2019), our normalization operation is applied on raw gradients instead of Adam states.
-Compared to Muon (Jordan et al., 2024), SWAN removes first-moment tracking and instead uses normalization.
-We show empirically that removing any one of these two pre-processing steps from SWAN results in significant performance degradation (Section 6.4).
+In this work, we focus on normalization and whitening.
+We apply normalization row-wise on gradient matrices, together with gradient whitening under a specific structural assumption of the Fisher Information (Hwang, 2024; Martens et al., 2018), recovering the orthogonalization step used in (Jordan et al., 2024; Tuddenham et al., 2022).
+Our key result is that composing normalization and whitening on stochastic gradients can be sufficient to enable the efficient training of LLMs in a stateless manner.
+In our setting, we show empirically that both removing any one of these two pre-processing steps from SWAN results in significant performance degradation (Section 6.4).
+Compared to Lamb (You et al., 2019), our normalization operation is applied on raw gradients row-wise; while LAMB applied layer-wise global normalization on Adam states instead.
+Compared to Muon (Jordan et al., 2024), SWAN removes first-moment tracking and instead uses row-wise normalization, making the optimizer fully statless.
+Moreover, our proposed efficient heuristic scheme for computing square root inverse requires $\mathcal{O}(m^{2})$ instead of $\mathcal{O}(m^{3})$ of the standard NS scheme (Song et al., 2022; Li et al., 2018; Huang et al., 2019).
 Low-rank methods.
@@
 (2019)), we treat each column of the gradient matrix ${\mathbf{G}}$ as a separate data sample.
-Then, $\mathtt{GradWhitening}$ can be seen as stretching/squeezing the gradient data along all eigen directions, such that the covariance matrix is the identity matrix.
+Then, $\mathtt{GradWhitening}$ can be seen as stretching/squeezing the data such that the covariance matrix is the identity across all eigen directions.
 Algorithm 1 SWAN Optimizer
@@
      ${\mathbf{G}}^{(t)}\in\mathbb{R}^{m\times n}\leftarrow\nabla_{{\mathbf{W}}^{(t)}}\mathcal{L}^{(t)}({\mathbf{W}}^{(t)})$
-     $\tilde{{\mathbf{G}}}^{(t)}\leftarrow\mathtt{GradNorm}({\mathbf{G}}^{(t)})$
+     $\tilde{{\mathbf{G}}}^{(t)}\leftarrow\mathtt{GradNorm}({\mathbf{G}}^{(t)})$ (Eq (1))
 […]
@@
 Algorithm 2 $\mathtt{GradWhitening}$ Operator
-  Input: ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of iterations $k$. Step size $\beta$ (default = 0.8).
+  Input: ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of iterations $K$. Step size $\beta$. Boolean diag indicating if use diagonal substitution scheme.
   Initialize ${\mathbf{Y}}\leftarrow{\mathbf{G}}{\mathbf{G}}^{\top}$, ${\mathbf{Z}}\leftarrow{\mathbf{I}}$
   for i = 1,…,K do
+     if diag then
+        ${\mathbf{Y}}\leftarrow\beta{\mathbf{Y}}\text{Diag}(3{\mathbf{I}}-{\mathbf{Z}}\text{Diag}({\mathbf{Y}}))$,
+        ${\mathbf{Z}}\leftarrow\beta(3{\mathbf{I}}-\text{Diag}({\mathbf{Z}}){\mathbf{Y}})\text{Diag}({\mathbf{Z}})$
+     else
         ${\mathbf{Y}}\leftarrow\beta{\mathbf{Y}}(3{\mathbf{I}}-{\mathbf{Z}}{\mathbf{Y}})$,
         ${\mathbf{Z}}\leftarrow\beta(3{\mathbf{I}}-{\mathbf{Z}}{\mathbf{Y}}){\mathbf{Z}}$
+     end if
   end for
   return  ${\mathbf{Z}}{\mathbf{G}}$
-Algorithm 3 $\mathtt{GradNorm}$ Operator
-  Input: ${\mathbf{G}}^{m\times n}$
-  return  $\frac{{\mathbf{G}}-\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}\mathbf{1}_{n}^{\top}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\frac{1}{n}\sum_{k=1}^{n}{\mathbf{G}}_{:,k})^{2}}\mathbf{1}_{n}^{\top}}$
-Figure 3: SWAN Optimizer.
-Note that for $\mathtt{GradWhitening}$, in principle, any approximate algorithms can be applied, such as matrix Pade approximate Song et al.
-(2022).
-Here we only show one particular working example based on Newton-Schulz, which introduces minimal overheads.
+Figure 3: SWAN Optimizer. When $\mathtt{diag}$ is set to $\mathtt{True}$, we cover the fast variant denoted by SWAN.
 As discussed in Section 3, we believe the key to designing stateless, adaptive, and effective optimizers lies in the incorporation of matrix-level operations that exploit rich information contained in the gradient matrix.
@@
 Let $1\leq i\leq m$ represent the input indices and $1\leq j\leq n$ represent the output indices.
-Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, as done in Adam, we propose to standardize across the output dimensions at each time step $t$, as follows:
-$\mathtt{GradNorm}({\mathbf{G}}):=\frac{{\mathbf{G}}-\bar{{\bm{g}}}\mathbf{1}_{n}^{\top}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
-where $\bar{{\bm{g}}}:=\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}$ is the mean across output dimension; ${\bm{s}}:=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\bar{{\bm{g}}})^{2}}$ is the standard deviation across dimension; $\bar{{\bm{g}}}$ and ${\bm{s}}$ are ${m}$-dimensional column vectors; $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
-$\mathtt{GradNorm}$ is the forward pass operator LayerNorm (Ba et al., 2016) (without learnable parameters) applied on backward gradients.
-Similar operations have been considered in the literature to process gradients (You et al., 2019).
-$\mathtt{GradNorm}$ allows the optimizer to be invariant under the following transformations on the gradient matrix: i) matrix-wise re-scaling; ii) matrix-wise re-centering; iii) row-wise re-scaling; and iv) row-wise re-centering.
-Compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) used in signed descent, $\mathtt{GradNorm}$ preserves richer information of the gradient scaling while offering invariance properties.
-In Section 5 we will show that when applied to a simplified transformer architecture, $\mathtt{GradNorm}$ also helps stabilize the gradient distribution.
+Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, as done in Adam, we propose to standardize across the output dimensions at each time step $t$:
+$\mathtt{GradNorm}({\mathbf{G}}):=\frac{{\mathbf{G}}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
+where ${\bm{s}}:=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}})$ is ${m}$-dimensional the root mean square across dimension ( ${m}$-dimensional column vector); $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
+$\mathtt{GradNorm}$ is the forward pass operator RMS-Norm (Zhang & Sennrich, 2019) applied on backward gradients.
+$\mathtt{GradNorm}$ allows the optimizer to be invariant under matrix-wise and row-wise rescaling.
+In Section 5 we show that this is the key to stabilizing the LLM gradient distribution.
+Meanwhile, compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) alternatives (Zhang et al., 2020; You et al., 2019), $\mathtt{GradNorm}$ preserves richer information of gradient scaling while offering invariance properties.
 On the $\mathtt{GradWhitening}$ Step.
 As discussed in Section 3, Adam relies on a second moment estimate to perform element-wise gradient whitening.
-More formally, the second moment of Adam estimate a diagonal approximation of the Fisher information matrix (FIM) via (Kingma & Ba, 2015; Hwang, 2024):
+More formally, its second moment estimate a diagonal approximation to the Fisher information matrix (FIM):  (Kingma & Ba, 2015; Hwang, 2024):
 $\mathbb{E}(\text{Vec}({\mathbf{G}})\text{Vec}({\mathbf{G}})^{\top})\approx\text{Diag}[\text{Vec}[\mathbb{E}({\mathbf{G}}^{\odot 2})]]$
@@
 The derivation of Equation 2 as structured FIM/Hessian approximation is discussed in Section 5.3.
+Apart from the FIM/Hessian approximation view, Equation 2 can also be interpreted as minimizing the condition number of ${\mathbf{G}}$ (defined as the ratio between the largest and smallest singular values).
+It can be shown that after whitening, its condition number achieves the minimum value which is 1.
 […]
-4.2 Practical Considerations
+4.2 Practical Considerations: Fast and Stable Implementation
+Fast $\mathtt{GradWhitening}$
 Computing $\mathtt{GradWhitening}$ exactly can be expensive, as it involves solving the matrix square-root inverse.
-We hence directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation.
+One option is to directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation.
 This is given by (Song et al., 2022; Li et al., 2018):
-${\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}),\quad{\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}$ (3)
+$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k})\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}\end{cases}$ (3)
 where ${\mathbf{Y}}_{0}={\mathbf{G}}{\mathbf{G}}^{\top}$ , ${\mathbf{Z}}_{0}={\mathbf{I}}$.
 At convergence, $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{Z}}{\mathbf{G}}$ (Algorithm 2).
-See Appendix H for implementation details.
+See Appendix J for implementation details.
 Note that the same N-S procedure has been discussed in (Mei et al., 2023) and  (Jackson, 2023) for preconditioned optimizers.
-Another N-S procedure has also been introduced in (Bernstein & Newhouse, 2024b), and optimized in (Jordan et al., 2024) under BF16 precision.
-Computational Overhead.
-[…]
-Rescaling $\mathtt{GradWhitening}$.
-The whitening operation $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, and as such might drastically change the effective learning rate.
-In practice, we propose the following re-scaling before updating the weight:
-$\Delta{\mathbf{W}}^{(t)}\leftarrow\frac{\|\tilde{{\mathbf{G}}}^{(t)}\|\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}$ (4)
-This helps rescale the norm of the whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, and allowing us to get reasonably good performance without extensively tuning learning rates; we found that the learning rates used by Adam tend to work well for SWAN.
-We can remove rescaling, but doing so would require additional hyper-parameter search to find the optimal learning rate.
+Another N-S procedure was been introduced in (Bernstein & Newhouse, 2024b), and optimized in (Jordan et al., 2024).
+However, estimating $({\mathbf{G}}{\mathbf{G}}^{\top})^{-1/2}$ with NS requires $\mathcal{O}(m^{3})$ (assuming $m<n$) complexity, making its scalability for larger models, especially under distributed settings unclear Jordan et al.
+(2024).
+Here, we propose a heuristic scheme that has $\mathcal{O}(m^{2})$ complexity to estimate square-root inverse (comparable to the $\mathcal{O}(mn)$ complexity of Adam second moment):
+$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}\text{Diag}(3{\mathbf{I}}-{\mathbf{Z}}_{k}\text{Diag}({\mathbf{Y}}_{k}))\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-\text{Diag}({\mathbf{Z}}_{k}){\mathbf{Y}}_{k})\text{Diag}({\mathbf{Z}}_{k})\end{cases}$ (4)
+where under a slight abuse of notation, $\text{Diag}(\cdot)$ returns a diagonal matrix that has the same diagonals as the input matrix.
+Intuitively, whenever we encounter matrix multiplication in NS iterations, we replace one of the matrices with its diagonal approximation.
+The particular choice of where to apply $\text{Diag}(\cdot)$ is determined by performance on synthetic datasets.
+We refer to this as the NS with diagonal substitution (NSDS) scheme.
+In Appendix B, we empirically demonstrate that NSDS performs well in minimizing the matrix condition number of gradients.
+As shown in Figure 7, on the gradient distribution induced by Llama model training, the performance of NSDS (when combined with $\mathtt{GradNorm}$) is comparable to or better than the standard NS scheme in reducing the condition number.
+We found that NSDS does have the side effect of slowing down the early stage convergence.
+However, it offers stronger tail convergence in return.
+For LLM pretraining NSDS helps SWAN achieve similar or better final validation loss performance than the original NS scheme (Section 6.1), enabling Adam-level training throughput without the help of distributed computation (Section 6.2).
+Rescaling for robustness.
+The operator $\mathtt{GradWhitening}$ maps the normalized gradient $\tilde{{\mathbf{G}}}^{(t)}$ onto the closest orthogonal matrix, and as such might drastically change its effective learning rate.
+In practice, we propose the following re-scaling before updating the weights:
+$\Delta{\mathbf{W}}^{(t)}\leftarrow\frac{\|\tilde{{\mathbf{G}}}^{(t)}\|\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}=\frac{\sqrt{mn}\Delta{\mathbf{W}}^{(t)}}{\|\Delta{\mathbf{W}}^{(t)}\|}$ (5)
+This helps rescale the norm of the whitened gradient back to the norm of $\tilde{{\mathbf{G}}}^{(t)}$, that is $\|\tilde{{\mathbf{G}}}^{(t)}\|=\sqrt{mn}$.
+Notice that this is exactly the norm of signed gradient descent ($=$ Adam without EMAs) (Bernstein et al., 2018).
+This relationship allows us to directly adapt the learning rates from Adam without tuning specifically for SWAN.
+This can be throught as a variation of LR grafting (Agarwal et al., ).
+Finally, the practical SWAN algorithm ($\mathtt{GradNorm}$ + NSDS-$\mathtt{GradWhitening}$ + rescaling) is presented in Algorithm 1 (under $\mathtt{diag}=\text{True}$). We denote it as SWAN.
+4.3 Saving analysis: balancing memory efficiency and expressiveness
+Table 1 shows the theoretical memory savings of SWAN.
+When applied on a 2D tensor of $m$ by $n$ ($m>n$), SWAN requires storing 00 optimizer states (i.e., intermediate variables maintained over time), compared with other state-of-the-art memory efficient optimizers.
+Notably, this is achieved without significantly compromising its capabilities to model the adaptive learning rates.
+SWAN can model element-wise adaptive learning rates, while the concurrent work (SGD-sal and Apollo) needs to resort to block-wise or even tensor-wise approximations.
+In our experiments in Section 6.1, we show that these approximations indeed has an negative impact on LLM training performance.
 5 Analysis: A LLM Learning Dynamics Perspective
@@
 Assuming we inherit the assumptions in Theorem 1 of Tian et al.
-(2023), as described in Appendix B.
+(2023), as described in Appendix C.
 Then consider ${\mathbf{U}}_{C}^{\top}{\mathbf{W}}$, the composition of the MLP project-up matrix and the embedding matrix as a whole.
@@
 Finally, Theorem 2 presented in Appendix will also show how $\mathtt{GradWhitening}$ helps to make the convergence rate of SGD more robust to the condition number of local curvatures, and outperforms both SGD and Adam in the ill-conditioned regime.
+5.4 Verifying the Theoretical Insights
+We empirically validate the theoretical benefits of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$.
+As detailed in Appendix I, we demonstrate that $\mathtt{GradNorm}$ effectively stabilizes stochastic gradient distributions during SGD training of a scaled-down LLaMA model on the C4 dataset, evidenced by significantly reduced KL divergence fluctuations compared to standard training.
+Additionally, $\mathtt{GradWhitening}$ significantly enhances optimization performance across high-dimensional and ill-conditioned classic optimization functions, consistently outperforming both SGD and Adam.
+These findings confirm that $\mathtt{GradNorm}$ promotes gradient stability and $\mathtt{GradWhitening}$ addresses local curvature challenges, enabling faster and more reliable performance.
 6 Experiments
-In this section, we report empirical results for SWAN. The main purposes are:
-  • Benchmarking the performance, training speed-ups and memory footprint of SWAN on pre-training LLaMA with C4 dataset (Raffel et al., 2020) against Adam.
-Results are presented in Section 6.1, Section 6.2, and Section 6.3.
-  • Investigating the contribution of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ to the performance of SWAN through ablation studies, as outlined in Section 6.4.
-  • Separately evaluating whether $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ achieve their intended purposes of stabilizing gradient noise and modeling/ counteracting local curvature, respectively.
-Detailed results are reported in Section 6.5.1 and Section 6.5.2.
-All experiments run on NVIDIA A100 GPUs.
-6.1 SWAN Performance on LLM Pre-training Tasks
+In this section, we report empirical results for SWAN. All experiments run on NVIDIA A100 GPUs.
+6.1 Memory-Efficient Pre-training of Llama Models
 Setup
-We evaluate SWAN on large language model (LLM) pre-training tasks using a LLaMa-based architecture Touvron et al.
-(2023a) with RMSNorm and SwiGLU activations (Zhang & Sennrich, 2019; Gao et al., 2023).
+We evaluate SWAN on memory-efficient Llama  (Touvron et al., 2023a) pre-training tasks, using the standardized settings of (Zhao et al., 2024a), which has been adopted by many recent works (Zhao et al., 2024a; Chen et al., 2024; Zhang et al., 2024c; Huang et al., ; Zhu et al., 2024).
 We consider models with 60M, 130M, 350M, and 1.3B parameters, all trained on the C4 dataset Raffel et al.
-(2020) using an effective token batch size of 130K tokens (batch size 512, context length 256).
+(2020) using an effective batch size of 130K tokens.
 Following the setup of Zhao et al.
 (2024a), SWAN is applied to all linear modules in both attention and MLP blocks.
-Training uses BF16 precision by default.
-The other evaluation settings are the same as Zhao et al.
+Training uses BF16 by default unless specified, see Appendix J.
+The other evaluation settings follows Zhao et al.
 (2024a).
@@
 Validation perplexity is reported, along with a memory estimate of the total of parameters and optimizer states based on BF16 format.
-Since we used train scripts and configs based on the code opensourced by Zhao et al.
-(2024a), all baseline results are taken from Zhao et al.
-(2024a).
-The exceptions are the results for Adam and Galore, as we cannot exactly reproduce the results reported in Zhao et al.
-(2024a); hence we reported results from our own runs instead.
+Baseline results are directly taken from the official numbers reported in Zhao et al.
+(2024a); Zhu et al.
+(2024), as they shares exactly the same setup.
+Note that for Adam we report both the official results from Zhao et al.
+(2024a) and our reproduced result.
 60M 130M 350M 1.3 B
-SWAN 32.28 (0.12G) 24.13 (0.25G) 18.22 (0.68G) 15.13 (2.60G)
-Adam 33.02 (0.36G) 24.44 (0.76G) 20.11 (2.06G) 16.44 (7.80G)
-Galore 33.09 (0.24G) 24.67 (0.52G) 19.74 (1.22G) 15.89 (4.38G)
-Low-Rank 78.18 (0.26G) 45.51 (0.54G) 37.41 (1.08G) 142.53 (3.57G)
+Adam 33.02 (0.32G) 24.44 (0.75G) 19.24 (2.05G) 16.44 (7.48G)
+Adam (cited) 34.06 (0.32G) 25.08 (0.75G) 18.80 (2.05G) 15.56 (7.48G)
+SWAN 30.59 (0.23G) 22.61 (0.43G) 16.63 (0.93G) 13.56 (2.98G)
+SWAN 30.00 (0.23G) 22.83 (0.43G) 17.14 (0.93G) 14.42 (2.98G)
+SWAN-0 32.28 (0.23G) 24.13 (0.43G) 18.22 (0.93G) 15.13 (2.98G)
+Momentum$+$$\mathtt{GradWhitening}$ 31.6 (0.27G) 24.59 (0.59G) 19.30 (1.49G) 16.08 (5.23G)
+Galore 34.88 (0.26G) 25.36 (0.57G) 18.95 (1.29G) 15.64 (4.43G)
+Apollo-mini 31.93 (0.23G) 23.53 (0.43G) 17.18 (0.93G) 14.17 (2.98G)
+Apollo 31.55 (0.26G) 22.94 (0.57G) 16.85 (1.29G) 14.20 (4.43G)
+Low-Rank 78.18 (0.26G) 45.51 (0.57G) 37.41 (1.29G) 142.53 (4.43G)
 LoRA 34.99 (0.36G) 33.92 (0.80G) 25.58 (1.76G) 19.21 (6.17G)
 ReLoRA 37.04 (0.36G) 29.37 (0.80G) 29.08 (1.76G) 18.33 (6.17G)
-SWAN speed up vs Adam 1.2 X 1.23 X > 2X > 2X
+SWAN speed up vs Adam 1.52 X 1.6 X > 2.3 X > 2.4 X
 $r$ of low-rank methods 128 256 256 512
@@
 Figure 4: Comparison of convergence rate of different methods on LLM pretraining tasks.
+The training curves of Adam, Galore and Apollo-mini are reproduced according to the opensource code of Zhao et al.
+(2024a); Zhu et al.
+(2024).
+We further compare to their official results in Table 2.
+SWAN optimizer
+We consider three variants: SWAN-0, which aims to show-case the robustness and effectiveness of our method out-of-the-box, with almost no tuning.
+It uses naive NS-iteration for whitening, disabled learning rate warmup, and use similar learning rates optimized for Adam.
+SWAN, the vanilla version of our method, in which we enabled learning rate warmup, and allowed the use of optimized learning rates that largely differ from Adam.
+Finally, SWAN, the strongest and most efficient variant that employs the proposed NSDS scheme for fast whitening (section 4.2).
+For all SWAN variants, we adopt a lazy-tuning approach (hyperparameters are set without extensive search) to reduce the possibility of unfair performance distortion.
+Notably, for SWAN, we share the same hyperparameters across all model sizes.
+Full details can be found in Appendix J.
 Baselines
-Since SWAN aims to explore EMA-free adaptive optimizers for LLMs, we focus on comparing it with Adam and other low-rank optimizers. The baselines are:
-  • Adam (Kingma & Ba, 2015): one standard choice for training large models.
-  • Galore Zhao et al. (2024a): a memory-efficient Adam variant with low-rank gradient projections.
-  • Low-Rank (Kamalakara et al., 2022): a low-rank factorization approach (${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$).
-  • LoRA Hu et al. (2021): the LoRA method applied for pre-training as in Zhao et al. (2024a).
-  • ReLoRA Lialin et al. (2023): a full-rank extension of LoRA with parameter merging.
-Hyperparameters
-We use the publicly available hyperparameter configurations from Zhao et al.
-(2024a) for all baselines [^2] This does not include Adam, since the authors of Zhao et al.
-(2024a) did not disclose their Adam configs.
-However, we used the same learning rate sweep procedure as described in their paper.
-This result in an optimal learning rate of 0.001 for 60M, 130M and 350M models, and 0.0007 for 1.3 B models., including method-specific learning rates selected through a grid search.
-For SWAN, we use a default setting without extensive tuning.
-We match the global learning rate (0.001) used by Adam, except for the 1.3B model (Adam diverges at 0.001, see Appendix H).
-We use 10 steps of Newton–Schulz iteration for $\mathtt{GradWhitening}$.
-Unlike Adam, SWAN does not require a warm-up phase to ensure stable training.
-All other scheduling parameters remain the same as in the baselines.
-Full details can be found in Appendix H.
+We focus on comparing SWAN with Adam (Kingma & Ba, 2015) and other memory-efficient optimizers.
+All baseline results are directly quoted from corresponding papers as they all share the same standard setup.
+The baselines include Adam (Kingma & Ba, 2015), which is a standard choice for training large models; and Galore Zhao et al.
+(2024a), a memory-efficient Adam variant with low-rank gradient projections.
+We also consider Low-Rank (Kamalakara et al., 2022), a low-rank factorization approach (${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$); and LoRA Hu et al.
+(2021), which applies the LoRA method for pre-training as in Zhao et al.
+(2024a).
+Additionally, we include ReLoRA Lialin et al.
+(2023), a full-rank extension of LoRA with parameter merging, and Momentum$+$$\mathtt{GradWhitening}$, which applies Newton-Schulz $\mathtt{GradWhitening}$ on top of momentum instead of $\mathtt{GradNorm}$; this is equivalent to Muon Jordan et al.
+(2024) with Nesterov acceleration turned off.
+Finally, we compare with Apollo-mini and Apollo (Zhu et al., 2024).
+Full details can be found in Appendix J.
 Fair Comparison
 We follow Kaddour et al.
-(2024) by allocating the same total training budget (in terms of processed tokens) for all methods.
-By the end of training, all methods decay to the same target learning rate.
-We also leave SWAN’s learning rates largely untuned for a more direct comparison.
+(2024) by allocating the same total training budget (in terms of tokens) for all methods.
+By the end of training, all methods decay to the same target learning rate, which is $10\%$ of the initial learning rates.
+This approach eliminates an often over-looked evaluation bias, where optimizers with earlier learning rate cool-downs tend to have an unfair advantage due to artificially fast convergence caused by learning rate decay Kaddour et al.
+(2024).
 Results
-Table 1 and Figure 4 show the results.
-SWAN achieves lower validation perplexities than the baselines with lower estimated memory consumptions.
-On the 350M and 1.3B models, SWAN reaches over a 2$\times$ speedup (in steps or tokens) relative to Adam (Figure 1 and Figure 5 (a)).
-Further details are in Section 6.2.
-6.2 Is the Improvement Multiplicative or Additive?
-A key question in assessing speedup factors is whether the improvement over Adam is multiplicative or additive.
+As shown in Table 2 and Figure 1, all SWAN variants achieve strong performance compared with baselines, requiring the lowest memory consumptions comparable to SGD.
+Across all models, SWAN-0 surpasses the performance of the Adam and Momentum-$\mathtt{GradWhitening}$ (muon-like) baseline in terms of validation perplexities.
+SWAN further delivers a comparable performance to Apollo series; and finally, SWAN with NSDS scheme delivers the strongest, SOTA-level performance under this setup.
+Notably, on the 350M and 1.3B models, all SWAN variants reaches at least 2$\times$ speedup (in steps or tokens) relative to Adam (Figure 1 and Figure 5 (a)).
+Finally, we observe that in general, SWAN has a slower convergence speed in the early stage than SWAN.
+However, it offers strong tail convergence in return and surpasses SWAN in terms of final performance.
+6.2 Memory Efficiency and Throughput Analysis
+Table 3: Raw and effective throughput analysis, under different model parallelization (MP) settings.
+Method Raw / eff. throughput (1B)
+Adam 117872 / 117872 (tokens/s)
+SWAN w/o MP   58600 / 117200 (tokens/s)
+SWAN w/o MP 107808 / 258739 (tokens/s)
+SWAN w/ MP 114160 / 228320 (tokens/s)
+SWAN w/ MP 115872 / 278092 (tokens/s)
+Memory Footprint
+We compare SWAN, Adam, and Galore on a single A100 GPU.
+Unlike Zhao et al.
+(2024a); Zhu et al.
+(2024), which report layer-wise training memory usage, we measure total end-to-end memory consumption under full-model training using a batch-size of $1$ for 1.3B, 7B, and 13B models.
+As shown in Figure 1 (c), SWAN’s memory usage is on par with SGD, providing nearly a 50% reduction in total memory.
+If we further incorporate the per-layer training technique of (Zhao et al., 2024a; Lv et al., 2023), SWAN further achieves $\approx 70\%$ total memory reduction on top of $\approx 100\%$ reduction on optimizer states.
+This underlines the benefit of the stateless design.
+Effective Throughput
+We assess throughput when training a 1.3B LLama model on 8 A100 GPUs with a batch size of 130K.
+All gradient processing are done in BF16.
+We use two metrics: raw throughput: number of tokens processed per second.
+Effective throughput: raw throughput adjusted by SWAN’s token efficiency relative to Adam.
+These metrics evaluate the impact of different $\mathtt{GradWhitening}$ schemes on training speed, and also account for the fact that some optimizers make more effective use of training tokens.
+As shown in Table 3, SWAN with naive NS $\mathtt{GradWhitening}$ requires model parallelization (where NS of different tensors are distributed to different GPUs) to achieve competitive throughput.
+Without model parallelization, the throughput of SWAN is 50 % lower than Adam.
+With the proposed NSDS scheme, SWAN achieves an raw throughput comparable to Adam, without any need for model parallelization.
+Consequently, SWAN exhibits a 2 $\times$ higher effective throughput than Adam.
+6.3 Is the Speed-up Multiplicative or Additive?
+A key question regarding speedup factors is whether the improvement over Adam is multiplicative or additive.
 A multiplicative speedup implies that the optimizer’s relative advantage remains proportionally consistent over time, while an additive speedup suggests a less desired constant step advantage.
-We explore these scenarios using two plots, the speed-up ratio comparison and the perplexity comparison (Figure 5), across different model sizes.
+To investigate this, we take the SWAN-0 setup as an example, and address this question using two plots, the speed-up ratio comparison and the perplexity comparison (Figure 5), across model sizes.
 (a) Speedup factor of SWAN vs Adam training steps
@@
 $R(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{SWAN{}}}(P)},$ (7)
-where $S_{\text{Adam}}(P)$ and $S_{\text{SWAN{}}}(P)$ are the training steps required by Adam and SWAN, respectively, to reach perplexity $P$.
+where $S_{\text{Adam}}(P)$/$S_{\text{SWAN{}}}(P)$ are the training steps required by Adam/SWAN to reach perplexity $P$.
 Counterfactual Additive Curve Estimation
@@
 $\text{PPL}_{\text{additive}}(S)=\text{PPL}_{\text{Adam}}\left(S+\Delta\right).$ (10)
-This represents the expected perplexity of SWAN if it only consistently outperforms Adam by $\Delta$ steps.
+This represents the expected perplexity of SWAN if it only outpaces Adam by fixed $\Delta$ steps.
 Results
@@
 If SWAN’s actual curves exceed these additive estimates, it indicates a tendency towards a multiplicative speedup, instead of the additive advantage.
-We summarize the observations for model sizes of 60M, 130M, 350M, and 1.3B parameters:
-  • For smaller models (60M and 130M), the actual speedup trajectories align closely with the additive baseline, indicating a primarily additive speedup.
-  • For larger models (350M and 1.3B), the actual curves rise noticeably above the additive estimates, suggesting a multiplicative speedup.
+We summarize the observations across model sizes as: 1) for smaller models (60M and 130M), the actual speedup trajectories align closely with the additive baseline, indicating a primarily additive speedup; 2) for larger models (350M and 1.3B), the actual curves rise noticeably above the additive estimates, suggesting a multiplicative speedup.
 This indicates that SWAN yields increasing efficiency gains as model size grows.
-6.3 Memory Efficiency and Throughput Analysis
-In this section, we measure both the memory footprint and throughput of SWAN.
-Memory Footprint
-We compare SWAN, Adam, and Galore on a single A100 GPU.
-Unlike Zhao et al.
-(2024a), which reports layer-wise training memory usage, we measure total end-to-end memory consumption under full-model training using batch size of $1$ for LLaMA with 1.3B, 7B, and 13B parameters.
-As shown in Figure 1 (c), SWAN’s memory usage is close to that of vanilla SGD, providing nearly a 50% reduction in total memory by removing the needs for first and second moments compared to Adam.
-This observation underlines the benefit of a stateless design.
-Effective Throughput
-We also assess throughput when training a 1.3B-parameter model on 4 A100 GPUs with a token batch size of 130K tokens. We use two metrics:
-  • Absolute throughput: number of tokens processed per second.
-  • Effective throughput: absolute throughput adjusted by SWAN’s token efficiency relative to Adam.
-These metrics evaluate whether $\mathtt{GradWhitening}$ (which uses Newton-Schulz iterations) significantly reduces training speed and also account for the fact that some optimizers make more effective use of training tokens.
-As shown in Figure 1 (d), SWAN achieves an absolute throughput comparable to Adam, indicating that the computational overhead of $\mathtt{GradWhitening}$ is minimal.
-Consequently, SWAN exhibits a higher effective throughput than Adam because of its more efficient use of tokens during training.
-All experiments use BF16 precision for $\mathtt{GradWhitening}$.
 6.4 Ablation Studies
+We take SWAN-0 from the SWAN series as an example and conduct the following ablation studies. All SWAN optimizer mentioned in this section specifically refers to the SWAN-0 setting.
 Effect of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ on Performance
@@
 Removing either results in performance degradation.
-Similarly, Adam also requires first and second moment estimates for optimal performance.
+Similarly, Adam also requires all moments for optimal performance.
 Figure 6: Ablation studies on 130M model.
@@
 These findings suggest that SWAN is more robust to warm-up schedule and can train effectively with or without it.
-6.5 Analysis of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
-6.5.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
-Figure 7: KL divergence comparison of gradient distributions against initial gradient distribution across training Steps.
-We use the projection weights in attention and MLP modules of the second layer as an example.
-The plots compare standard training with $\mathtt{GradNorm}$-augmented training.
-Lower KL divergence values indicate greater stability in gradient distributions.
-To examine whether $\mathtt{GradNorm}$ stabilizes the distribution of the stochastic gradients as suggested by Theorem 1, we conduct controlled experiments using a scaled-down LLaMA-based model (about 10 million parameters) (Lialin et al., 2023), trained on the C4 dataset.
-Our goal is to measure how $\mathtt{GradNorm}$ affects the distribution of stochastic gradients over multiple training steps.
-Specifically, we employ a small-scale LLaMA-based model with approximately 10 million parameters  (Lialin et al., 2023).
-Training is conducted on the C4 dataset (Raffel et al., 2020).
-Baselines
-We compare:
-  • Standard training: This uses an SGD optimizer with a learning rate of $5\times 10^{-4}$ and a linear learning rate scheduler, including a 10% warm-up of total training steps (10,000 steps).
-  • $\mathtt{GradNorm}$-processed training: This applies $\mathtt{GradNorm}$ to pre-process the stochastic gradient before the parameter update.
-All other settings match the standard training baseline.
-Methodology
-We measure gradient statistics in the presence of mini-batch noise.
-At step $t=0$, we sample 16 additional mini-batches (batch size 64 each) and compute the mean and standard deviation of the corresponding raw or $\mathtt{GradNorm}$ gradients in each batch, and obtain the approximated initial gradient distribution.
-After each training step of baseline methods, we perform the same procedure and calculate the Kullback-Leibler divergence between the resulting gradient distributions and the initial gradient distributions.
-This process tracks how the gradient distribution changes over time.
-Results
-Figure 7 shows the KL divergence of gradient distributions for standard and $\mathtt{GradNorm}$-augmented training, relative to the corresponding initial approximated distributions.
-Apart from early spikes, $\mathtt{GradNorm}$ reduces fluctuations in the gradient distribution throughout training.
-6.5.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
-This subsection evaluates the optimization performance of gradient descent when combined with $\mathtt{GradWhitening}$. We use three classic problem settings:
-(a) Quadratic optimization, well-conditioned
-(b) Quadratic optimization, ill-conditioned
-(c) Rastrigin function optimization
-(d) Effect of initialization
-Figure 8: Comparison of convergence rate of different methods on quadratic and non-convex optimization problems.
-(a): 2500-dimensional quadratic optimization with well-conditioned ${\mathbf{H}}$.
-(b): 2500-dimensional quadratic optimization with ill-conditioned ${\mathbf{H}}$ (c): 2500-dimensional Rastrigin function optimization.
-(d): 2500-dimensional Rastrigin function optimization, but forcing all methods to use the same orthogonal initial location.
-  • High-dimensional quadratic optimization. A quadratic problem of the form in Equation 10, where ${\mathbf{W}}\in\mathbb{R}^{50\times 50}$.
-  • Ill-conditioned quadratic optimization. Same setup as above, but with a deliberately chosen ill-conditioned ${\mathbf{H}}$.
-  • Non-convex optimization with multiple local optima. We use the multivariate Rastrigin function:
-    $f({\mathbf{W}})=m^{2}{\mathbf{A}}+\frac{1}{2}\text{Tr}[{\mathbf{W}}^{\top}{\mathbf{W}}]-{\mathbf{A}}\sum_{ij}\text{cos}(2\pi W_{ij}),$
-    where ${\mathbf{W}}$ is an $m\times m$ matrix and $m=50$. This function has $10^{m^{2}}$ possible local optima.
-Baselines
-We compare five methods on all three problems: gradient descent (GD) with the theoretical optimal learning rate in Theorem 3, Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and a hand-tuned learning rate, Newton’s method with a tuned learning rate, and two $\mathtt{GradWhitening}$-based variants (with and without orthogonal initialization).
-This is to verify, under orthogonal initialization, $\mathtt{GradWhitening}$-processed GD behaves similarly to Newton’s method, as discussed in Section 5.3 and Theorem 2.
-All methods share the same initialization, except for the orthogonal $\mathtt{GradWhitening}$ variant, which projects the initial parameters onto an orthogonal matrix.
-Results
-From Figure 8 (a)–(c), we summarizes the following outcomes:
-  • Quadratic problems (Figure 8 (a) and (b)).
-$\mathtt{GradWhitening}$ with orthogonal initialization and Newton’s method converge to optimum in one step, aligning with the theoretical predictions in Section 5.3 and Theorem 2.
-  • Well- vs.
-ill-conditioned cases (Figure 8 (a) and (b)).
-In the well-conditioned setting (a), standard GD outperforms both Adam and $\mathtt{GradWhitening}$ (non-orthogonal initialization).
-In the ill-conditioned setting (b), $\mathtt{GradWhitening}$ (non-orthogonal initialization) outperforms GD by a large margin, while GD experiences slow convergence.
-  • Comparison with Adam (Figure 8 (a)–(c)). In all three settings, $\mathtt{GradWhitening}$ with non-orthogonal initialization consistently outperforms Adam, consistent with Proposition 3.
-  • Rastrigin function (Figure 8 (c)). On this non-convex problem, $\mathtt{GradWhitening}$ performs comparably to Newton’s method, with or without orthogonal initialization.
-  • Effect of initialization (Figure 8 (d)) Furthermore, we force all methods to share the same orthogonalized initialization.
-As shown by the result, $\mathtt{GradWhitening}$ GD with orthogonal initialization still consistently outperforms all baselines, confirming that this initialization is only beneficial to $\mathtt{GradWhitening}$ GD among all baselines.
+7 Open Questions and Extensions
+In Section 5 we analyzed the design choice of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ from the perspective of LLM dynamics.
+However, the compositional effect of these two operators is still an open question, which is not discussed in this work.
+Specifically, in Appendix B we empirically show that when comparing with standard, standalone NS processing of the gradients, both $\mathtt{GradWhitening}$ and (NSDS)-$\mathtt{GradWhitening}$ of SWAN introduces significant changes, rotating the update by a relatively large magnitude.
+This indicates that additional efforts are needed to further explain the novel dynamics of SWAN, as well as the question of “what makes a good preprocessing chain for gradients”.
+After the v1 version of this paper on arXiv, we posted a follow-up work of SWAN (Scetbon et al., 2025), in which we have shown that SWAN is a special case of a general framework called multi-normalized gradient descent (MNGD).
+MNGD aims at normalizing gradients according to multiple norms, generalizing the steepest descent viewpoint of Bernstein & Newhouse (2024b) that recasts popular first-order optimizers as normalization of gradients under a single norm.
 8 Conclusion
@@
 Therefore, we hypothesize that Adam approximately whitens the gradients element-wise, leading to well-conditioned regions.
-Appendix B Additional Analysis
-B.1 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
+Appendix B Acceleration of Newton-Schulz iteration via diagonal substitution
+B.1 Algorithnms
+Computing $\mathtt{GradWhitening}$ exactly can be expensive, as it involves solving the matrix square-root inverse.
+One option is to directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation.
+This is given by (Song et al., 2022; Li et al., 2018):
+$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k})\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}\end{cases}$
+where ${\mathbf{Y}}_{0}={\mathbf{G}}{\mathbf{G}}^{\top}$ , ${\mathbf{Z}}_{0}={\mathbf{I}}$.
+At convergence, $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{Z}}{\mathbf{G}}$ (Algorithm 2).
+However, estimating $({\mathbf{G}}{\mathbf{G}}^{\top})^{-1/2}$ with NS requires $\mathcal{O}(m^{3})$ (assuming $m<n$) complexity.
+Here, we propose a heuristic scheme that has $\mathcal{O}(m^{2})$ complexity to estimate square-root inverse:
+$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}\text{Diag}(3{\mathbf{I}}-{\mathbf{Z}}_{k}\text{Diag}({\mathbf{Y}}_{k})))\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-\text{Diag}({\mathbf{Z}}_{k}){\mathbf{Y}}_{k})\text{Diag}({\mathbf{Z}}_{k})\end{cases}$
+where $\text{Diag}(\cdot)$ returns a diagonal matrix that has the same diagonal elements as the input matrix.
+Basically, whenever we encounter matrix multiplication in NS iterations, we replace one of them by its diagonal approximation.
+We refer to this as the NS with diagonal substitution (NSDS) scheme.
+Note that in the above standard presentation we have fixed that both NS and NSDS uses coefficients $=0.5$ for ${\mathbf{Y}}$ updates and ${\mathbf{Z}}$ updates, respectively.
+In practice we may further tune these coefficients to compensate for short number of iterations (usually under 10).
+B.2 Experiment: LLM Gradient Condition Number Reduction
+Setup
+In this synthetic experiment, we assessed the effectiveness of two whitening methods, Newton-Schulz and the proposed Newton-Schulz with diagonal substitution (NSDS) scheme, on gradient matrices obtained from LLM training.
+As discussed in Section 4.1, the exact $\mathtt{GradWhitening}$ operator results in matrices with optimal condition number ($=1$); therefore, we hereby investigate the matrix condition numbers of the processed gradients obtained from different methods.
+Specifically, both methods use 5 NS iterations with NS step size optimized.
+We train a 130M LLama model following the architecture setting of Section 6.1 on randomly generated sequences for 1000 steps, and take the MLP weights of a middle layer (we take the fifth layer without loss of generality) and use different methods to whiten the gradient matrices.
+We consider three methods: standard NS; NSDS; and SWAN with NSDS (that is, composing $\mathtt{GradNorm}$ with NSDS-$\mathtt{GradWhitening}$).
+At each training step, the condition number reduction ratio of different method was calculated for both whitening methods (the higher the better).
+Note that for all methods, the gradients have been pre-normalized by its norm.
+Results
+Results are shown in Figure 7 (a).
+We notice that NSDS alone (orange curve) is not sufficient to reach a good condition number reduction ratio.
+However, when combined with $\mathtt{GradNorm}$ (i.e., SWAN with NSDS, the green curve), its performance started to catch up and even outperform the standard NS method after 500 training steps.
+This show the effectiveness of the proposed scheme.
+One potential caveat that we spot is that the condition number produced by SWAN with NSDS is more noisy than the standard NS iteration; which might lead to improvements that will be addressed in future work.
+The significance of $\mathtt{GradNorm}$ and NSDS
+[…]
+Figure 7: Comparison between different whitening schemes.
+(a) Performance comparison by condition number reduction ratio (higher the better).
+(b) Cosine similarities between the whitened matrix produced by different preprocessing schemes, respectively.
+B.3 Ablation: effect of NSDS iterations
+We examine the impact of the number of iterations of our Newton Schulz with Diagonal Substitution (NSDS) scheme.
+With SWAN, we compare the test PPL performance on 130M model.
+Results are shown in Table 4.
+As we can see, the improvement brought by additional NSDS iteration is marginal; hence in this paper, we only apply 2 iterations of NSDS.
+Table 4: Effect of NSDS iterations on test PPL. Results are obtain using FP32 precision.
+# NSDS iterations Test PPL
+1 - (LLM loss diverge)
+2 22.63
+5 22.62
+10 22.61
+B.4 Ablation: Precision
+We compare the performance of different precisions of $\mathtt{GradWhitening}$ with NSDS scheme.
+As shown in Table 5, on 130M model (2 step NSDS iterations) ablation we show that BF16 can be used without major performance degrade compared to FP32.
+Table 5: BF16 vs FP 32 NSDS
+NSDS Precision Test PPL on 130M
+BF16 22.61
+FP32 22.63
+Appendix C Additional Analysis
+C.1 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
 In this section, we present main results regarding the convergence rate of the $\mathtt{GradWhitening}$ method, understand its implications, and compare it with the lower bounds of GD and Adam.
@@
 Therefore, we will not discuss the case of $\beta_{2}<1$ to avoid the complication.
-B.2 Numerical Verification of Proposition 1
+C.2 Numerical Verification of Proposition 1
 Given a STB, we consider the following standard full-batch learning dynamics (Tian et al., 2023).
@@
 As predicted by Proposition 1, we see very similar structures across the diagonal blocks of the Hessian.
-Appendix C Proof of Theorem 1
+Appendix D Proof of Theorem 1
 Proof.
@@
 Therefore, after row-wise standardization, the new dynamics becomes
-$\dot{\tilde{{\bm{u}}}}_{i}$ $=\frac{\alpha_{i}(t)(\mu_{i}+{\bm{\xi}}_{i})-\alpha_{i}(t)(\frac{1}{n}\sum_{j}\mu_{ij}+\frac{1}{n}\sum_{j}\xi_{ij})}{\alpha_{i}(t)(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij}-\frac{1}{n}\sum_{j}\mu_{ij}-\frac{1}{n}\sum_{j}\xi_{ij})^{2})}$
-$=\frac{(\mu_{i}+{\bm{\xi}}_{i})-(\frac{1}{n}\sum_{j}\mu_{ij}+\frac{1}{n}\sum_{j}\xi_{ij})}{(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij}-\frac{1}{n}\sum_{j}\mu_{ij}-\frac{1}{n}\sum_{j}\xi_{ij})^{2})}$
+$\dot{\tilde{{\bm{u}}}}_{i}$ $=\frac{\alpha_{i}(t)(\mu_{i}+{\bm{\xi}}_{i})}{\alpha_{i}(t)(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij})^{2})}$
+$=\frac{(\mu_{i}+{\bm{\xi}}_{i})}{(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij})^{2})}$
 Therefore, the normalized noisy gradient $\dot{\tilde{{\bm{u}}}}_{i}$ no longer depend on the time-variant component $\alpha(t)$. Hence, we have proved:
@@
 The corresponding result for $\tilde{{\mathbf{G}}}_{{\bm{z}}_{q}}^{(t)}:=\mathtt{GradNorm}(\frac{\partial{\mathcal{L}}_{{\mathbf{W}},{\bm{z}}_{q}}(\top{\bm{x}}^{(t)})}{{\bm{z}}_{q}})$ can be trivially derived due to Equation 18.
-Appendix D Proof of Theorem 2
+Appendix E Proof of Theorem 2
 Proof.
@@
 Hence, the proof is complete. ∎
-Appendix E Proof of Proposition 2
+Appendix F Proof of Proposition 2
 Proof.
@@
 This inequality holds even as the condition number $\kappa\xrightarrow{}+\infty$. ∎
-Appendix F Proof of Proposition 3
+Appendix G Proof of Proposition 3
 To prove Proposition 3, we first generalize existing work on the convergence rate lower bound (via contraction factor) of gradient descent and Adam (we only consider $\beta_{2}=1$) under the same setting:
@@
 for all $l\in[L]$. ∎
-Appendix G Proof of Proposition 1
+Appendix H Proof of Proposition 1
 Proof.
@@
 This holds for all $1\leq s,s^{\prime}\leq d$ and $1\leq k,k^{\prime}\leq n$. ∎
-Appendix H Implementation details
-Setup
+Appendix I Empirical verification of theoretical insights from Section 5 regarding $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
+I.0.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
+Figure 9: KL divergence comparison of gradient distributions against initial gradient distribution across training Steps.
+We use the projection weights in attention and MLP modules of the second layer as an example.
+The plots compare standard training with $\mathtt{GradNorm}$-augmented training.
+Lower KL divergence values indicate greater stability in gradient distributions.
+To examine whether $\mathtt{GradNorm}$ stabilizes the distribution of the stochastic gradients as suggested by Theorem 1, we conduct controlled experiments using a scaled-down LLaMA-based model (about 10 million parameters) (Lialin et al., 2023), trained on the C4 dataset.
+Our goal is to measure how $\mathtt{GradNorm}$ affects the distribution of stochastic gradients over multiple training steps.
+Specifically, we employ a small-scale LLaMA-based model with approximately 10 million parameters  (Lialin et al., 2023).
+Training is conducted on the C4 dataset (Raffel et al., 2020).
+Baselines
+We compare:
+  • Standard training: This uses an SGD optimizer with a learning rate of $5\times 10^{-4}$ and a linear learning rate scheduler, including a 10% warm-up of total training steps (10,000 steps).
+  • $\mathtt{GradNorm}$-processed training: This applies $\mathtt{GradNorm}$ to pre-process the stochastic gradient before the parameter update.
+All other settings match the standard training baseline.
+Methodology
+We measure gradient statistics in the presence of mini-batch noise.
+At step $t=0$, we sample 16 additional mini-batches (batch size 64 each) and compute the mean and standard deviation of the corresponding raw or $\mathtt{GradNorm}$ gradients in each batch, and obtain the approximated initial gradient distribution.
+After each training step of baseline methods, we perform the same procedure and calculate the Kullback-Leibler divergence between the resulting gradient distributions and the initial gradient distributions.
+This process tracks how the gradient distribution changes over time.
+Results
+Figure 9 shows the KL divergence of gradient distributions for standard and $\mathtt{GradNorm}$-augmented training, relative to the corresponding initial approximated distributions.
+Apart from early spikes, $\mathtt{GradNorm}$ reduces fluctuations in the gradient distribution throughout training.
+I.0.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
+This subsection evaluates the optimization performance of gradient descent when combined with $\mathtt{GradWhitening}$. We use three classic problem settings:
+(a) Quadratic optimization, well-conditioned
+(b) Quadratic optimization, ill-conditioned
+(c) Rastrigin function optimization
+(d) Effect of initialization
+Figure 10: Comparison of convergence rate of different methods on quadratic and non-convex optimization problems.
+(a): 2500-dimensional quadratic optimization with well-conditioned ${\mathbf{H}}$.
+(b): 2500-dimensional quadratic optimization with ill-conditioned ${\mathbf{H}}$ (c): 2500-dimensional Rastrigin function optimization.
+(d): 2500-dimensional Rastrigin function optimization, but forcing all methods to use the same orthogonal initial location.
+  • High-dimensional quadratic optimization. A quadratic problem of the form in Equation 11, where ${\mathbf{W}}\in\mathbb{R}^{50\times 50}$.
+  • Ill-conditioned quadratic optimization. Same setup as above, but with a deliberately chosen ill-conditioned ${\mathbf{H}}$.
+  • Non-convex optimization with multiple local optima. We use the multivariate Rastrigin function:
+    $f({\mathbf{W}})=m^{2}{\mathbf{A}}+\frac{1}{2}\text{Tr}[{\mathbf{W}}^{\top}{\mathbf{W}}]-{\mathbf{A}}\sum_{ij}\text{cos}(2\pi W_{ij}),$
+    where ${\mathbf{W}}$ is an $m\times m$ matrix and $m=50$. This function has $10^{m^{2}}$ possible local optima.
+Baselines
+We compare five methods on all three problems: gradient descent (GD) with the theoretical optimal learning rate in Theorem 3, Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and a hand-tuned learning rate, Newton’s method with a tuned learning rate, and two $\mathtt{GradWhitening}$-based variants (with and without orthogonal initialization).
+This is to verify, under orthogonal initialization, $\mathtt{GradWhitening}$-processed GD behaves similarly to Newton’s method, as discussed in Section 5.3 and Theorem 2.
+All methods share the same initialization, except for the orthogonal $\mathtt{GradWhitening}$ variant, which projects the initial parameters onto an orthogonal matrix.
+Results
+From Figure 10 (a)–(c), we summarizes the following outcomes:
+  • Quadratic problems (Figure 10 (a) and (b)).
+$\mathtt{GradWhitening}$ with orthogonal initialization and Newton’s method converge to optimum in one step, aligning with the theoretical predictions in Section 5.3 and Theorem 2.
+  • Well- vs.
+ill-conditioned cases (Figure 10 (a) and (b)).
+In the well-conditioned setting (a), standard GD outperforms both Adam and $\mathtt{GradWhitening}$ (non-orthogonal initialization).
+In the ill-conditioned setting (b), $\mathtt{GradWhitening}$ (non-orthogonal initialization) outperforms GD by a large margin, while GD experiences slow convergence.
+  • Comparison with Adam (Figure 10 (a)–(c)). In all three settings, $\mathtt{GradWhitening}$ with non-orthogonal initialization consistently outperforms Adam, consistent with Proposition 3.
+  • Rastrigin function (Figure 10 (c)). On this non-convex problem, $\mathtt{GradWhitening}$ performs comparably to Newton’s method, with or without orthogonal initialization.
+  • Effect of initialization (Figure 10 (d)) Furthermore, we force all methods to share the same orthogonalized initialization.
+As shown by the result, $\mathtt{GradWhitening}$ GD with orthogonal initialization still consistently outperforms all baselines, confirming that this initialization is only beneficial to $\mathtt{GradWhitening}$ GD among all baselines.
+Appendix J Implementation details
+General setup
 We describe the implementation setups for SWAN used in LLM pre-training tasks.
 To enable a more straightforward and comparable analysis, we simply replicate the setting of Zhao et al.
-(2024a), under exactly the same model configs and optimizer configs.
-This includes the same model architecture, tokenizer, batch size, context length, learning rate scheduler, etc.
-The only exception is that since SWAN does not need warm-up, we use zero-warm-up ratio for SWAN.
-But we still ensure that the learning rates of all methods are fully decayed to the same target learning rate (0.1 $\times$ initial learning rate).
-Across all model sizes, we use the same hyperparameter settings and the same learning rate tuning procedure.
+(2024a), under exactly the same model configs and optimizer hyperparameter configs, whenever possible.
+This includes the same model architecture, tokenizer, batch size, context length, learning rate scheduler, learning rates, subspace scaling, etc.
+Precision
+[…]
+Learning rate scheduling
+we use exactly the same scheduler as in Zhao et al.
+(2024a), with the exception of SWAN-0, which does not require any learning rate warmup.
+Therefore, for SWAN-0, we directly start with maximum learning rate, and enter the learning rate decay phase, using the same decay parameters as Zhao et al.
+(2024a).
+Reproducing baseline results
+Most baseline results are cited from respective papers (Zhao et al., 2024a; Zhu et al., 2024) as we share the exact same setup.
+We also tried to reproduce their results using the same opensourced code, and generally obtain slightly worse results for Galore, Apollo and Adam for larger models (350M and 1B, see Table 6).
+Therefore in the main paper we only compare with the official results.
+For the reproduced results of Adam, we specify the details below as it was not disclosed in Zhao et al.
+(2024a).
+We use same learning rate tuning procedure as suggested by Zhao et al.
+(2024a) (i.e., performing grid search over $\{0.01,0.005,0.001,0.0005,0.0001\}$).
+We found that the optimal learning rates for Adam is 0.001.
 The only exception is that for a model of size 1.3B: as we already know that a larger model requires smaller learning rates, we conduct a learning search for Adam over a smaller but more fine-grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$.
 As a result, the optimal learning rate found for Adam on 1.3B is 0.0007.
-Applying SWAN on LLMs
+Finally, one baseline that does not exist in the literature is the Momentum + $\mathtt{GradWhitening}$ (Muon-like optimizer without Nestrov acceleration) baseline; and we report our own results.
+We start from the default learning rates used by Muon Jordan et al.
+(2024) and tuned them over a grid of ${0.01,0.02,0.03,0.04,0.05}$.
+Table 6: Reproduced results.
+130M 350M 1.3 B
+Adam (reproduced) 24.44 (0.75G) 19.24 (2.05G) 16.44 (7.48G)
+Apollo-mini (reproduced) 23.97 (0.43G) 17.60 (0.93G) 14.37 (2.98G)
+Galore (reproduced) 24.67 (0.57G) 19.74 (1.29G) 15.89 (4.43G)
+$r$ of low-rank methods 256 256 512
+Training Steps 20K 60K 100K
+SWAN Settings and hyperparameters
 Since SWAN utilizes matrix-level operations on gradients, it can only be applied to 2D parameters.
@@
 Similar to Galore (Zhao et al., 2024a), the rest of the non-linear parameters still uses Adam as the default choice.
-Therefore, we follow the learning rate setup of Galore, where we fix a global learning rate 0.001 across all model sizes and all modules.
+Therefore, we follow the learning rate setup of Galore, where we fix some global learning rate across all model sizes and all modules.
 Then, for the linear projection modules where SWAN is applied, we simply apply a scaling factor $\alpha$ on top of the global learning rate.
-We choose $\alpha=1$ for all model sizes except for 1.3 B case.
+For all SWAN variants, we adopt a lazy-tuning approach (hyperparameters are set without extensive search), as detailed below.
+This helps to reduce the possibility of unfair performance distortion due to excessive tuning.
+  • SWAN-0 uses naive NS-iteration for whitening, disabled learning rate warmup, and use similar learning rates optimized for Adam.
+We fix the global learning rate to be the same as Adam, and fix $\alpha=1$.
+The only exception is the 1.3 B case.
 This is because we observe that the optimal learning rate of Adam under 1.3B becomes smaller than 0.001, hence we also reduce the learning rate on SWAN, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$.
-Note that this choice is quite arbitrary, and it is not extensively grid-searched.
-This is to show how the method work out-of-the box, without extensive problem-specific tuning.
-Implementation of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
-For $\mathtt{GradNorm}$, we simply set $\bar{{\bm{g}}}=0$ to simplify computation, as we found in practice this often has small scales.
+To summarize the hyperparameter of SWAN-0 is set to be similar to Adam, without any tuning.
+This is to demonstrate the robustness of SWAN series optimizers and their capability to work out-of-the-box as a replacement for Adam.
+  • SWAN, is the vanilla version of our method, in which we enabled learning rate warmup, and allowed the use of optimized learning rates that largely differ from Adam.
+We notice that SWAN allows larger learning rates than Adam.
+We use a global learning rate of 0.02, as well as the scaling factor $\alpha=0.05$.
+This is selected by simply searching the learning rate over a constraint grid $\{0.01,0.02,0.05\}$, and then setting $\alpha=0.05$ such that the effective learning rate is scaled back to 0.001.
+There is no guarantee that this heuristic rule is optimal; but we found that this usually does not make a run fail (e.g., with loss divergence).
+  • Finally, SWAN, the most efficient version of SWAN that employs the proposed NSDS scheme for fast whitening (section 4.2).
+Similar to SWAN, we use the same global learning rate of 0.02, as well as the scaling factor $\alpha=0.05$ across all model sizes.
+We suspect with more careful tuning, its performance can be significantly improved; however, this is out of the scope of the paper.
+The configurations of $\mathtt{GradWhitening}$ is discussed next.
+Implementation of $\mathtt{GradWhitening}$
 For $\mathtt{GradWhitening}$, before N-S iteration, we further normalize its input matrix by its Frobenius norm.
 This is applied in all our ablation studies as well, regardless of whether the gradient is processed by $\mathtt{GradNorm}$.
-For the N-S iteration, we found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence.
-We set $\beta=0.8$ with iteration number $k=10$ and it generally gives satisfactory performance.
-$\mathtt{GradWhitening}$ can be performed in both FP32 and BF16 format.
-In FP32 format, we first convert the gradient matrice to FP32, perform $\mathtt{GradWhitening}$, and then convert back to BF16.
-This is the default setup in Section 6.1 unless stated otherwise.
-Alternatively, $\mathtt{GradWhitening}$ can also be done in BF16 precision.
-The setup we use for BF16 precision is $\beta=0.85$, $k=6$.
-These hyperparameters are selected on randomly sampled matrices via minimizing inverse square root estimation loss; then directly applied to LLM without any further tuning.
-We found that this usually translates well.
-Note that this might be further optimized similar to (Jordan et al., 2024), we leave it to future work as N-S performance optimization is orthogonal to our key contribution.
+For our proposed Newton-Schulz with Diagonal Substitution (NSDS) used in SWAN, we only run it for 2 steps across all model sizes.
+We found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence.
+We set $\beta=0.4$.
+We found that NSDS is generally robust to $\beta$ as long as it is not too large; and our specific choice of parameters usually already gives satisfactory performance in LLM pretraining.
+For the naive N-S iteration used in SWAN-0 and SWAN, we run 10 steps which is usually sufficient.
+We set $\beta=0.8$.
+The naive NS is run in FP32 precision, on top of BF16 gradients and weights; while the NSDS is run in BF16.
+Computational Overhead.
+[…]
+However, note that whether the above analysis hold for large scale, distributed LLM training is still unclear.
+This is due to a few factors.
+First, the N-S iteration is after all a $\mathcal{O}(m^{3}$ complexity operation and might need to scale as model size increases.
+Second, the distributed computation of N-S steps needs extra non-trivial effort, which brings additional infrastructure challenges.
+This motivates us to propose the NSDS scheme, which enables Adam-level throughput without performance compromise.